### v1

# Quantum Pufferfish Privacy: A Flexible Privacy Framework for Quantum Systems

Abstract
We propose a versatile privacy framework for quantum systems, termed quantum pufferfish privacy (QPP). Inspired by classical pufferfish privacy, our formulation generalizes and addresses limitations of quantum differential privacy by offering flexibility in specifying private information, feasible measurements, and domain knowledge. We show that QPP can be equivalently formulated in terms of the Datta–Leditzky information spectrum divergence, thus providing the first operational interpretation thereof. We reformulate this divergence as a semi-definite program and derive several properties of it, which are then used to prove convexity, composability, and post-processing of QPP mechanisms. Parameters that guarantee QPP of the depolarization mechanism are also derived. We analyze the privacy-utility tradeoff of general QPP mechanisms and, again, study the depolarization mechanism as an explicit instance. The QPP framework is then applied to privacy auditing for identifying privacy violations via a hypothesis testing pipeline that leverages quantum algorithms. Connections to quantum fairness and other quantum divergences are also explored and several variants of QPP are examined.
Index Terms:
Auditing privacy, privacy-utility tradeoff, pufferfish privacy, quantum differential privacy.
I Introduction
Figure 1: Representation of a setup in which the goal is to hide the resource value $\mathsf{V}$ being $a$ or $b$ on the set of states $\{\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}\}$. In this abstract diagram, large squares represent the entire quantum state, while small rectangles correspond to a specific attribute of the quantum state. The specific attribute can take on one of two values, either $a$ or $b$, where $a$ is represented by solid lines and $b$ is represented by dotted lines.
With a surging interest in quantum and classical–quantum systems, ensuring privacy of both classical and quantum data has become pivotal. Privacy-preserving data analysis has been widely studied for classical systems by means of statistical privacy frameworks. Differential privacy (DP) is an important such framework, which enables answering aggregate queries about a database while keeping individual records private [1, 2]. However, DP accounts for one type of private information only (namely, records of individual users), and it does not allow encoding domain knowledge into the framework. To address these limitations, a versatile generalization of DP, termed Pufferfish Privacy (PP), has been proposed [3]. PP allows for customizing which information is regarded as private and explicitly integrates distributional assumptions into the definition [3, 4, 5]. Information-theoretic formulations of classical DP and PP have been proposed in [6] and [7], respectively.
Quantum DP (QDP) is a generalization of the classical DP notion and has been proposed in [8]. See also [9] for DP of quantum measurements and [10] for an information-theoretic interpretation of QDP. Connections to quantum stability through private learning have been studied in [11, 12]. Moreover, [13] has explored how quantum classifiers can be made private by using the intrinsic noise of existing quantum systems. However, similar to the classical case, the versatility of QDP is limited.
In this paper, we propose a flexible privacy framework for quantum systems, termed quantum PP (QPP), that addresses these limitations. We provide a comprehensive study of QPP, encompassing properties, mechanisms, privacy-utility tradeoffs, as well as the first operational meaning of the Datta–Leditzky information spectrum divergence [14] (hereafter abbreviated as the DL divergence), which arises from our framework.
I-A Motivation
We seek to address key limitations of QDP by exploring more flexible privacy frameworks for quantum information processing. As delineated next, flexible secrets, embedding domain knowledge, and relaxing the need for worst-case measurements are considerations central to our approach.
Flexible secrets. QDP guarantees that any pair of states that are classified as neighbors are approximately indistinguishable, i.e., cannot be identified under any possible measurement. However, scenarios may arise in which one wants to hide specific properties of the states, as opposed to the state itself (e.g., whether the states possess a certain symmetry or property, have any entanglement with a special subsystem, or have secret correlations with other systems). In such situations, QDP may be an overly pessimistic notion of privacy, which, in turn, hinders utility. As an example, consider hiding a resource value $\mathsf{V}$ of the states $\{\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}\}$, for which $\mathsf{V}(\rho_{1})=\mathsf{V}(\rho_{2})=a$ and $\mathsf{V}(\sigma_{1})=\mathsf{V}(\sigma_{2})=b$. As illustrated in Fig. 1, hiding whether $\mathsf{V}$ equals $a$ or $b$ amounts to making the classes $\{\rho_{1},\rho_{2}\}$ and $\{\sigma_{1},\sigma_{2}\}$ indistinguishable. This can be achieved by applying a QDP mechanism to the state space $\{\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}\}$, but doing so provides a stricter guarantee than required. The source of the issue is the inability of QDP to account for secrets concerning collections of states (as opposed to singletons), which is the first issue we aim to address.
Domain knowledge. In QDP, a worst-case privacy guarantee is provided for all neighboring states. However, one may possess knowledge about the likelihood of observing different states, e.g., via expert feedback. Referring back to the setting from Fig. 1, if we have domain knowledge such that observing the states $\rho_{1},\rho_{2},\sigma_{1},\sigma_{2}$ is prescribed by the probability vector $(p/2,\left(1-p\right)/2,1/2,0)$, for $p\in(0,1)$, then the requirement simplifies to the indistinguishability of $\{\rho_{1},\rho_{2}\}$ versus $\{\sigma_{1}\}$. Classically, it has been demonstrated that domain knowledge can be leveraged to design privacy mechanisms with increased accuracy and utility [15, 3]. This calls for a quantum privacy framework that can also encode domain knowledge.
Relaxing worst-case measurements. Another worst-case aspect of QDP is its account of all possible measurements. However, such a requirement might be too stringent in practice, especially in quantum systems. As an example, while a joint measurement can accurately distinguish between entangled but physically separated states, oftentimes only local operations and classical communications (LOCC) are available (e.g., as considered in quantum data-hiding protocols [16, 17, 18, 19, 20, 21, 22]). In such cases, one may achieve improved accuracy and utility by relaxing the privacy requirement to account for LOCC measurements only.
In sum, the rapid advancements in quantum technologies requires designing flexible privacy frameworks that can be adjusted to timely needs. Furnishing such a framework is the main objective of our paper.
I-B Contributions
This work proposes a quantum analog of the PP framework that accounts for the three aforementioned aspects. Our formalism enables reasoning about privacy of quantum systems using information-theoretic tools. We provide a comprehensive study of QPP, encompassing properties, mechanisms, and privacy-utility tradeoffs. Our paradigm also gives rise to the first operational interpretation of the DL divergence [14]. The proposed QPP framework comprises four key ingredients:
  1. 1.
    the set of potential secrets,
  2. 2.
    the set of discriminative pairs that are required to be indistinguishable at the output of the mechanism,
  3. 3.
    the set of data distributions, which encodes domain knowledge on the occurrence of quantum or classical data, and
  4. 4.
    the set of measurements to be accounted for, which is specified based on physical, ethical, or any other constraints.
See Definition 4 for a formal definition. QPP guarantees the indistinguishability, under any allowable measurement, of sets of states formed based on the above ingredients.
After defining the operational privacy framework, we observe that when the measurement class contains all possible measurements, QPP can be equivalently posed as a DL divergence constraint. To the best of our knowledge, this provides the first operational interpretation of the DL divergence. We then derive an efficiently computable formulation of the DL divergence as a semi-definite program (SDP), which may be of independent interest. This SDP is utilized to prove properties of the DL divergence, which are then used in the analysis of QPP mechanisms. These properties include joint quasi-convexity and the data-processing inequality under positive and trace non-increasing maps (see Section IV). Our results also generalize the connection between the hockey-stick divergence and QDP, originally established in [10]. Moreover, we show that existing privacy frameworks such as classical DP [1, 2], classical PP [3], utility-optimized local DP (not subsumed by classical PP) [23], and QDP [8, 10] are special cases of our QPP framework.
We then move on to derive properties of QPP mechanisms, encompassing convexity, post-processing, and composability (both parallel and adaptive). As a specific example, we characterize the flip parameter that guarantees QPP of the depolarization mechanism. We also describe how QPP mechanisms implementable on quantum devices can be instantiated to achieve classical PP. We consider the associated privacy-utility tradeoff for QPP mechanisms. Our utility metric captures how invertible the privacy mechanism is, which is formulated as the infimized diamond distance between a post-processing of the mechanism’s output and the identity channel. We show that this utility metric can be computed as an SDP and analyze the privacy-utility tradeoff of the depolarization mechanism. Lastly, we study optimal privacy-utility tradeoffs of QPP mechanisms and characterize the achievable region in several settings.
Another application we consider is privacy auditing, which refers to certifying whether a black-box mechanism satisfies a target privacy guarantee. While several auditing methods are available for classical frameworks, there is currently no approach that can handle quantum data. We fill this gap by proposing the first auditing pipeline for quantum privacy mechanisms. In contrast to existing approaches for classical DP and PP, which require first relaxing the privacy notion and only then auditing, our approach audits for QDP directly. Extensions of these ideas to the QPP setting are also considered.
Finally, we explore connections between QPP, existing quantum privacy frameworks, and figures of merit. First, we examine the connection between quantum privacy and fairness [24, 25], showing that private quantum mechanisms are fair, and under certain conditions, fair algorithms are private. We also provide bounds on quantum Rényi divergences and the trace distance, which stem from QPP. This inspires relaxations of QPP that are defined via these divergences, which, in particular, provide other operational interpretations thereof as privacy metrics. Lastly, we present a variant of QPP that can incorporate entanglement into the framework with the use of reference systems.
I-C Organization
The rest of our paper is organized as follows. In Section II, we introduce notation and preliminaries in quantum information theory and privacy. Section III presents the QPP framework, its equivalent formulations in terms of the DL divergence, and special cases of it. In Section IV, we focus further on the DL divergence, reformulate it as an SDP, and use this SDP to prove several properties of it. Properties of general QPP mechanisms and the depolarization mechanism are studied in Section V. We analyze the privacy-utility tradeoff of QPP mechanisms in Section VI, while the privacy auditing application is considered in Section VII. Connections to existing privacy frameworks and to other quantum divergences are explored in Section VIII. Then, we propose several relaxations and variants of QPP in Section IX. Section X summarizes our main contributions and provides concluding remarks.
II Preliminaries and Background
II-A Notation
Sets are denoted by calligraphic letters, e.g., $\mathcal{X}$. For $k,n\in\mathbb{N}$, we use $\mathcal{X}^{n\times k}$ to denote the database space of $n\times k$ matrices; columns correspond to different attributes while rows to different individuals. The $(i,j)$th entry of $x\in\mathcal{X}^{n\times k}$ is denoted as $x(i,j)$. The $i$th row and $j$th column of $x$ are denoted by $x(i,\cdot)$ and $x(\cdot,j)$, respectively.
We denote by $(\Omega,\mathcal{F},\mathbb{P})$ the underlying probability space on which all random variables (RVs) are defined, with $\mathbb{E}$ designating expectation. RVs are denoted by upper case letters, e.g., $X$, with $P_{X}$ representing the corresponding probability law. For $X\sim P_{X}$, we interchangeably use $\mathrm{supp}(X)$ and $\mathrm{supp}(P_{X})$ for the support. The joint law of $(X,Y)$ is denoted by $P_{XY}$, while $P_{Y|X}$ designates the (regular) conditional probability of $Y$ given $X$. Conventions for $n\times k$-dimensional random variables are the same as for deterministic elements. The space of all Borel probability measures on $\mathcal{S}\subseteq\mathbb{R}^{d}$ is denoted by $\mathcal{P}(\mathcal{S})$. The Kullback–Leibler (KL) divergence between $P,Q\in\mathcal{P}(\mathcal{X})$ with $P\ll Q$ is given by $\mathsf{D}(P\|Q)\coloneqq\mathbb{E}_{P}\!\left[\ln\!\left(\frac{dP}{dQ}\right)\right],$ where $\frac{dP}{dQ}$ is the Radon–Nikodym derivative of $P$ with respect to $Q$. For $(X,Y)\sim P_{XY}$, the mutual information between $X$ and $Y$ is denoted by $\mathsf{I}(X;Y)\coloneqq\mathsf{D}(P_{XY}\|P_{X}\otimes P_{Y})$.
We now review basic concepts from quantum information theory and refer the reader to [26, 27] for more details. A (classical or quantum) system $R$ is identified with a finite-dimensional Hilbert space $\mathcal{H}_{R}$. We denote the set of linear operators acting on $\mathcal{H}_{R}$ by $\mathcal{L}(\mathcal{H}_{R})$. The support of a linear operator $X\in\mathcal{L}(\mathcal{H}_{R})$ is defined to be the orthogonal complement of its kernel, and it is denoted by $\mathrm{supp}(X)$. Let $\mathrm{T}(C)$ denote the transpose of $C$. The partial transpose of $C\in\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ on the subsystem $A$ is represented as $\mathrm{T}_{A}(C)$. Let $\mathrm{Tr}\!\left[C\right]$ denote the trace of $C$, and let $\mathrm{Tr}_{A}\!\left[C\right]$ denote the partial trace of $C$ over the subsystem $A$. The trace norm of a matrix $B$ is defined as $\left\|B\right\|_{1}\coloneqq\mathrm{Tr}[\sqrt{B^{\dagger}B}]$. For operators $A$ and $B$, the notation $A\geq B$ indicates that $A-B$ is a PSD operator, while $A>B$ indicates that $A-B$ is a positive definite operator.
A quantum state $\rho_{R}\in\mathcal{L}(\mathcal{H}_{R})$ on $R$ is a positive semi-definite (PSD), unit-trace operator acting on $\mathcal{H}_{R}$. We denote the set of all density operators in $\mathcal{L}(\mathcal{H}_{R})$ by $\mathcal{D}(\mathcal{H}_{R})$. A state $\rho_{R}$ of rank one is called pure, and we may choose a normalized vector $|\psi\rangle\in\mathcal{H}_{R}$ satisfying $\rho_{R}=|\psi\rangle\!\langle\psi|$ in this case. Otherwise, $\rho_{R}$ is called a mixed state. By the spectral decomposition theorem, every mixed state can be written as a convex combination of pure, orthogonal states. A quantum channel $\mathcal{N}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ is a linear, completely-positive and trace-preserving (CPTP) map from $\mathcal{L}(\mathcal{H}_{A})$ to $\mathcal{L}(\mathcal{H}_{B})$. We denote the adjoint of $\mathcal{N}$ by $\mathcal{N}^{\dagger}$. A measurement of a quantum system $R$ is described by a positive operator-valued measure (POVM) $\{\mathrm{M}_{y}\}_{y\in\mathcal{Y}}$, which is defined to be a collection of PSD operators satisfying $\sum_{y\in\mathcal{Y}}\mathrm{M}_{y}=\mathrm{I}_{\mathcal{H}_{R}}$, where $\mathcal{Y}$ is a finite alphabet. The Born rule dictates that, after applying the above POVM to $\rho\in\mathcal{D}(\mathcal{H}_{R})$, the probability of observing the outcome $y$ is given by $\mathrm{Tr}\!\left[\mathrm{M}_{y}\rho\right]$.
II-B Quantum Divergences
We define several quantum divergences that will be used throughout this work. We call a distinguishability measure $\mathsf{D}(\cdot\|\cdot)$ a generalized divergence [28] if it satisfies the data-processing inequality; i.e., for every channel $\mathcal{N}$, state $\rho$, and PSD operator $\sigma$,
$\boldsymbol{\mathsf{D}}(\rho\|\sigma)\geq\boldsymbol{\mathsf{D}}\!\left(\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\right).$ (1)
The normalized trace distance between the states $\rho$ and $\sigma$ is defined as
$\mathsf{T}(\rho,\sigma)\coloneqq\frac{1}{2}\left\|\rho-\sigma\right\|_{1},$ (2)
while the fidelity between them is defined as [29]
$\mathsf{F}(\rho,\sigma)\coloneqq\left\|\sqrt{\rho}\sqrt{\sigma}\right\|^{2}_{1}.$ (3)
The diamond distance between the two channels $\mathcal{N},\mathcal{M}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ is defined as [30]
$\left\|\mathcal{N}-\mathcal{M}\right\|_{\diamond}\coloneqq\sup_{\rho_{RA}}\left\|\mathcal{N}_{A\to B}(\rho_{RA})-\mathcal{M}_{A\to B}(\rho_{RA})\right\|_{1},$ (4)
where the optimization in the definition is over every bipartite density operator $\rho_{RA}$, with an arbitrarily large system $R$. It is well known, however, that it suffices to perform the optimization over pure bipartite states such that the dimension of the reference system $R$ is equal to the dimension of the channel input system $A$.
The Petz–Rényi quantum relative entropy of order $\alpha\in(0,1)\cup(1,\infty)$ of a state $\rho$ with respect to a PSD operator $\sigma$ is given by [31, 32]
$\mathsf{D}_{\alpha}(\rho\|\sigma)\coloneqq\\ \begin{cases}\frac{1}{\alpha-1}\ln\mathrm{Tr}[\rho^{\alpha}\sigma^{1-\alpha}],&\mbox{if }\alpha\in(0,1)\lor\\ &(\alpha>1\land\mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma))\\ +\infty,&\mbox{otherwise}\end{cases}$ (5)
The special case of $\alpha\to 1$ is called the quantum relative entropy and amounts to
$\mathsf{D}(\rho\|\sigma)\equiv\mathsf{D}_{1}(\rho\|\sigma)\coloneqq\lim_{\alpha\to 1}\mathsf{D}_{\alpha}(\rho\|\sigma)=\mathrm{Tr}\!\left[\rho(\ln\rho-\ln\sigma)\right]$ (6)
when $\mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma)$ and it is equal to +$\infty$+ ∞ otherwise. The quantum entropy of a state $\rho$ is defined as
$\mathsf{S}(\rho)\coloneqq-\mathrm{Tr}\!\left[\rho\ln\rho\right].$ (7)
Equivalently, $\mathsf{S}(\rho)=-\mathsf{D}_{1}(\rho\|I)$, where $I$ is the identity operator.
Fix $\alpha\in(0,1)\cup(1,\infty)$. The sandwiched Rényi relative entropy of a state $\rho$ and a PSD operator $\sigma$ is defined as [33, 34]
$\widetilde{D}_{\alpha}(\rho\|\sigma)\coloneqq\\ \begin{cases}\frac{1}{\alpha-1}\ln\widetilde{Q}_{\alpha}(\rho\|\sigma)&\mbox{if}\ \alpha\in(0,1)\,\lor\\ &(\alpha\in(1,\infty)\land\ \mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma)),\\ +\infty&\mbox{otherwise},\end{cases}$ (8)
where
$\widetilde{Q}_{\alpha}(\rho\|\sigma)\coloneqq\mathrm{Tr}\!\left[\left(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}}\right)^{\alpha}\right].$ (9)
It is a generalized divergence for $\alpha\in[1/2,1)\cup(1,\infty)$ [35] (see also [36]).
Fix $\delta\in[0,1]$, a state $\rho$, and a PSD operator $\sigma$. The Datta–Leditzky information spectrum divergences are defined as follows [14]:
$\underline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $\coloneqq\sup\left\{\gamma\in\mathbb{R}:\operatorname{Tr}[\left(\rho-e^{\gamma}\sigma\right)_{+}]\geq 1-\delta\right\},$ (10a)
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $\coloneqq\inf\left\{\gamma\in\mathbb{R}:\operatorname{Tr}[\left(\rho-e^{\gamma}\sigma\right){}_{+}]\leq\delta\right\},$ (10b)
where
$\left(A\right)_{+}\coloneqq\sum_{i:a_{i}\geq 0}a_{i}|i\rangle\!\langle i|$ (11)
for a Hermitian operator $A~{}=~{}\sum_{i}a_{i}|i\rangle\!\langle i|$. Hereafter we abbreviate these divergences as DL divergences. Proposition 4.3 of [14] shows that
$\underline{\mathsf{D}}^{\delta}(\rho\|\sigma)=\overline{\mathsf{D}}^{1-\delta}(\rho\|\sigma),$ (12)
and so we can speak of a single DL divergence, which we set hereafter to be $\overline{\mathsf{D}}^{\delta}$ from (10b). Slightly rewriting (II-B), we have the equivalent representations:
$\underline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $=\ln\sup\left\{\lambda\geq 0:\operatorname{Tr}[\left(\rho-\lambda\sigma\right)_{+}]\geq 1-\delta\right\}$ (13a)
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $=\ln\inf\left\{\lambda\geq 0:\operatorname{Tr}[\left(\rho-\lambda\sigma\right)_{+}]\leq\delta\right\}.$ (13b)
The latter forms are helpful for making connections to privacy.
The max-relative entropy of a state $\rho$ and a PSD operator $\sigma$ is defined as [37]
$\mathsf{D}_{\max}(\rho\|\sigma)$ $\coloneqq\ln\inf\left\{\lambda:\rho\leq\lambda\sigma\right\}$ (14)
$=\log\sup_{0\leq\mathrm{M}\leq\mathrm{I}}\frac{\mathrm{Tr}\!\left[\mathrm{M}\rho\right]}{\mathrm{Tr}\!\left[\mathrm{M}\sigma\right]},$ (15)
and the smooth max-relative entropy is defined for $\delta\in[0,1]$ as
$\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)\coloneqq\inf_{\widetilde{\rho}\,:\frac{1}{2}\left\|\widetilde{\rho}-\rho\right\|_{1}\leq\delta}\mathsf{D}_{\max}(\widetilde{\rho}\|\sigma),$ (16)
with the optimization taken over every state $\widetilde{\rho}$. These quantities have been given an operational meaning in [38].
The Thompson metric [39] is defined in terms of the max-relative entropy as
$\mathsf{D}_{T}(\rho\|\sigma)\coloneqq\max\{\mathsf{D}_{\max}(\rho\|\sigma),\mathsf{D}_{\max}(\sigma\|\rho)\},$ (17)
and it has been given an operational meaning in [40, 41].
II-C Classical and Quantum Privacy
In this section, we provide background on the existing definitions of privacy for both classical and quantum systems, starting from classical DP and proceeding to quantum DP.
II-C1 Classical Differential and Pufferfish Privacy
DP allows for answering queries about aggregate quantities while protecting the individual entries in a database [1]. To this end, the output of a differentially privacy mechanism should be indistinguishable for neighboring databases, defined as those that differ only in a single record (row). Formally, we say that $x,x^{\prime}\in\mathcal{X}^{n\times k}$ are neighbors, denoted $x\sim x^{\prime}$, if $x(i,\cdot)\neq x^{\prime}(i,\cdot)$ for some $i\in\{1,\ldots,n\}$, and they agree on all other rows. We also note that a randomized privacy mechanism $A$, as mentioned below, is described by a (regular) conditional probability distribution $P_{A|X}$ for its output given the data.
Definition 1 (Classical differential privacy).
Fix $\varepsilon\geq 0$ and $\delta\in[0,1]$. A randomized mechanism ${A}:\mathcal{X}^{n\times k}\to\mathcal{Y}$ is $(\varepsilon,\delta)$-differentially private if
$\mathbb{P}\big{(}A(x)\in\mathcal{B}\big{)}\leq e^{\varepsilon}\hskip 2.84526pt\mathbb{P}\big{(}A(x^{\prime})\in\mathcal{B}\big{)}+\delta,$ (18)
for all $x\sim x^{\prime}$ with $x,x^{\prime}\in\mathcal{X}^{n\times k}$ and $\mathcal{B}\subseteq\mathcal{Y}$ measurable.
As is evident from the above definition, DP aims to conceal whether any particular individual (row) is in fact part of the database or not. While being a powerful and widely applicable privacy framework, it is often appropriate to consider even broader frameworks. Pufferfish privacy [3] is a versatile generalization of DP that not only allows flexibility in the definition of secrets but also enables the integration of domain knowledge of the database space $\mathcal{X}^{n\times k}$. The PP framework consists of three components:
  1. 1.
    A set of secrets $\mathcal{S}\subseteq\mathcal{X}^{n\times k}$ of measurable subsets;
  2. 2.
    A set of secret pairs $\mathcal{Q}\subseteq\mathcal{S}\times\mathcal{S}$ that need to be indistinguishable in the $(\varepsilon,\delta)$ sense (cf., (19) below),
  3. 3.
    A class of data distributions $\Theta\subseteq\mathcal{P}(\mathcal{X}^{n\times k})$ that captures prior beliefs or domain knowledge.
As formulated next, PP aims to guarantee that all secret pairs in $\mathcal{Q}$ are indistinguishable with respect to the prior beliefs $P_{X}\in\Theta$.
Definition 2 (Classical pufferfish privacy).
Fix $\varepsilon\geq 0$ and $\delta\in[0,1]$. A randomized mechanism ${A}:\mathcal{X}^{n\times k}\to\mathcal{Y}$ is $(\varepsilon,\delta)$-private in the pufferfish framework $(\mathcal{S},\mathcal{Q},\Theta)$ if
$\mathbb{P}\big{(}A(X)\in\mathcal{B}\big{|}\mathcal{R}\big{)}\leq e^{\varepsilon}\hskip 2.84526pt\mathbb{P}\big{(}A(X)\in\mathcal{B}\big{|}\mathcal{T}\big{)}+\delta,$ (19)
for all $P_{X}\in\Theta$, $(\mathcal{R},\mathcal{T})\in\mathcal{Q}_{c}$ with $P_{X}(\mathcal{R}),P_{X}(\mathcal{T})>0$, and $\mathcal{B}\subseteq\mathcal{Y}$ measurable.
DP from Definition 1 is a special case of PP when $\mathcal{S}=\mathcal{X}^{n\times k}$, the set $\mathcal{Q}$ contains all neighboring pairs of databases, and $\Theta=\mathcal{P}(\mathcal{X}^{n\times k})$ (i.e., there are no distributional assumptions, and privacy is guaranteed in the worst case). Other important examples that are subsumed by PP include (i) generic DP [42], which allows for arbitrary neighboring relationships, and (ii) attribute privacy [5], which privatizes global properties of a database (e.g., a column that corresponds to some sensitive information, such as salary).
II-C2 Quantum Differential Privacy
QDP lifts the notion of DP to the space of quantum states, with the neighboring relation typically defined either in terms of closeness in trace distance [8] or reachability by a single local operation [9]. [^1] Given two quantum states $\rho$ and $\sigma$ of $n$ registers each, call them neighbors if it is possible to reach either $\sigma$ from $\rho$ or $\rho$ from $\sigma$ by performing a general quantum channel on a single register only. We denote two states being neighbors by $\rho\sim\sigma$.
Definition 3 (Quantum differential privacy [8, 10]).
Fix $\varepsilon\geq 0$ and $\delta\in[0,1]$. Let $\mathcal{D}$ be a set of quantum states, and let $\mathcal{A}$ be a quantum algorithm (viz., a quantum channel). The algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private if
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho)\right]\leq e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\sigma)\right]+\delta.$ (20)
for every measurement operator $\mathrm{M}$ (i.e., satisfying $0\leq\mathrm{M}\leq\mathrm{I}$) and all $\rho,\sigma\in\mathcal{D}$ such that $\rho\sim\sigma$.
This definition reduces to classical DP for discrete-output mechanisms with an appropriate choice of the measurement set. See Section III-C3 below for further details.
III Quantum Pufferfish Privacy (QPP)
Inspired by the versatility of the classical PP framework, we propose a quantum variant thereof. Termed QPP, our framework allows for customizing the notion of private states, tailoring the feasible set of measurements to the application of interest, and incorporating domain knowledge of the state distribution into the model. As such, the QPP framework can generate a rich class of privacy definitions for both classical and quantum systems, and for hybrid classical–quantum systems as well.
III-A Framework
The QPP framework requires a domain expert to specify four components: a set $\mathcal{S}$ of potential secrets, a set $\mathcal{Q}\subseteq\mathcal{S}\times\mathcal{S}$ of discriminative pairs, a set $\Theta$ of data distributions, and a set $\mathcal{M}$ of measurements. We expand on and explicitly define each component next.
Set $\mathcal{S}$ of potential secrets: Secrets are modeled as subsets of density operators that share a certain property (these subsets are merely singletons in the QDP case). The set $\mathcal{S}$ is a collection of such secret subsets. For example, if one aims to privatize the resource value $\mathsf{V}$ of a state, then the corresponding set of secrets is $\mathcal{S}=\bigcup_{i=1}^{n}\mathcal{T}_{i}$, where
$\mathcal{T}_{i}=\big{\{}\rho\in\mathcal{D}(\mathcal{H}):\mathsf{V}(\rho)=a_{i}\big{\}}$ (21)
and $\{a_{i}\}_{i=1}^{n}$ are the possible values that $\mathsf{V}$ can take (recall that, in Fig. 1, we considered a setup relevant to hiding the resource value $\mathsf{V}$ being $a$ or $b$).
Set $\mathcal{Q}$ of discriminative pairs: This is a subset of $\mathcal{S}\times\mathcal{S}$ that specifies which pairs of elements from $\mathcal{S}$ should be indistinguishable. Namely, if $(\mathcal{T}_{1},\mathcal{T}_{2})\in\mathcal{Q}$, then the goal of the privacy mechanism is to conceal whether the input belongs to $\mathcal{T}_{1}$ or $\mathcal{T}_{2}$. Here, note that $\rho\in\mathcal{T}_{1}\Rightarrow\rho\notin\mathcal{T}_{2}$. Note that $(\mathcal{T}_{i},\mathcal{T}_{j})\in\mathcal{Q}$ if and only if $(\mathcal{T}_{j},\mathcal{T}_{i})\in\mathcal{Q}$. Proceeding with the same example, we can set
$\mathcal{Q}=\bigcup_{i\neq j}\{(\mathcal{T}_{i},\mathcal{T}_{j})\}.$ (22)
Set $\Theta$ of data distributions: A collection of probability distributions $P_{X}\in\mathcal{P}(\mathcal{X})$ over a finite space $\mathcal{X}$ that indexes an ensemble of density operators $\{\rho^{x}\}_{x\in\mathcal{X}}$. Taking $X\sim P_{X}\in\Theta$, the matrix-valued random variable $\rho^{X}$ models a density operator that is randomly chosen according to $P_{X}$. The set $\Theta$ can be understood as capturing beliefs that the adversary has regarding the state of the system.
Set $\mathcal{M}$ of measurements: This set is a subset of all possible measurements, i.e., $\mathcal{M}\subseteq\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$. The choice of $\mathcal{M}$ gives the flexibility to consider only measurements that are possible under physical, legal, or ethical constraints.
Remark 1 (Designing QPP frameworks).
QPP allows system designers to explicitly encode their assumptions into the privacy framework. Setting the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ to accurately reflect the application of interest is crucial for obtaining meaningful privacy guarantees and to optimize utility. Explicit assumptions can also help account for ethical or fairness concerns associated with quantum systems; cf. Remark 17 for a concrete example concerning quantum fairness and how it is incorporated within QPP.
Now, we are ready to present a formal definition of the quantum analog of PP, which we call QPP.
Definition 4 (Quantum pufferfish privacy).
Fix $\varepsilon\geq 0$ and $\delta\in[0,1]$. A quantum algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-private in the quantum pufferfish framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ if for all $P_{X}\in\Theta$, $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ with $P_{X}(\mathcal{R}),P_{X}(\mathcal{T})>0$, and all $\mathrm{M}\in\mathcal{M}$, the following inequality holds:
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]\leq e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\delta,$ (23)
where
$\rho^{\mathcal{R}}$$\coloneqq\sum_{\{x:\rho^{x}\in\mathcal{R}\}}q_{\mathcal{R}}(x)\rho^{x},$ (24)
$q_{\mathcal{R}}(x)$ $\coloneqq\frac{P_{X}(x)}{P_{X}(\mathcal{R})},$ (25)
$P_{X}(\mathcal{R})$$\coloneqq\sum_{\{x:\rho^{x}\in\mathcal{R}\}}P_{X}(x),$ (26)
and $\rho^{\mathcal{T}}$ is defined similarly but with $\mathcal{T}$ instead of $\mathcal{R}$. We say that an algorithm ${\mathcal{A}}$ satisfies $\varepsilon$-QPP if it satisfies $(\varepsilon,0)$-QPP.
Evidently, discriminative secret pairs in $\mathcal{Q}$ are indistinguishable at the output of a QPP mechanism $\mathcal{A}$ in the $(\varepsilon,\delta)$-sense, under every measurement from the class $\mathcal{M}$.
Remark 2 (Semantics of the QPP framework).
Informally, the QPP framework provides the following privacy guarantee for fixed $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ and $P_{X}\in\Theta$: For a state $\rho^{X}$ chosen according to $X\sim P_{X}$ and input to the quantum channel $\mathcal{A}$, an adversary applying a measurement $\mathrm{M}\in\mathcal{M}$ on the channel output $\mathcal{A}(\rho^{X})$ draws the same conclusions regardless of whether $\rho^{X}$ belongs to $\mathcal{R}$ or $\mathcal{T}$.
Remark 3 (Incorporating entanglement).
We can incorporate entanglement in the QPP framework by introducing a reference system. Specifically, we can modify the QPP framework from $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ to $(\mathcal{S},\mathcal{G},\Theta,\mathcal{M}^{\prime})$, where
$\mathcal{G}\coloneqq\left\{\begin{array}[c]{c}(\omega_{RA}^{\mathcal{R}},\omega_{RA}^{\mathcal{T}}):\omega_{RA}^{\mathcal{R}},\omega_{RA}^{\mathcal{T}}\in\mathcal{D}(\mathcal{H}_{R}\otimes\mathcal{H}_{A}),\\ \mathrm{Tr}_{R}\!\left[\omega_{RA}^{\mathcal{R}}\right]=\rho^{\mathcal{R}},\ \mathrm{Tr}_{R}\!\left[\omega_{RA}^{\mathcal{T}}\right]=\rho^{\mathcal{T}},\\ (\mathcal{R},\mathcal{T})\in\mathcal{Q}\end{array}\right\}$ (27)
is a set of pairs of bipartite states. We then say that $\mathcal{A}$ is $(\varepsilon,\delta)$-QPP with respect to that framework if for all $P_{X}\in\Theta$, Msuperscriptnormal-$\mathrm{M}^{\prime}\in\mathcal{M}^{\prime}$, and $(\omega_{RA}^{\mathcal{R}},\omega_{RA}^{\mathcal{T}})\in\mathcal{G}$, we have
$\mathrm{Tr}\!\left[\mathrm{M}^{\prime}(\mathcal{I}\otimes\mathcal{A})(\omega_{RA}^{\mathcal{R}})\right]\leq e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}^{\prime}(\mathcal{I}\otimes\mathcal{A})(\omega_{RA}^{\mathcal{T}}))\right]+\delta,$ (28)
with $\rho^{\mathcal{R}}$ and $\rho^{\mathcal{T}}$ defined similar to Definition 4.
However, it is unclear whether such a stronger privacy notion would be useful in practical applications. For example, consider $\sigma_{1}\coloneqq|0\rangle\!\langle 0|\otimes\rho^{\mathcal{R}}$ and $\sigma_{2}\coloneqq|1\rangle\!\langle 1|\otimes\rho^{\mathcal{T}}$ with $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$. If a measurement on the reference system can be applied, then a computational-basis measurement distinguishes $\sigma_{1}$ and $\sigma_{2}$ perfectly. Thus, it is important to choose $\mathcal{A}$ appropriately with a practically applicable $\mathcal{M}^{\prime}$, such that the required indistinguishability is achieved. We shall revisit a variant of this framework with quantum divergences in Section IX-B.
III-B Equivalent Formulation of QPP with DL Divergence
We present an equivalent formulation for $(\varepsilon,\delta)$-QPP by means of the DL divergence from (10b). To the best of our knowledge, this provides the first operational interpretation of the DL divergence.
Proposition 1 (Equivalent formulation of $(\varepsilon,\delta)$-QPP).
Fix the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, with $\bar{\mathcal{M}}$ corresponding to the set of all possible measurements. Then algorithm $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-QPP with respect to the framework $(\mathcal{S},\mathcal{Q},\mathcal{M},\Theta)$ if and only if for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, we have
$\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\middle\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\varepsilon.$ (29)
Proof:
We first show that $(\varepsilon,\delta)$-QPP implies (29). For fixed $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, observe that $(\varepsilon,\delta)$-QPP corresponds to
$\sup_{\mathrm{M}\in\bar{\mathcal{M}}}\mathrm{Tr}\!\left[\mathrm{M}\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\varepsilon}\mathcal{A}(\rho^{\mathcal{T}})\right)\right]\leq\delta.$ (30)
Since
[…]
the inequality in (30) is equivalent to
$\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\varepsilon}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]\leq\delta.$ (32)
By the definition in (10b), this leads to $\varepsilon$ being a possible candidate for the optimization in $\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\middle\|\mathcal{A}(\rho^{\mathcal{T}})\right)$, and thus implies
$\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\varepsilon.$ (33)
As this holds for every $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, we obtain the desired implication $(\varepsilon,\delta)$-QPP $\Rightarrow$ (29).
Next, we show that (29) implies $(\varepsilon,\delta)$-QPP. Suppose that for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, we have $\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\varepsilon$. Then, for fixed $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, let
$\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\middle\|\mathcal{A}(\rho^{\mathcal{T}})\right)=\nu,$ (34)
which implies that $\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\nu}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]\leq\delta$. Recalling that $\nu\leq\varepsilon$ and noting that $\lambda\mapsto\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\lambda}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]$ is a monotonically decreasing function (cf. [14, Lemma 4.2]), we have
$\sup_{\mathrm{M}\in\bar{\mathcal{M}}}$ $\mathrm{Tr}\!\left[\mathrm{M}\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\varepsilon}\mathcal{A}(\rho^{\mathcal{T}})\right)\right]$
$\qquad=\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\varepsilon}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]$ (35)
$\qquad\leq\mathrm{Tr}\!\left[\left(\mathcal{A}(\rho^{\mathcal{R}})-e^{\nu}\mathcal{A}(\rho^{\mathcal{T}})\right)_{+}\right]$ (36)
$\qquad\leq\delta.$ (37)
As $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ are arbitrary, $(\varepsilon,\delta)$-QPP follows. ∎
Remark 4 (Operational interpretation of DL divergence).
For fixed $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, the DL divergence $\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)$ is equal to the minimal $\varepsilon$ that can be achieved for fixed $\delta$ via the indistinguishability condition of the QPP framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ stated in (23).
Remark 5 (Equivalent formulation with hockey-stick divergence).
Another equivalent formulation of QPP arises as a generalization of the information-theoretic equivalence for QDP [10]. Specifically, $\mathcal{A}$ is $(\varepsilon,\delta)$-QPP with respect to the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, where ¯=$\bar{\mathcal{M}}=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$, if
$\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\middle\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\delta,$ (38)
for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, where $\mathsf{E}_{\nu}(\rho\|\sigma)\coloneqq\mathrm{Tr}\!\left[(\rho-\nu\sigma)_{+}\right]$ is the hockey-stick divergence for $\nu\geq 1$ [28]. Fixing $P_{X}$ and $(\mathcal{R},\mathcal{T})$, the quantity $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)$ is the minimal $\delta$ that can be achieved for fixed $\varepsilon$ under the indistinguishability condition from (23).
III-C Reduction to Existing Privacy Frameworks
The proposed QPP framework subsumes other important privacy frameworks as special cases. These reductions are presented next.
III-C1 Quantum DP
In QDP (Definition 3), secrets are singletons, discriminative pairs comprise states satisfying a neighboring relation, while the measurement class $\mathcal{M}$ includes all possible measurements. QPP recovers the QDP setting by making the following choices:
$\mathcal{S}$ $=\mathcal{D},$ (39)
$\mathcal{Q}$ =$=\{(\rho,\sigma):\rho,\sigma\in\mathcal{D},\rho\sim\sigma\},$
$\Theta$ $=\mathcal{P}(\mathcal{X}),$
$\mathcal{M}$ =$=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}.$
More generally, one may add flexibility to the QDP formulation by considering other subsets $\Theta$. This can be used, for instance, to treat situations in which only certain neighboring pairs are of interest, namely, by choosing the distributions that assign positive probabilities only to those selected density operators. This can be interpreted as adding domain knowledge to the original QDP framework.
III-C2 Quantum Local DP
In quantum local DP (QLDP) [10], we choose secret pairs to be pairs of arbitrary distinct states, while the measurement class includes all possible measurements. Thus, QLDP realizes the same $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ framework as QDP, except that $\mathcal{Q}=\{(\rho,\sigma):\rho,\sigma\in\mathcal{D}\}$ for QLDP.
III-C3 Classical PP
Consider a classical PP framework $(\mathcal{S}_{c},\mathcal{Q}_{c},\Theta_{c})$, as specified in Definition 2. Let the encoding of the database $x\in\mathcal{X}^{n\times k}$ be $\rho^{x}\coloneqq|x\rangle\!\langle x|$, and denote a projective measurement operator corresponding to outcome $y$ as $|y\rangle\!\langle y|$. Here note that $\{|x\rangle\}_{x\in\mathcal{X}}$ and $\{|y\rangle\}_{y\in\mathcal{Y}}$ are respective orthonormal bases formed related to the input and output alphabets of the classical PP mechanism $\mathcal{A}_{c}$. Then classical PP is obtained from QPP by setting
$\mathcal{S}$ =$=\big{\{}\{\rho^{x}:x\in\mathcal{R}_{c}\}:\mathcal{R}_{c}\in\mathcal{S}_{c}\big{\}},$ (40)
$\mathcal{Q}$ ={($=\big{\{}\big{(}\{\rho^{x}:x\in\mathcal{R}_{c}\},\{\rho^{x}:x\in\mathcal{T}_{c}\}\big{)}:(\mathcal{R}_{c},\mathcal{T}_{c})\in\mathcal{Q}_{c}\}\big{\}},$
$\Theta$ $=\Theta_{c},$
$\mathcal{M}$ =$=\left\{\sum_{y\in\mathcal{B}}|y\rangle\!\langle y|:\mathcal{B}\subseteq\mathcal{Y}\right\}.$
In this scenario, assuming the output of the algorithm is discrete, we have that
$\mathcal{A}(\rho^{x})=\sum_{y,x^{\prime}}p(y|x)|y\rangle\!\langle x^{\prime}|\rho^{x}|x^{\prime}\rangle\!\langle y|$ (41)
where $p(y|x)=\mathbb{P}\big{(}\mathcal{A}_{c}(x)=y\big{)}$.
Remark 6 (Utility-optimized privacy models).
As is evident from above, the measurement set corresponding to classical PP entails every subset $\mathcal{B}\subseteq\mathcal{Y}$. However, when some of the outcomes are not sensitive, we may want to relax this requirement to gain utility (cf., e.g., [23]). While classical PP does not allow for that, QPP gives extra flexibility in choosing $\mathcal{M}$ and adapting it to the application of interest. Indeed, if we only need to privatize outcomes within the set $\mathcal{Y}^{\prime}\subsetneq\mathcal{Y}$, the smaller measurement set =$\mathcal{M}=\left\{\sum_{y\in\mathcal{Y}^{\prime}}|y\rangle\!\langle y|:\mathcal{B}\subseteq\mathcal{Y}^{\prime}\right\}$ is sufficient.
IV Datta–Leditzky Information Spectrum Divergence
We now focus on the DL divergence [14], whose operational interpretation in terms of QPP was provided in the previous section, and we study structural properties thereof, which will be useful when analyzing the QPP framework. We first formulate a primal and dual SDP to compute the DL divergence and then use that to prove joint-quasi convexity, the data-processing inequality under positive, trace-preserving maps, and connections to the smooth max-relative entropy.
IV-A SDP Formulations
We now present several SDPs for computing the DL divergence in (10b), which may be of independent interest. (Recall that the other DL divergence in (10a) is easily obtained by applying the equality in (12).)
Lemma 1 (SDP formulation of the DL divergence).
For $\delta\in(0,1)$, a state $\rho$, and a PSD operator $\sigma$, the following equalities hold
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $=\ln\inf_{\lambda,Z\geq 0}\left\{\lambda:\operatorname{Tr}[Z]\leq\delta,\ Z\geq\rho-\lambda\sigma\right\}$ (42a)
$=\ln\sup_{\mu,W\geq 0}\left\{\begin{array}[c]{c}\operatorname{Tr}[W\rho]-\mu\delta:\\ \operatorname{Tr}[W\sigma]\leq 1,\ W\leq\mu I\end{array}\right\}.$ (42d)
Proof:
Considering (13b), fix $\lambda>0$ and first observe that
[…]
Indeed, this follows because, for every $0\leq\Lambda\leq I$, we have that
$\operatorname{Tr}[\Lambda\left(\rho-\lambda\sigma\right)]$ $=\operatorname{Tr}\!\left[\Lambda\left(\left(\rho-\lambda\sigma\right)_{+}-\left(\rho-\lambda\sigma\right)_{-}\right)\right]$
$\leq\operatorname{Tr}\!\left[\Lambda\left(\rho-\lambda\sigma\right)_{+}\right]$
$\leq\operatorname{Tr}\!\left[\left(\rho-\lambda\sigma\right)_{+}\right],$ (44)
and the inequalities above are all attained by setting $\Lambda$ to be the projection onto the support of $\left(\rho-\lambda\sigma\right)_{+}$. The SDP dual of this quantity is given by
$\operatorname{Tr}\!\left[\left(\rho-\lambda\sigma\right)_{+}\right]=\inf_{Z\geq 0}\left\{\operatorname{Tr}[Z]:Z\geq\rho-\lambda\sigma\right\}.$ (45)
Intuitively, $Z=\left(\rho-\lambda\sigma\right)_{+}$ is the smallest choice of a PSD operator that satisfies the constraint $Z\geq\rho-\lambda\sigma$.
We then find from (10b), (13b), and (45) that
$\overline{\mathsf{D}}_{s}^{\delta}(\rho\|\sigma)$ $=\ln\inf\left\{\lambda\geq 0:\operatorname{Tr}[\left(\rho-\lambda\sigma\right)_{+}]\leq\delta\right\}$
$=\ln\inf_{\lambda,Z\geq 0}\left\{\lambda:\operatorname{Tr}[Z]\leq\delta,\ Z\geq\rho-\lambda\sigma\right\},$ (46)
which completes the proof of (42a).
The dual forms of these optimization problems are derived from the canonical primal and dual formulations of SDPs, which are respectively given by (cf. [27, Definition 2.20])
$\begin{split}&\inf_{Y\geq 0}\left\{\operatorname{Tr}[BY]:\Phi^{{\dagger}}(Y)\geq A\right\},\\ &\sup_{X\geq 0}\left\{\operatorname{Tr}[AX]:\Phi(X)\leq B\right\},\end{split}$ (47)
where $A$ and $B$ are Hermitian matrices and $\Phi$ is a Hermiticity-preserving superoperator. Comparing the former to (46), we make the following choices so that the general optimization problem recovers (46) (inside the logarithm):
$Y$ $=\begin{bmatrix}\lambda&0\\ 0&Z\end{bmatrix},\qquad B=\begin{bmatrix}1&0\\ 0&0\end{bmatrix},$ (52)
$\Phi^{{\dagger}}(Y)$ $=\begin{bmatrix}-\operatorname{Tr}[Z]&0\\ 0&Z+\lambda\sigma\end{bmatrix},$ (55)
$A$ $=\begin{bmatrix}-\delta&0\\ 0&\rho\end{bmatrix}.$ (58)
Then, setting
$X=\begin{bmatrix}\mu&0\\ 0&W\end{bmatrix},$ (59)
we solve for the map $\Phi(X)$ to find that
$\operatorname{Tr}[X\Phi^{{\dagger}}(Y)]$ $=\operatorname{Tr}\!\left[\begin{bmatrix}\mu&0\\ 0&W\end{bmatrix}\begin{bmatrix}-\operatorname{Tr}[Z]&0\\ 0&Z+\lambda\sigma\end{bmatrix}\right]$ (64)
$=-\mu\operatorname{Tr}[Z]+\operatorname{Tr}\!\left[W\left(Z+\lambda\sigma\right)\right]$ (65)
$=\operatorname{Tr}[\left(W-\mu I\right)Z]+\lambda\operatorname{Tr}[W\sigma]$ (66)
$=\operatorname{Tr}\!\left[\begin{bmatrix}\lambda&0\\ 0&Z\end{bmatrix}\begin{bmatrix}\operatorname{Tr}[W\sigma]&0\\ 0&W-\mu I\end{bmatrix}\right]$ (71)
$=\operatorname{Tr}[Y\Phi(X)],$ (72)
so that
$\Phi(X)=\begin{bmatrix}\operatorname{Tr}[W\sigma]&0\\ 0&W-\mu I\end{bmatrix}.$ (73)
Plugging into the dual form, we obtain
$\sup_{X\geq 0}\left\{\operatorname{Tr}[AX]:\Phi(X)\leq B\right\}\\ =\sup_{\mu,W\geq 0}\left\{\operatorname{Tr}[W\rho]-\mu\delta:\operatorname{Tr}[W\sigma]\leq 1,W\leq\mu I\right\}.$ (74)
Choose $\mu=\mu_{1}\in(0,1)$ and $W=\mu_{2}I$ such that $\mu_{1}\delta<\mu_{2}<\mu_{1}$, as a strictly feasible solution to the above. For the other SDP formulation from (46), set $\lambda$ such that $\mathrm{Tr}\!\left[(\rho-\lambda\sigma)_{+}\right]\leq\delta$, and $Z~{}=~{}(\rho-\lambda\sigma)_{+}\geq 0$ as a feasible solution. By Slater’s condition, we conclude that strong duality holds, and the primal and dual optimal values coincide. ∎
IV-B Properties
We derive several properties of the DL divergence from (13b), which are subsequently used in the analysis of the QPP framework. Basic properties of the DL divergence, including the data-processing inequality, have been proven in [14, Proposition 4.3]. Here, we generalize the data-processing inequality to hold for arbitrary positive, trace non-increasing maps (beyond the set of quantum channels) and also establish joint-quasi convexity of the DL divergence, along with its connection to the smooth max-relative entropy. The proofs of these properties rely on the SDP formulation from Lemma 1.
Proposition 2 (Properties of the DL divergence).
Fix $\delta\in(0,1)$, and let $\rho,\rho_{1},\ldots,\rho_{k}$ and $\sigma,\sigma_{1},\ldots,\sigma_{k}$ be two collections of states and PSD operators, respectively. The DL divergence in (13b) satisfies the following properties:
  1. 1.
    Data-processing inequality: For every positive, trace non-increasing map $\mathcal{N}$, we have
    $\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)\geq\overline{\mathsf{D}}^{\delta}\big{(}\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\big{)}.$ (75)
  2. 2.
    Joint-quasi convexity: Let $p_{i}\in[0,1]$, for $i\in\{1,\ldots,k\}$, with $\sum_{i=1}^{k}p_{i}=1$. Then
    $\overline{\mathsf{D}}^{\delta}\!\left(\sum_{i=1}^{k}p_{i}\rho_{i}\middle\|\sum_{i=1}^{k}p_{i}\sigma_{i}\right)\leq\max_{i}\overline{\mathsf{D}}^{\delta}\!\left(\rho_{i}\|\sigma_{i}\right),$ (76)
    and, more generally,
    $\overline{\mathsf{D}}^{\delta^{\prime}}\!\left(\sum_{i=1}^{k}p_{i}\rho_{i}\middle\|\sum_{i=1}^{k}p_{i}\sigma_{i}\right)\leq\max_{i}\overline{\mathsf{D}}^{\delta_{i}}\!\left(\rho_{i}\|\sigma_{i}\right),$ (77)
    where $\delta^{\prime}\coloneqq\sum_{i=1}^{k}p_{i}\delta_{i}$ with $\delta_{1},\ldots,\delta_{k}\in(0,1)$.
  3. 3.
    Relation to smooth max-relative entropy:
    $\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)\leq\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)\leq\overline{\mathsf{D}}^{\delta^{\prime}}\!(\rho\|\sigma)-\ln\!\left(1-\delta^{\prime}\right),$ (78)
    where $\delta^{\prime}\coloneqq 1-\sqrt{1-\delta^{2}}\in(0,1)$, and the second inequality above can be equivalently written as
    $\mathsf{D}_{\max}^{\sqrt{\delta(2-\delta)}}(\rho\|\sigma)\leq\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)-\ln\!\left(1-\delta\right).$ (79)
  4. 4.
    Quasi subadditivity: Let $\delta_{1},\delta_{2}\in(0,1)$ satisfy $\delta_{1}^{\prime}+\delta_{2}^{\prime}<1$, with $\delta^{\prime}_{i}\coloneqq\sqrt{\delta_{i}(2-\delta_{i})}\in(0,1)$ for $i\in\{1,2\}$. Then
    $\overline{\mathsf{D}}^{\delta_{1}^{\prime}+\delta_{2}^{\prime}}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})\\ \leq\overline{\mathsf{D}}^{\delta_{1}}(\rho_{1}\|\sigma_{1})+\overline{\mathsf{D}}^{\delta_{2}}(\rho_{2}\|\sigma_{2})-\ln\big{(}(1-\delta_{1})(1-\delta_{2})\big{)}.$ (80)
    Furthermore,
    1. (a)
      if $\delta_{1}=\delta_{2}=0$, then
      $\overline{\mathsf{D}}^{0}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})\leq\overline{\mathsf{D}}^{0}(\rho_{1}\|\sigma_{1})+\overline{\mathsf{D}}^{0}(\rho_{2}\|\sigma_{2}).$ (81)
    2. (b)
      if $\sigma_{1},\sigma_{2}$ are states, then
      $\overline{\mathsf{D}}^{\delta}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})\leq\overline{\mathsf{D}}^{\delta_{1}}(\rho_{1}\|\sigma_{1})+\overline{\mathsf{D}}^{\delta_{2}}(\rho_{2}\|\sigma_{2}),$ (82)
      where
      $\delta\coloneqq\min\!\left\{\delta_{1}+e^{\overline{\mathsf{D}}^{\delta_{1}}(\rho_{1}\|\sigma_{1})}\delta_{2},\ \delta_{2}+e^{\overline{\mathsf{D}}^{\delta_{2}}(\rho_{2}\|\sigma_{2})}\delta_{1}\right\}.$ (83)
Proof:
Property 1: The statement was proven in [14, Proposition 4.3] for a quantum channel $\mathcal{N}$ by using the inequality
$\mathrm{Tr}\!\left[(\rho-e^{\lambda}\sigma)_{+}\right]\geq\mathrm{Tr}\!\left[\!\left(\mathcal{N}(\rho)-e^{\lambda}\mathcal{N}(\sigma)\right)_{+}\right],$ (84)
which holds for all $\lambda\in\mathbb{R}$. Here, we prove the data-processing inequality, but we generalize it to hold for a positive, trace non-increasing map $\mathcal{N}$. Our derivation relies on the SDP formulation of the DL divergence from (42a).
Let $\lambda^{\star}$ and $Z^{\star}$ be optimal choices [^2] When the DL divergence is finite, the infimum is achieved by a standard continuity plus compactness argument. The stated relations trivially hold when the DL divergence is infinite. in the optimization for $\overline{\mathsf{D}}^{\delta}\!\left(\rho\|\sigma\right)$, so that $\overline{\mathsf{D}}^{\delta}\!\left(\rho\|\sigma\right)=\ln\lambda^{\star}$, $Z^{\star}\geq\rho-\lambda^{\star}\sigma$ with $\mathrm{Tr}\!\left[Z^{\star}\right]\leq\delta$, and $Z^{\star}\geq 0$ (indeed, note that the infimum is achieved with $\mathrm{Tr}[(\rho-\lambda^{\star}\sigma)_{+}]=\delta$). Since $Z^{\star}-(\rho-\lambda^{\star}\sigma)\geq 0$, it follows that $\mathcal{N}\!\big{(}Z^{\star}-(\rho-\lambda^{\star}\sigma)\big{)}\geq 0$ from the assumption that $\mathcal{N}$ is a positive map. Consequently, we obtain
$Z^{\prime}\coloneqq\mathcal{N}(Z^{\star})\geq\mathcal{N}(\rho)-\lambda^{\star}\mathcal{N}(\sigma).$ (85)
Furthermore, $Z^{\prime}\geq 0$ since $Z^{\star}\geq 0$ and $\mathcal{N}$ is a positive map. Additionally, since $\mathcal{N}$ is trace non-increasing, it follows that
$\operatorname{Tr}[Z^{\prime}]\leq\operatorname{Tr}[Z^{\star}]\leq\delta.$ (86)
Thus, $\lambda^{\star}$ is a feasible point for $\overline{\mathsf{D}}^{\delta}\!\big{(}\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\big{)}$. We conclude the proof by noting that the quantity $\overline{\mathsf{D}}^{\delta}\big{(}\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\big{)}$ involves a minimization over all such feasible points, implying the desired inequality:
$\overline{\mathsf{D}}^{\delta}\big{(}\mathcal{N}(\rho)\|\mathcal{N}(\sigma)\big{)}\leq\ln(\lambda^{\star})=\overline{\mathsf{D}}^{\delta}\!\left(\rho\|\sigma\right).$ (87)
Property 2: We again consider the SDP from (42a). Let $\lambda^{\star}_{i}$ and $Z^{\star}_{i}$ be optimal for $\overline{\mathsf{D}}^{\delta}\!\left(\rho_{i}\|\sigma_{i}\right)$, so that $\overline{\mathsf{D}}^{\delta}\!\left(\rho_{i}\|\sigma_{i}\right)=\ln(\lambda^{\star}_{i})$, $Z^{\star}_{i}\geq\rho_{i}-\lambda^{\star}_{i}\sigma_{i}$ with $\operatorname{Tr}\!\left[Z^{\star}_{i}\right]\leq\delta$, and $Z^{\star}_{i}\geq 0$. Define
$Z\coloneqq\sum_{i=1}^{k}p_{i}Z^{\star}_{i}\geq\sum_{i=1}^{k}p_{i}\rho_{i}-\sum_{i=1}^{k}\lambda^{\star}_{i}p_{i}\sigma_{i}.$ (88)
and observe that $\mathrm{Tr}\!\left[Z\right]\leq\delta$ and $Z\geq 0$. This implies that $Z\geq\sum_{i=1}^{k}p_{i}\rho_{i}-\max_{i}\lambda^{\star}_{i}\sum_{j=1}^{k}p_{j}\sigma_{j}$, which suggests that $\max_{i}\lambda^{\star}_{i}$ and $Z$ are (candidate) infimizers in the SDP formulation of $\overline{\mathsf{D}}^{\delta}\!\left(\sum_{i=1}^{k}p_{i}\rho_{i}\middle\|\sum_{i=1}^{k}p_{i}\sigma_{i}\right)$. Consequently, we obtain
$\overline{\mathsf{D}}^{\delta}\!\left(\sum_{i=1}^{k}p_{i}\rho_{i}\middle\|\sum_{i=1}^{k}p_{i}\sigma_{i}\right)\leq\ln\!\left(\max_{i}\lambda^{\star}_{i}\right)=\max_{i}\overline{\mathsf{D}}^{\delta}\!\left(\rho_{i}\|\sigma_{i}\right).$ (89)
The proof of the general case follows along the same lines by observing that $\mathrm{Tr}[Z]\leq\sum_{i=1}^{k}p_{i}\delta_{i}$.
Property 3: From [38, Appendix B], we have that
${\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)}$ $=\ln\inf_{\lambda,\widetilde{\rho},Y\geq 0}\left\{\begin{array}[c]{c}\lambda:\widetilde{\rho}\leq\lambda\sigma,\ \ \operatorname{Tr}[Y]\leq\delta,\\ \operatorname{Tr}[\widetilde{\rho}]=1,\ \ Y\geq\rho-\widetilde{\rho}\end{array}\right\}.$ (92)
Let $\lambda$, $Y$, and $\widetilde{\rho}$ be arbitrary operators satisfying the constraints for ${\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)}$. Then by combining the inequalities $\widetilde{\rho}\leq\lambda\sigma$ and $Y\geq\rho-\widetilde{\rho}$, we get
$Y\geq\rho-\lambda\sigma.$ (93)
We see that $\lambda$ and $Y$ satisfy the constraints needed for $\lambda$ and $Z$, respectively, in the SDP for $\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$, whereby
${\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)}\leq\lambda.$ (94)
Since the argument holds for all $\lambda$, $Y$, and $\widetilde{\rho}$ satisfying the constraints in the definition of ${\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)}$, we further obtain
${\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)}\leq{\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)}.$ (95)
The proof is concluded by invoking the following lemma (proven in Appendix A).
Lemma 2.
Fix $\lambda>0$, let $\rho$ be a state and $\sigma$ a positive semi-definite operator, and define $\delta\coloneqq\operatorname{Tr}[\left(\rho-\lambda\sigma\right)_{+}]$. Then
$\mathsf{D}_{\max}^{\sqrt{\delta\left(2-\delta\right)}}(\rho\|\sigma)\leq\ln\lambda-\ln(1-\delta).$ (96)
For fixed $\delta\in(0,1)$, by definition, we have $\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)=\ln(\lambda^{\star})$ with $\delta=\operatorname{Tr}[\left(\rho-\lambda^{\star}\sigma\right)_{+}]$. With that, Lemma 2 with the reparametrization $\delta\to 1-\sqrt{1-\delta^{2}}$, yields
$\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)\leq\overline{\mathsf{D}}^{1-\sqrt{1-\delta^{2}}}(\rho\|\sigma)+\ln\!\left(\frac{1}{\sqrt{1-\delta^{2}}}\right).$ (97)
This completes the proof.
Property 4: This follows by invoking Property 3 and using the fact that the smooth max-relative entropy satisfies subadditivity (Appendix B) with
$\mathsf{D}_{\max}^{\delta_{1}+\delta_{2}}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})\leq\mathsf{D}_{\max}^{\delta_{1}}(\rho_{1}\|\sigma_{1})+\mathsf{D}_{\max}^{\delta_{2}}(\rho_{2}\|\sigma_{2}).$ (98)
Part (a) now follows by taking the limits $\delta_{1}\to 0$ and $\delta_{2}\to 0$ in (98), and applying Property 3.
To prove Part (b), we use the SDP formulation in (42a). Let $\overline{\mathsf{D}}^{\delta_{i}}\!\left(\rho_{i}\|\sigma_{i}\right)=\ln(\lambda^{\star}_{i})$ for $i\in\{1,2\}$. It follows that $Z_{i}\geq\rho_{i}-\lambda^{\star}_{i}\sigma_{i}$ with $\mathrm{Tr}\!\left[Z_{i}\right]\leq\delta$ and $Z_{i}\geq 0$. Consider that
$(\rho_{1}\otimes\rho_{2})-\lambda^{\star}_{1}\lambda^{\star}_{2}(\sigma_{1}\otimes\sigma_{2})$
$=(\rho_{1}\otimes\rho_{2})-\lambda^{\star}_{1}\sigma_{1}\otimes\rho_{2}+\lambda^{\star}_{1}\sigma_{1}\otimes\rho_{2}-\lambda^{\star}_{1}\lambda^{\star}_{2}(\sigma_{1}\otimes\sigma_{2})$
$=(\rho_{1}-\lambda^{\star}_{1}\sigma_{1})\otimes\rho_{2}+\lambda^{\star}_{1}\sigma_{1}\otimes(\rho_{2}-\lambda^{\star}_{2}\sigma_{2})$
$\leq Z_{1}\otimes\rho_{2}+\lambda^{\star}_{1}\sigma_{1}\otimes Z_{2}\eqqcolon Z.$ (99)
Observe that $Z\geq 0$ and $\mathrm{Tr}\!\left[Z\right]=\mathrm{Tr}\!\left[Z_{1}\right]+\lambda^{\star}_{1}\mathrm{Tr}\!\left[Z_{2}\right]$, since $\mathrm{Tr}\!\left[\rho_{1}\right]=\mathrm{Tr}\!\left[\sigma_{1}\right]=1$. Consequently, we have $\mathrm{Tr}\!\left[Z\right]\leq\delta_{1}+\lambda^{\star}_{1}\delta_{2}$, and $\lambda^{\star}_{1}\lambda^{\star}_{2}$ is a candidate infimizer. For $\delta^{\prime}=\delta_{1}+\lambda^{\star}_{1}\delta_{2}$, we now arrive at
$\overline{\mathsf{D}}^{\delta^{\prime}}\!\left(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2}\right)$ $\leq\ln(\lambda^{\star}_{1}\lambda^{\star}_{2})$ (100)
$=\ln(\lambda^{\star}_{1})+\ln(\lambda^{\star}_{2})$ (101)
$=\overline{\mathsf{D}}^{\delta_{1}}\!\left(\rho_{1}\|\sigma_{1}\right)+\overline{\mathsf{D}}^{\delta_{2}}\!\left(\rho_{2}\|\sigma_{2}\right).$ (102)
The above holds for $\delta^{\prime}=\delta_{2}+\lambda^{\star}_{2}\delta_{1}$ as well, by adding and subtracting $\lambda_{2}^{\star}\rho_{1}\otimes\sigma_{1}$ instead of $\lambda_{1}^{\star}\sigma_{1}\otimes\rho_{2}$, and then following the same argument. ∎
V Properties and Mechanisms for QPP
V-A Properties of QPP Mechanisms
Modern guidelines for privacy frameworks [43] render properties such as convexity and post-processing (also known as transformation invariance) as basic requirements for privacy frameworks. Composability is another important property, which implies that a combination of privacy mechanisms is itself private. These properties are known to hold for the classical mutual information PP framework, and all of them, except for composability, hold for the classical PP framework; cf. [7, Theorem 2] and [3, Theorem 5.1], respectively.
Before proving these properties for the QPP framework, we discuss their operational interpretation. Convexity means that applying a QPP mechanism that is randomly chosen from a given set of such mechanisms still satisfies QPP. Post-processing ensures that passing the output of a QPP mechanism $\mathcal{A}$ through a channel $\mathcal{N}$ preserves QPP; see Fig. 1(a). Parallel composability is illustrated in Fig. 1(b) and guarantees that QPP holds after applying
$\mathcal{A}^{(k)}\coloneqq\bigotimes_{i=1}^{k}\mathcal{A}_{i}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}\otimes\cdots\otimes\mathcal{A}_{k}$ (103)
to the input $\rho^{X_{1}}\otimes\rho^{X_{2}}\otimes\cdots\otimes\rho^{X_{k}}$, with $X_{i}\sim P_{X}\in\Theta$, where each $X_{i}$ is independently chosen. Informally, the semantic meaning of this property is that after applying $\mathcal{A}^{(k)}$, the same conclusions can be drawn about the input $\rho^{X_{1}}\otimes\rho^{X_{2}}\otimes\cdots\otimes\rho^{X_{k}}$ regardless of whether each $\rho^{X_{i}}$ belongs to $\mathcal{R}_{i}$ or $\mathcal{T}_{i}$, where $(\mathcal{R}_{i},\mathcal{T}_{i})\in\mathcal{Q}$ for all $i\in\{1,\ldots,k\}$. In this setting, the set of discriminative pairs is taken as
$\mathcal{Q}^{(k)}\coloneqq\left\{(\mathcal{R}^{(k)},\mathcal{T}^{(k)}):\begin{aligned} &\mathcal{R}^{(k)}\coloneqq(\mathcal{R}_{1},\ldots,\mathcal{R}_{k})\\ &\mathcal{T}^{(k)}\coloneqq(\mathcal{T}_{1},\ldots,\mathcal{T}_{k})\\ &\forall i\in\{1,\ldots,k\}\\ &(\mathcal{R}_{i},\mathcal{T}_{i})\in\mathcal{Q}\end{aligned}\right\}.$ (104)
Furthermore, the class of product measurements is $\bigotimes_{i=1}^{k}\mathcal{M}_{i}$ (i.e., the output of algorithm $\mathcal{A}_{i}$ is followed by a measurement from $\mathcal{M}_{i}$, for all $i\in\{1,\ldots,k\}$), while the set of all possible measurements on the $k$ systems, including joint measurements, is denoted by ¯$\bar{\mathcal{M}}^{k}$. We note here that one could consider other classes of limited measurements, such as local operations and classical communication (LOCC) measurements and positive-partial-transpose (PPT) measurements [44].
Figure 2: Properties of QPP mechanisms: (a) refers to post-processing of QPP algorithm $\mathcal{A}$; If $\mathcal{A}$ satisfies QPP, then $\mathcal{N}\circ\mathcal{A}$ also satisfies QPP. (b) refers to parallel composition of $k$ QPP mechanisms; composition of $k$ mechanisms independently in a parallel fashion satisfies QPP if each $\mathcal{A}_{i}$ satisfies QPP.
The formal statement of these properties is as follows.
Theorem 1 (Properties of QPP mechanisms).
The following properties hold:
  1. 1.
    Convexity: Let $\mathcal{A}_{1},\ldots,\mathcal{A}_{k}$ be $(\varepsilon,\delta)$-QPP mechanisms in the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$. Take $I$ to be a $k$-ary categorical random variable with probability distribution $(p_{1},\ldots,p_{k})$. Then the mechanism $\mathcal{A}\coloneqq\mathcal{A}_{I}$ (i.e., $\mathcal{A}=\mathcal{A}_{i}$ with probability $p_{i}$, for $i\in\{1,\ldots,k\}$) also satisfies $(\varepsilon,\delta)$-QPP in the same framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$.
  2. 2.
    Post-processing: If a mechanism $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$, then, for a quantum channel $\mathcal{N}$, the processed mechanism $\mathcal{N}\circ\mathcal{A}$ also satisfies $(\varepsilon,\delta)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M}^{\prime})$, where $\mathcal{M}^{\prime}\subseteq\left\{\mathrm{M}^{\prime}:\mathcal{N}^{\dagger}(\mathrm{M}^{\prime})\in\mathcal{M}\right\}$.
  3. 3.
    Parallel composability (non-adaptive): Let $\mathcal{A}_{1},\ldots,\mathcal{A}_{k}$ be mechanisms such that $\mathcal{A}_{i}$ is $(\varepsilon_{i},\delta_{i})$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M}_{i})$, for each $i\in\{1,\ldots,k\}$. Then the composed mechanism
    $\mathcal{A}^{(k)}:\bigotimes_{i=1}^{k}\sigma_{i}\mapsto\mathcal{A}_{1}(\sigma_{1})\otimes\cdots\otimes\mathcal{A}_{k}(\sigma_{k})$ (105)
    satisfies $\left(\sum_{i=1}^{k}\varepsilon_{i},\sum_{i=1}^{k}\delta_{i}\right)$-QPP in the framework $\left(\mathcal{S},\mathcal{Q}^{(k)},\Theta,\bigotimes_{i=1}^{k}\mathcal{M}_{i}\right)$.
Proof:
See Appendix C. ∎
The latter two properties above change if one considers a measurement class that comprises all possible measurements $\mathcal{M}^{k}$, as opposed to only product measurements. This is one of the main distinctions between the semi-classical and quantum cases, where, for the latter, joint measurements can infer more information and thus privacy degrades. The following theorem accounts for this latter scenario.
Theorem 2 (Properties of QPP with $\mathcal{M}=\bar{\mathcal{M}}$).
The following properties hold for the case in which the measurement class is $\bar{\mathcal{M}}$:
  1. 1.
    Convexity: Let $\mathcal{A}_{1},\ldots,\mathcal{A}_{k}$ be $(\varepsilon,\delta)$-QPP mechanisms in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$. Take $I$ to be a $k$-ary categorical random variable with parameters $(p_{1},\ldots,p_{k})$. Then the mechanism $\mathcal{A}\coloneqq\mathcal{A}_{I}$ (i.e., $\mathcal{A}=\mathcal{A}_{i}$ with probability $p_{i}$, for $i\in\{1,\ldots,k\}$) also satisfies $(\varepsilon,\delta)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$.
  2. 2.
    Post-processing: If a mechanism $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, then, for a quantum channel $\mathcal{N}$, the mechanism $\mathcal{N}\circ\mathcal{A}$ also satisfies $(\varepsilon,\delta)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$.
  3. 3.
    Parallel composability: If $\mathcal{A}_{i}$ satisfies $(\varepsilon_{i},\delta_{i})$-QPP in $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ for $i\in\{1,2\}$, then the composed mechanism $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ satisfies $(\varepsilon^{\prime},\delta^{\prime})$-QPP in the framework $\left(\mathcal{S},\mathcal{Q}^{(2)},\Theta,\bar{\mathcal{M}}^{2}\right)$ where
    $\varepsilon^{\prime}$ $\coloneqq\varepsilon_{1}+\varepsilon_{2}+\ln\!\left(\frac{1}{(1-\delta_{1})(1-\delta_{2})}\right),$ (106)
    $\delta^{\prime}$ $\coloneqq\sqrt{\delta_{1}(2-\delta_{1})}+\sqrt{\delta_{2}(2-\delta_{2})}.$ (107)
    and $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ also satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta)$-QPP where
    $\delta\coloneqq\min\{\delta_{1}+e^{\varepsilon_{1}}\delta_{2},\delta_{2}+e^{\varepsilon_{2}}\delta_{1}\}.$ (108)
    Observe that, if $\delta_{i}=0$ for $i\in\{1,2\}$, then $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ satisfies $(\varepsilon_{1}+\varepsilon_{2})$-QPP for the parallel composed framework.
Proof:
The proof of Theorem 2 relies on properties of the DL divergence established in Proposition 2. Items 1), 2), and 3) follow from joint quasi-convexity (Property 2), data processing (Property 1), and quasi subadditivity (Property 4), respectively. ∎
Remark 7 (Comparison to existing results).
In [10, Corollary III.3], the parallel composition of two mechanisms that satisfy $(\varepsilon_{i},\delta_{i})$-QDP for $i\in\{1,2\}$ is shown to be $(\varepsilon_{1}+\varepsilon_{2},\delta)$-QDP, where $\delta$ is given in (108). The proof technique is, however, different from ours. Property 3 of Theorem 1 also reveals that if one considers a restricted class of measurements (e.g., product measurements), then it is possible to achieve tighter privacy guarantees (namely, $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-QDP) than those obtained when allowing all joint measurements on the two systems. Also note that $\delta^{\prime}$ in (107) is independent of $\varepsilon_{1}$ and $\varepsilon_{2}$, whereas $\delta$ in (108) depends on them. Depending on the particular values that the parameters $\varepsilon_{i}$ and $\delta_{i}$ take, for $i\in\{1,2\}$, one of these aforementioned results provides sharper privacy guarantees.
Example 1.
Here we provide an example to illustrate the distinction between the case of joint measurements and product measurements, and more generally PPT measurements (which contain the set of LOCC measurements, as well as product measurements). Let $\alpha_{d}$ be the maximally mixed state on the antisymmetric subspace of two $d$-dimensional systems, and let $\sigma_{d}$ be the maximally mixed state on the symmetric subspace, i.e.,
$\alpha_{d}$ $\coloneqq\frac{I-F}{d\left(d-1\right)},$ (109)
$\sigma_{d}$ $\coloneqq\frac{I+F}{d\left(d+1\right)},$ (110)
where $F\coloneqq\sum_{i,j}|i\rangle\!\langle j|\otimes|j\rangle\!\langle i|$ is the unitary swap operator. These states are orthogonal and thus perfectly distinguishable by a joint measurement. Indeed, this measurement is given by $\left\{\Pi^{\alpha_{d}},\Pi^{\sigma_{d}}\right\}$, where $\Pi^{\alpha_{d}}\coloneqq\left(I-F\right)/2$ and $\Pi^{\sigma_{d}}\coloneqq\left(I+F\right)/2$. By setting $\mathrm{M}=\Pi^{\alpha_{d}}$, we find that $\operatorname{Tr}[\mathrm{M}\alpha_{d}]=1$ and $\operatorname{Tr}[\mathrm{M}\sigma_{d}]=0$.
[…]
However, we can alternatively restrict the measurement operators to PPT measurement operators, i.e., those $\mathrm{M}$ which satisfy $0\leq\mathrm{M}\leq I$ and $0\leq\mathrm{M}^{\Gamma}\leq I$, where the $\Gamma$ superscript denotes the partial transpose. In this case, we find for every such PPT measurement operator $M$ that
$\operatorname{Tr}[\mathrm{M}\alpha_{d}]$ $=\operatorname{Tr}[\mathrm{M}^{\Gamma}\alpha_{d}^{\Gamma}]$ (111)
$=\operatorname{Tr}[\mathrm{M}^{\Gamma}\left(I-d\,\Phi^{d}\right)/\left(d\left(d-1\right)\right)]$ (112)
$\leq\operatorname{Tr}[\mathrm{M}^{\Gamma}\left(I+d\,\Phi^{d}\right)/\left(d\left(d-1\right)\right)]$ (113)
$=\operatorname{Tr}[\mathrm{M}\left(I+F\right)/\left(d\left(d-1\right)\right)]$ (114)
$=\frac{d+1}{d-1}\operatorname{Tr}[\mathrm{M}\sigma_{d}].$ (115)
The first equality follows because the partial transpose is its own adjoint, and the second equality follows by introducing the maximally entangled state $\Phi^{d}\coloneqq\frac{1}{d}\sum_{i,j}|i\rangle\!\langle j|\otimes|i\rangle\!\langle j|$. The inequality follows because $0\leq\mathrm{M}^{\Gamma}$. By applying the above inequality, we conclude that $\left(\varepsilon,\delta\right)$-QPP holds with $\varepsilon=\ln\!\left(\frac{d+1}{d-1}\right)$ and $\delta=0$, so that privacy improves as dimension increases.
V-A1 Adaptive Composability
Adaptive composition refers to the case when each subsequently composed mechanism is chosen based on the outputs of the preceding ones. The goal is to quantify the overall privacy leakage at the output of the adaptively composed mechanism. This ideas has been studied in detail for classical privacy settings [2], and here we explore it for QPP.
Figure 3: Setup for adaptive composition: On the top system, the channel $\mathcal{A}_{1}$ is followed by the quantum instrument $\{\mathcal{E}_{y}\}_{y\in\mathcal{Y}}$, and then the random classical outcome $Y$ is used to choose the channel $\mathcal{A}_{2}^{Y}$. In this setting, we analyse how well an adversary can learn properties of the input state $\sigma_{I}$ by applying measurements on the output state.
We first focus on the setting depicted in Fig. 3. Fix $X_{i}\sim P_{X}\in\Theta$ for $i\in\{1,2\}$, which are independently chosen, and let the input state be
$\sigma_{I}\coloneqq\rho^{X_{1}}\otimes\rho^{X_{2}}.$ (116)
On the top subsystem in Fig. 3, the channel $\mathcal{A}_{1}$ is followed by the quantum instrument $\{\mathcal{E}_{y}\}_{y\in\mathcal{Y}}$, which is a collection of completely positive maps such that the sum map
¯$\overline{\mathcal{E}}\coloneqq\sum_{y\in\mathcal{Y}}\mathcal{E}_{y}$ (117)
is trace preserving [45, 46, 47]. Depending on the measurement outcome $y$, the channel $\mathcal{A}_{2}^{y}$ is chosen and applied to the bottom subsystem. The combined output state at stage $O$, as marked in the figure, is
$\sigma_{O}$ $\coloneqq\sum_{y\in\mathcal{Y}}\mathcal{E}^{y}\!\left(\mathcal{A}_{1}(\rho^{X_{1}})\right)\otimes|y\rangle\!\langle y|\otimes\mathcal{A}^{y}_{2}(\rho^{X_{2}}).$ (118)
We focus on adaptive composition of two quantum mechanisms in the above described setting. Suppose that $\mathcal{A}_{1}$ is an $(\varepsilon_{1},\delta_{1})$-QPP mechanism in the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M}_{1})$. Suppose furthermore that, for each outcome $y\in\mathcal{Y}$, the mechanism $\mathcal{A}_{2}^{y}$ satisfies $(\varepsilon_{2},\delta_{2})$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M}_{2})$ in the following sense: for all $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ and $\mathrm{M}\in\mathcal{M}_{2}$,
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}_{2}^{y}(\rho^{\mathcal{R}})\right]\leq e^{\varepsilon_{2}}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}})\right]+\delta_{2}.$ (119)
Under adaptive composition, we want to guarantee indistinguishability of pairs of states
$\sigma_{I}^{\mathcal{R}}\coloneqq\rho^{\mathcal{R}_{1}}\otimes\rho^{\mathcal{R}_{2}}\ \textnormal{ and }\ \sigma_{I}^{\mathcal{T}}\coloneqq\rho^{\mathcal{T}_{1}}\otimes\rho^{\mathcal{T}_{2}},$ (120)
where $(\mathcal{R}_{i},\mathcal{T}_{i})\in\mathcal{Q}$ for $i\in\{1,2\}$. This means, informally, that the adversary would draw the same conclusions regardless of whether $\rho^{X_{i}}$ belongs to $\mathcal{R}_{i}$ or $\mathcal{T}_{i}$, for $i\in\{1,2\}$, when the initial input $\sigma_{I}$ to the system in Fig. 3 is given by (116). The following proposition provides parameters under which QPP of the adaptively composed mechanism is guaranteed.
Proposition 3 (Adaptive composition of QPP).
Fix the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$. Suppose that $\mathcal{A}_{1}$ satisfies $(\varepsilon_{1},\delta_{1})$-QPP and $\mathcal{A}_{2}^{y}$ satisfies $(\varepsilon_{2},\delta_{2})$-QPP for every measurement outcome $y$, as in (119). Then the mechanism in Fig. 3 satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta_{2}+\delta_{1}|\mathcal{Y}|)$-QPP in the framework $(\mathcal{S},\mathcal{Q}\times\mathcal{Q},\Theta,\bar{\mathcal{M}}\otimes\bar{\mathcal{M}})$ where $|\mathcal{Y}|$ denotes the cardinality of the set $\mathcal{Y}$.
Proof:
See Appendix D. ∎
Note that, when $\delta_{i}=0$ for $i\in\{1,2\}$, the privacy parameters are additive. However, when $\delta_{i}\neq 0$, the privacy parameter $\delta_{2}+\delta_{1}|\mathcal{Y}|$ degrades linearly with an increasing number of measurement outcomes.
V-A2 Composability with Correlated States
In Property 3 of Theorem 1 and Proposition 3, we considered the case in which two mechanisms, composed either in parallel or adaptively, receive independent inputs (i.e., the input being $\rho^{X_{1}}\otimes\rho^{X_{2}}$ where $X_{i}\sim P_{X}\in\Theta$ for $i=\{1,2\}$, which are chosen independently). We now focus on the setting in which the inputs are correlated. The input is chosen as a separable state of the form
$\sigma_{I}\coloneqq\sum_{z\in\mathcal{Z}}q(z)\,\omega^{z}\otimes\tau^{z},$ (121)
where $q$ represents a probability distribution with $q(z)\geq 0$ and $\sum_{z\in\mathcal{Z}}q(z)=1$, and $\omega^{z}$ and $\tau^{z}$ are quantum states for all $z\in\mathcal{Z}$ [^3] Note that (121) covers the case of having input states of the form $\sigma_{I}\coloneqq\sum_{(x,y)\in\mathcal{X}\times\mathcal{Y}}q(x,y)\ \omega^{x}\otimes\tau^{y}$ where for all $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ $\omega^{x}$ and $\tau^{y}$ are states, by considering $z$ to be an index for multiple variables, i.e., setting $z=(x,y)$. One special case of interest is as follows:
$\sigma_{I}\coloneqq\sum_{x\in\mathcal{X}}P_{X}(x)\,\rho^{x}\otimes\rho^{x}.$ (122)
In this setting, QDP ensures indistinguishability of the input states
$\sigma^{1}_{I}\coloneqq\sum_{z\in\mathcal{Z}}q(z)\,\omega_{1}^{z}\otimes\tau_{1}^{z}\ \textnormal{and}\ \sigma^{2}_{I}\coloneqq\sum_{z\in\mathcal{Z}}q(z)\,\omega_{2}^{z}\otimes\tau_{2}^{z},$ (123)
where $\omega_{1}^{z}\sim\omega_{2}^{z}$ and $\tau_{1}^{z}\sim\tau_{2}^{z}$ are neighbors for all $z\in\mathcal{Z}$.
We consider an instance of the QPP framework, called flexible QDP, where $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ is such that $\Theta$ and $\mathcal{M}$ are chosen based on user needs, while the other parameters are as given in Section III-C1. Flexible QDP then satisfies the following composability properties.
Corollary 1 (Composability of flexible QDP).
Let the initial input to the two mechanisms $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be of the form $\sum_{z\in\mathcal{Z}}q(z)\,\omega^{z}\otimes\tau^{z}$. The following composability properties hold for the QDP framework.
Parallel composability: Consider the parallel composed mechanism $\sum_{z\in\mathcal{Z}}q(z)\,\mathcal{A}_{1}(\omega^{z})\otimes\mathcal{A}_{2}(\tau^{z})$.
  1. 1.
    If $\mathcal{A}_{i}$ is $(\varepsilon_{i},\delta_{i})$-QDP in the framework $\left(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M}_{i}\right)$, for $i\in\{1,2\}$, then the composed mechanism satisfies $\left(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2}\right)$-QDP in $\left(\mathcal{S},\mathcal{Q}^{(2)},\Theta,\bigotimes_{i=1}^{2}\mathcal{M}_{i}\right)$
  2. 2.
    If $\mathcal{A}_{i}$ is $(\varepsilon_{i},\delta_{i})$-QDP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, for $i\in\{1,2\}$, then the composed mechanism satisfies $(\varepsilon^{\prime},\delta^{\prime})$-QDP in $\left(\mathcal{S},\mathcal{Q}^{(2)},\Theta,\bar{\mathcal{M}}^{2}\right)$ with
    $\varepsilon^{\prime}$ $\coloneqq\varepsilon_{1}+\varepsilon_{2}+\ln\!\left(\frac{1}{(1-\delta_{1})(1-\delta_{2})}\right),$ (124)
    $\delta^{\prime}$ $\coloneqq\sqrt{\delta_{1}(2-\delta_{1})}+\sqrt{\delta_{2}(2-\delta_{2})}.$ (125)
    and also satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta)$ in the same framework with $\delta\coloneqq\min\{\delta_{1}+e^{\varepsilon_{1}}\delta_{2},\delta_{2}+e^{\varepsilon_{2}}\delta_{1}\}$.
Adaptive composability: Suppose that $\mathcal{A}_{1}$ satisfies $(\varepsilon_{1},\delta_{1})$-QDP and $\mathcal{A}_{2}$ chosen adaptively satisfies $(\varepsilon_{2},\delta_{2})$-QDP, as in (119). Then, the composed mechanism in Fig. 3 with $\sigma_{I}$ in (121) satisfies $(\varepsilon_{1}+\varepsilon_{2},\delta_{2}+\delta_{1}|\mathcal{Y}|)$ in the framework $(\mathcal{S},\mathcal{Q}\times\mathcal{Q},\Theta,\bar{\mathcal{M}}\otimes\bar{\mathcal{M}})$.
Proof:
Item 1 in the parallel composability part follows by a similar argument as given in the proof of Property 3 from Theorem 1. For the proof of Item 2, first, we use quasi-convexity of the DL divergence (property 2 in Proposition 2) and then adapt Item 3 of Theorem 2. The adaptive composition result follows along the same lines as the proof of Proposition 3 for fixed $z$, and then averaging over all $z\in\mathcal{Z}$ gives the desired result. ∎
Remark 8 (Extensions beyond flexible QDP).
Corollary 1 does not hold for the general QPP framework. Indeed, it fails to hold, for instance, for the classical PP framework [3, Theorem 9.1]. Nevertheless, Corollary 1 can be extended to account for input states $\sum_{x\in\mathcal{X}}P_{X}(x)\,\rho^{x}\otimes\rho^{x}$ subjected to additional structural assumptions on the class of admissible distributions:
$\Theta\subseteq\left\{P_{X}\in\mathcal{P}(\mathcal{X}):\begin{array}[]{cc}\forall\ (\mathcal{R},\mathcal{T})\in\mathcal{Q},\ \exists\ x,x^{\prime}\in\mathcal{X}&\\ \textnormal{s.t.}\ q_{\mathcal{R}}(x)=q_{\mathcal{T}}(x^{\prime})=1&\end{array}\right\},$ (126)
where $q_{\mathcal{R}}$ and $q_{\mathcal{T}}$ are defined as in Definition 4. The classical version of this condition for PP is known as “universally composable scenarios” [3, Corollary 9.4].
V-B Mechanisms for QPP
We propose mechanisms to achieve $\varepsilon$-QPP and $(\varepsilon,\delta)$-QPP using the depolarization channel. In addition, we provide a general procedure to generate $(\varepsilon,\delta)$-(classical) PP mechanisms using a quantum mechanism satisfying $(\varepsilon,\delta)$-QPP.
V-B1 Depolarization Mechanism
Let
$\mathcal{A}^{p}_{\mathrm{Dep}}(\rho)\coloneqq(1-p)\rho+\frac{p}{d}I,$ (127)
where $p\in[0,1]$ and $d$ is the dimension of the Hilbert space on which $\rho$ acts.
Figure 4: Depolarization mechanism to achieve QPP: This corresponds to a channel $\mathcal{E}$ followed by a depolarizing channel. Note that we can choose $\mathcal{E}=\mathcal{I}$ to be the identity channel as well.
Theorem 3 ($\varepsilon$-QPP depolarization mechanism).
Fix $p\in[0,1]$ and a privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$. Let $\mathcal{E}$ be a quantum channel. Then $\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\cdot)\right)$ (in Fig. 4) is $\varepsilon$-QPP if
$p\geq\frac{dK}{dK+e^{\varepsilon}-1},$ (128)
where
$K\coloneqq\sup_{\mathrm{M}\in\mathcal{M}}\frac{\left\|\mathrm{M}\right\|_{\infty}}{\mathrm{Tr}\!\left[\mathrm{M}\right]}\times\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\frac{\left\|\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right\|_{1}}{2}.$ (129)
This further implies that the depolarization channel with parameter $p$ achieves $\varepsilon$-QPP whenever
$\varepsilon\geq\ln\!\left(1+\frac{(1-p)dK}{p}\right).$ (130)
Proof:
Fix $P_{X}\in\Theta$, $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, and $\mathrm{M}\in\mathcal{M}$, and consider that
$\frac{\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\rho^{\mathcal{R}})\right)\right]}{\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\rho^{\mathcal{T}})\right)\right]}-1$
$=\frac{(1-p)\mathrm{Tr}\!\left[\mathrm{M}\mathcal{E}(\rho^{\mathcal{R}})\right]+\frac{p}{d}\mathrm{Tr}\!\left[\mathrm{M}\right]}{(1-p)\mathrm{Tr}\!\left[\mathrm{M}\mathcal{E}(\rho^{\mathcal{T}})\right]+\frac{p}{d}\mathrm{Tr}\!\left[\mathrm{M}\right]}-1$ (131)
$=\frac{(1-p)\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]}{(1-p)\mathrm{Tr}\!\left[\mathrm{M}\mathcal{E}(\rho^{\mathcal{T}})\right]+\frac{p}{d}\mathrm{Tr}\!\left[\mathrm{M}\right]}$ (132)
$\leq\frac{(1-p)\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]}{\frac{p}{d}\mathrm{Tr}\!\left[\mathrm{M}\right]}$ (133)
Given the above, if
$\varepsilon\geq\ln\!\left(1+\frac{d(1-p)\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]}{p\mathrm{Tr}\!\left[\mathrm{M}\right]}\right),$ (134)
then
$\frac{\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\rho^{\mathcal{R}})\right)\right]}{\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\rho^{\mathcal{T}})\right)\right]}\leq e^{\varepsilon}.$ (135)
Recalling that $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ if and only if $(\mathcal{T},\mathcal{R})\in\mathcal{Q}$, the roles of $\rho^{\mathcal{R}}$ and $\rho^{\mathcal{T}}$ can be interchanged, and we conclude that
$e^{-\varepsilon}\leq\frac{\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\rho^{\mathcal{R}})\right)\right]}{\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\rho^{\mathcal{T}})\right)\right]}.$ (136)
Consider that
$\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]\leq\left\|\mathrm{M}\right\|_{\infty}\frac{\left\|\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right\|_{1}}{2}.$ (137)
Indeed, consider the following Jordan–Hahn decomposition $\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})=P-Q$, where $P$ and $Q$ are the positive and negative parts of $\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})$, respectively, satisfying $P,Q\geq 0$ and $PQ=0$. Then
$\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]$ (138)
$=\mathrm{Tr}\!\left[\mathrm{M}\!\left(P-Q\right)\right]$ (139)
$\leq\mathrm{Tr}\!\left[\mathrm{M}P\right]$ (140)
$\leq\left\|\mathrm{M}\right\|_{\infty}\frac{\left\|\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right\|_{1}}{2},$ (141)
where the last inequality follows from Hölder’s inequality and because $\left\|\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right\|_{1}=\mathrm{Tr}[P]+\mathrm{Tr}[Q]=2\mathrm{Tr}[P]$ since $\mathrm{Tr}\!\left[\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right]=0=\mathrm{Tr}[P-Q]$.
Collecting all terms and supremizing over $\mathcal{M}$ and $\Theta$ and all secret pairs of $\mathcal{Q}$ yields the desired result. ∎
Corollary 2 ($\varepsilon$-QDP with domain knowledge).
Fix $p\in[0,1]$, and a privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ for QDP that encodes domain knowledge. Let $\mathcal{E}$ be a quantum channel. Then $\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\cdot)\right)$ is $\varepsilon$-QPP with
$\varepsilon\geq\ln\!\left(1+\frac{(1-p)d}{2p}k^{\prime}\sup_{\mathrm{M}\in\mathcal{M}}\frac{\|\mathrm{M}\|_{\infty}}{\mathrm{Tr}\!\left[\mathrm{M}\right]}\right),$ (142)
where
$k^{\prime}$ $\coloneqq\sup_{(\rho^{x_{1}},\rho^{x_{2}})\in\mathcal{W}_{\Theta}}\left\|\mathcal{E}(\rho^{x_{1}})-\mathcal{E}(\rho^{x_{2}})\right\|_{1}$ (143)
$\mathcal{W}_{\Theta}$$\coloneqq\{(\rho^{x_{1}},\rho^{x_{2}})\in\mathcal{Q}\ |\ \exists P_{X}\in\Theta\ \!\ P_{X}(x_{1}),P_{X}(x_{2})>0\}.$ (144)
Note that the domain knowledge encoded into the QDP framework may guide towards an improved accuracy/utility, as opposed to considering all neighboring states as secret pairs and all possible measurements. For the QDP framework without domain knowledge, [8, Theorem 3] shows that
$\varepsilon\geq\ln\!\left(1+\frac{(1-p)d}{2p}\sup_{\rho\sim\sigma}\left\|\rho-\sigma\right\|_{1}\right)$ (145)
is a sufficient condition to ensure $(\varepsilon,0)$-QDP. Compared with that due to the condition
$\sup_{\mathrm{M}\in\mathcal{M}}\frac{\left\|\mathrm{M}\right\|_{\infty}}{\mathrm{Tr}\!\left[\mathrm{M}\right]}\times\sup_{(\rho^{x_{1}},\rho^{x_{2}})\in\mathcal{W}_{\Theta}}\left\|\mathcal{E}(\rho^{x_{1}})-\mathcal{E}(\rho^{x_{2}})\right\|_{1}\\ \leq\sup_{\rho\sim\sigma}\left\|\rho-\sigma\right\|_{1},$ (146)
a QDP framework that has the capability to incorporate domain knowledge may cause less perturbation to the useful channel output of $\mathcal{E}$ in some cases. The rightmost inequality holds because $\mathcal{W}_{\Theta}$ includes only the neighboring pairs of states such that their occurrence has a positive probability, while $\rho\sim\sigma$ denotes all possible neighboring pairs.
Remark 9 (Local DP).
For the setup in Section III-C2, Theorem 3 reduces to
$p\geq\frac{d}{d+e^{\varepsilon}-1},$ (147)
with the choice of the identity channel instead of $\mathcal{E}$ in Fig. 4. This occurs because $\left\|\rho-\sigma\right\|_{1}\leq 2$, with equality for pairs of orthogonal states, and $\left\|\mathrm{M}\right\|_{\infty}\leq\mathrm{Tr}\!\left[\mathrm{M}\right]$, with equality holding when considering all possible measurements. This is analogous to a version of the randomized response technique used to achieve classical local DP [48]. For a finite alphabet $\mathcal{X}$ with cardinality $|\mathcal{X}|$, the randomized response mechanism outputs the true value with probability $(1-q)$, and outputs a randomly chosen realization with probability $q/|\mathcal{X}|$. Then, if
$q\geq\frac{|\mathcal{X}|}{|\mathcal{X}|+e^{\varepsilon}-1},$ (148)
$\varepsilon$-local differential privacy is achieved. This analogy further suggests that the depolarization mechanism can be considered as a quantum version of the randomized response mechanism that achieves classical privacy guarantees.
Considering the scenario in which we want to provide a privacy guarantee for all possible measurements (i.e., $\mathcal{M}=\bar{\mathcal{M}}$), next we derive the parameter $p$ to achieve $(\varepsilon,\delta)$-QPP.
Proposition 4 ($(\varepsilon,\delta)$-QPP depolarization mechanism).
Fix $p\in[0,1]$, and the privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ with =$\mathcal{M}=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$. Let $\mathcal{E}$ be a quantum channel. Then $\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\cdot)\right)$is $\varepsilon$-QPP if
$p\geq\min\!\left\{0,\frac{d(K^{\prime}-\delta)}{dK^{\prime}+e^{\varepsilon}-1}\right\},$ (149)
where
$K^{\prime}\coloneqq\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\frac{\left\|\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right\|_{1}}{2}.$ (150)
Proof:
The proof follows from the use of the equivalent formulation through the hockey-stick divergence and the properties of this divergence. By [10, Lemma IV.I], we have
$\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}_{\mathrm{Dep}}\!\left(\mathcal{E}(\rho^{\mathcal{R}})\right)\|\mathcal{A}_{\mathrm{Dep}}\!\left(\mathcal{E}(\rho^{\mathcal{T}})\right)\right)\\ \leq(1-e^{\varepsilon})\frac{p}{d}+(1-p)\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{E}(\rho^{\mathcal{R}})\|\mathcal{E}(\rho^{\mathcal{T}})\right).$ (151)
We also have the property [10, Lemma II.4]
$\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{E}(\rho^{\mathcal{R}})\|\mathcal{E}(\rho^{\mathcal{T}})\right)\leq\frac{\left\|\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right\|_{1}}{2}.$ (152)
Combining these relations, and supremizing over $\Theta$, and secret pairs $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, we can choose
$\delta\geq(1-e^{\varepsilon})\frac{p}{d}+(1-p)K^{\prime}.$ (153)
Then, rearranging the terms we arrive at
$p\geq\frac{d(K^{\prime}-\delta)}{dK^{\prime}+e^{\varepsilon}-1}.$ (154)
Since $p\geq 0$, when $K^{\prime}-\delta\leq 0$, we set $p=0$. ∎
V-B2 Classical PP Mechanisms from QPP Mechanisms
The QPP formalism provides a direct methodology to design classical PP mechanisms with the assistance of QPP mechanisms. In this case, we use quantum encoding to convert classical data to quantum data. We denote the quantum encoding of classical data $x\in\mathcal{X}^{n\times k}$ as $\rho^{x}\coloneqq|x\rangle\!\langle x|$. Then, we ensure the privacy for the quantum data (quantum encoding) such that the privacy is ensured for the underlying classical data.
Proposition 5 ($(\varepsilon,\delta)$- (classical) PP mechanism).
Given an ($\varepsilon,\delta$)-QPP mechanism $\mathcal{A}$ within the framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ when $\rho^{x}\coloneqq|x\rangle\!\langle x|$ with
$\mathcal{S}$ =$=\left\{\{\rho^{x}:x\in\mathcal{R}_{c}\}:\mathcal{R}_{c}\in\mathcal{S}_{c}\right\},$ (155)
$\mathcal{Q}$ ={($=\left\{\left(\{\rho^{x}:x\in\mathcal{R}_{c}\},\{\rho^{x}:x\in\mathcal{T}_{c}\}\right):(\mathcal{R}_{c},\mathcal{T}_{c})\in\mathcal{Q}_{c}\}\right\},$
$\Theta$ =$=\{\{p_{X}(x),\rho^{x}\}_{x}:p_{X}\in\Theta_{c}\},$
$\mathcal{M}$ =$=\left\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\right\},$
any post-processing of $\mathcal{A}$ by a quantum channel $\mathcal{J}$ followed by applying a POVM $\{\mathrm{M}_{y}\}_{y\in\mathcal{Y}}$ denoted as $A:~{}\mathcal{X}^{n\times k}\to\mathcal{Y}$ as shown in Fig. 5 is $(\varepsilon,\delta)$-PP in the framework $(\mathcal{S}_{c},\mathcal{Q}_{c},\Theta_{c})$.
Furthermore, for a selected post-processing $\mathcal{J}$ and POVM $\{\mathrm{M}_{y}\}_{y\in\mathcal{Y}}$, it is sufficient for $\mathcal{M}^{\mathcal{J}}_{c}\subseteq\mathcal{M}$ for $A$ to be $(\varepsilon,\delta)$-PP, where
$\mathcal{M}^{\mathcal{J}}_{c}\coloneqq\left\{\mathcal{J}^{\dagger}\left(\sum_{y\in\mathcal{B}}\mathrm{M}_{y}\right):\mathcal{B}\in\mathcal{Y}\right\}.$ (156)
Figure 5: Generation of classical PP mechanisms from QPP mechanism $\mathcal{A}$: First the classical data is encoded using quantum encoding techniques, then the QPP mechanism $\mathcal{A}$, and if needed any other channel $\mathcal{J}$, and finally the measurement channel.
Proof:
Fix $\mathcal{B}\subseteq\mathcal{Y}$. Consider that
$\mathbb{P}\!\left(A(X)\in\mathcal{B}\middle|\mathcal{R}_{c}\right)$
=($=\frac{\mathbb{P}\!\left(\{A(X)\in\mathcal{B}\}\cap\mathcal{R}_{c}\right)}{\mathbb{P}\!\left(\mathcal{R}_{c}\right)}$ (157)
=$\stackrel{{\scriptstyle(a)}}{{=}}\frac{\sum_{x\in\mathcal{R}_{c}}p(x)\mathbb{P}\!\left(A(x)\in\mathcal{B}\right)}{P_{X}(\mathcal{R})}$ (158)
=$\stackrel{{\scriptstyle(b)}}{{=}}\frac{\sum_{x\in\mathcal{R}_{c}}p(x)\sum_{y\in\mathcal{B}}\mathbb{P}\!\left(A(x)=y\right)}{P_{X}(\mathcal{R})}$ (159)
=$\stackrel{{\scriptstyle(c)}}{{=}}\frac{\sum_{\rho^{x}\in\mathcal{R}}p(x)\sum_{y\in\mathcal{B}}\mathrm{Tr}\!\left[\mathrm{M}_{y}\mathcal{J}\circ\mathcal{A}(\rho^{x})\right]}{P_{X}(\mathcal{R})}$ (160)
=$\stackrel{{\scriptstyle(d)}}{{=}}\mathrm{Tr}\!\left[\sum_{y\in\mathcal{B}}\mathrm{M}_{y}\mathcal{J}\circ\mathcal{A}\left(\sum_{\rho^{x}\in\mathcal{R}}\frac{p(x)}{P_{X}(\mathcal{R})}\rho^{x}\right)\right]$ (161)
=$\stackrel{{\scriptstyle(e)}}{{=}}\mathrm{Tr}\!\left[\mathcal{J}^{\dagger}\left(\sum_{y\in\mathcal{B}}\mathrm{M}_{y}\right)\mathcal{A}\left(\rho^{\mathcal{R}}\right)\right]$ (162)
=$\stackrel{{\scriptstyle(f)}}{{=}}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right],$ (163)
where: (a) from $\mathcal{R}\coloneqq\{\rho^{x}:x\in\mathcal{R}_{c}\}$; (b) from $\mathcal{B}$ being a collection of $y\in\mathcal{Y}$; (c) from $\mathrm{M}_{y}$ being the measurement applied to obtain the outcome $y$; (d) from the linearity of trace operator and quantum channels $\mathcal{A},\mathcal{J}$; (e) from the definition of $\rho^{\mathcal{R}}$, and $\mathcal{J}^{\dagger}$ being the adjoint of $\mathcal{J}$; and finally (f) from $\mathrm{M}~{}\coloneqq~{}\mathcal{J}^{\dagger}\left(\sum_{y\in\mathcal{B}}\mathrm{M}_{y}\right)$.
Similarly, $\mathbb{P}\!\left(A(X)\in\mathcal{B}\middle|\mathcal{T}_{c}\right)=\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]$. Then with the assumption that $\mathcal{A}$ is $(\varepsilon,\delta)$-QPP for $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ mentioned in the proposition statement, we have
$\mathbb{P}\!\left(A(X)\in\mathcal{B}\middle|\mathcal{R}\right)\leq e^{\varepsilon}\hskip 2.84526pt\mathbb{P}\!\left(A(X)\in\mathcal{B}\middle|\mathcal{T}\right)+\delta.$ (164)
concluding the proof. ∎
Depolarization is a common kind of noise considered in quantum information processing. Thus, this provides a method for designing classical PP mechanisms by combining the results presented in Proposition 5 and Theorem 3. It is important to note that quantum encoding of classical data would require computational resources, shifting the complexity of the mechanism design phase to the encoding phase.
VI Quantifying Privacy-Utility Tradeoff
It is important to identify how the performance or utility degrades with the application of a privacy mechanism. Although the application of a privacy mechanism has the intended effect of increasing the privacy of data, it can have the detrimental effect of corrupting the data as well, thus reducing the utility of the data. Thus, it is a goal to preserve utility as much as possible while maintaining privacy. However, there is a tradeoff when trying to attain both goals. Here we define a utility metric to help quantify this tradeoff, and we prove that it can be characterized by an SDP. We then use that metric to analyse privacy-utility tradeoffs related to privacy mechanisms.
VI-A Defining Utility Metrics
We now propose a utility metric based on an operational approach. Let $\mathcal{A}$ denote a privacy mechanism. We focus on the capability of reversing the action of $\mathcal{A}$ by post-processing it with $\mathcal{B}$, in order to recover the original state input to $\mathcal{A}$ up to an error $\gamma$. Based on that operational perspective, we define $\gamma$-utility in terms of how distinguishable $\mathcal{B}\circ\mathcal{A}$ is from the identity channel, and we use the normalized diamond distance as the distinguishability measure.
Definition 5 ($\gamma$-utility).
Let $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{C})$ be a privacy mechanism, and fix $\gamma\in[0,1]$. We say that $\mathcal{A}$ satisfies $\gamma$-utility if
$\mathsf{U}(\mathcal{A})\coloneqq\inf_{\mathcal{B}}\frac{1}{2}\left\|\mathcal{I}-\mathcal{B}\circ\mathcal{A}\right\|_{\diamond}\leq\gamma,$ (165)
where the infimum is taken over every quantum channel $\mathcal{B}:\mathcal{L}(\mathcal{H}_{C})\to\mathcal{L}(\mathcal{H}_{D})$.
The defined utility metric can be reformulated as an SDP by using the dual SDP form of the diamond distance [49, Section 4] and rewriting the quantum channel $\mathcal{B}$ in terms of its Choi matrix, as well as translating the conditions for $\mathcal{B}$ to be a channel to conditions on its Choi matrix. In the following proposition, we employ the following facts. The Choi matrix of the composed channel $\mathcal{B}\circ\mathcal{A}$ is given by [27, Eq. (3.2.22)]
$\Gamma_{AD}^{\mathcal{B}\circ\mathcal{A}}\coloneqq\mathrm{Tr}_{C}\!\left[\left(I_{A}\otimes\Gamma_{CD}^{\mathcal{B}}\right)\left(\mathrm{T}_{C}(\Gamma_{AC}^{\mathcal{A}})\otimes I_{D}\right)\right],$ (166)
and $\Gamma$ represents the Choi matrix with the subscripts showing the input and the output system of the channel while the superscript indicating the channel considered, with no superscript for the identity channel. Recall that $\mathrm{T}_{C}$ is the partial transpose operation acting on system $C$, and $\mathcal{B}$ is a quantum channel if and only if its Choi matrix $\Gamma_{CD}^{\mathcal{B}}$ satisfies $\Gamma_{CD}^{\mathcal{B}}\geq 0$ and $\mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{c}$.
Proposition 6 (SDP formulation of $\gamma$-utility).
$\mathsf{U}(\mathcal{A})=\inf_{\begin{subarray}{c}\mu\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\mu:\\ Z_{AD}\geq\Gamma_{AD}-\Gamma_{AD}^{\mathcal{B}\circ\mathcal{A}},\\ \mu I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{c}\end{array}\right\}\leq\gamma.$ (167)
The above presented SDP formulation can be used to compute the utility of a privacy mechanism $\mathcal{A}$ of interest.
VI-B Analysis of Depolarization Mechanism
In this section, we instantiate $\mathcal{A}$ to be the depolarizing channel with parameter $p$, denoted as $\mathcal{A}^{p}_{\mathrm{Dep}}$ and defined previously in (127).
Proposition 7 (Utility from depolarization mechanism).
Fix $p\in[0,1]$. The depolarization mechanism satisfies $\gamma$-utility if and only if
$\mathsf{U}(\mathcal{A}_{\mathrm{Dep}}^{p})=\frac{p(d^{2}-1)}{d^{2}}\leq\gamma.$ (168)
Proof:
Our proof of Proposition 7 relies on identifying that the utility metric is minimized by setting $\mathcal{B}=\mathcal{I}$, and then evaluating $\left\|\mathcal{I}-\mathcal{A}_{\mathrm{Dep}}^{p}\right\|_{\diamond}$ using the Choi states of the channels $\mathcal{A}_{\mathrm{Dep}}^{p}$ and $\mathcal{I}$, due to the joint covariance of the two channels under unitaries (i.e., $\mathcal{A}_{\mathrm{Dep}}^{p}\circ\mathcal{U}=\mathcal{U}\circ\mathcal{A}_{\mathrm{Dep}}^{p}$ and $\mathcal{I}\circ\mathcal{U}=\mathcal{U}\circ\mathcal{I}$ for every unitary channel $\mathcal{U}$).
Consider that
$\left\|\mathcal{I}-\mathcal{B}\circ\mathcal{A}_{\mathrm{Dep}}^{p}\right\|_{\diamond}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left\|\mathcal{U}\circ(\mathcal{I}-\mathcal{B}\circ\mathcal{A}_{\mathrm{Dep}}^{p})\circ\mathcal{U}^{\dagger}\right\|_{\diamond}$ (169)
=$\stackrel{{\scriptstyle(b)}}{{=}}\left\|\mathcal{I}-\mathcal{U}\circ\mathcal{B}\circ\mathcal{U}^{\dagger}\circ\mathcal{A}_{\mathrm{Dep}}^{p})\right\|_{\diamond}$ (170)
=$\stackrel{{\scriptstyle(c)}}{{=}}\int\,\mathsf{d}\mathcal{U}\,\left\|\mathcal{I}-\mathcal{U}\circ\mathcal{B}\circ\mathcal{U}^{\dagger}\circ\mathcal{A}_{\mathrm{Dep}}^{p})\right\|_{\diamond}$ (171)
$\stackrel{{\scriptstyle(d)}}{{\geq}}\left\|\mathcal{I}-\left(\int\,\mathsf{d}\mathcal{U}\ \mathcal{U}\circ\mathcal{B}\circ\mathcal{U}^{\dagger}\right)\circ\mathcal{A}_{\mathrm{Dep}}^{p})\right\|_{\diamond},$ (172)
where: (a) follows from the unitary invariance of the diamond norm with $\mathcal{U}$ representing a unitary channel; (b) from the commutative property of $\mathcal{A}_{\mathrm{Dep}}^{p}$ with every unitary channel; (c) from the left-hand side being independent of $\mathcal{U}$; and (d) from the convexity of the diamond norm.
Next, observe that $\mathcal{B}^{\star}\coloneqq\int\,\mathsf{d}\mathcal{U}\ \mathcal{U}\circ\mathcal{B}\circ\mathcal{U}^{\dagger}$ is a quantum channel, and this is equal to a depolarization channel [50]. Then, $\mathcal{B}^{\star}=\mathcal{A}_{\mathrm{Dep}}^{q}$ for some $q\in[0,1]$.
The composition of two depolarization channels
$\mathcal{B}^{\star}=\mathcal{A}_{\mathrm{Dep}}^{q}\circ\mathcal{A}_{\mathrm{Dep}}^{p}$ (173)
is also a depolarization channel with parameter $p^{\star}\coloneqq 1-(1-p)(1-q)$. The minimum value is attained by the choice $q=0$, where $\mathcal{A}_{\mathrm{Dep}}^{q}$ in that case corresponds to the identity channel.
With that, we arrive at
$\inf_{\mathcal{B}}\left\|\mathcal{I}-\mathcal{B}\circ\mathcal{A}_{\mathrm{Dep}}^{p}\right\|_{\diamond}=\left\|\mathcal{I}-\mathcal{A}_{\mathrm{Dep}}^{p}\right\|_{\diamond}.$ (174)
With the property of joint covariance of $\mathcal{I}$ and $\mathcal{A}_{\mathrm{Dep}}^{p}$ under unitaries [27, Proposition 7.82], we simplify this to
$\left\|\mathcal{I}-\mathcal{A}_{\mathrm{Dep}}^{p}\right\|_{\diamond}=\frac{1}{d}\left\|\Gamma_{AD}-\Gamma_{AD}^{\mathcal{A}_{\mathrm{Dep}}^{p}}\right\|_{1}.$ (175)
Then consider that
$\frac{1}{d}\left\|\Gamma_{AD}-\Gamma_{AD}^{\mathcal{A}_{\mathrm{Dep}}^{p}}\right\|_{1}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\frac{1}{d}\left\|\Gamma_{AD}-\left((1-p)\Gamma_{AD}+\frac{p}{d}I_{d^{2}}\right)\right\|_{1}$ (176)
$=\frac{p}{d}\left\|\Gamma_{AD}-\frac{1}{d}I_{d^{2}}\right\|_{1}$ (177)
$=p\left\|\frac{\Gamma_{AD}}{d}\left(1-\frac{1}{d^{2}}\right)-\frac{1}{d^{2}}\left(I_{d^{2}}-\frac{\Gamma_{AD}}{d}\right)\right\|_{1}$ (178)
=$\stackrel{{\scriptstyle(b)}}{{=}}p\left(1-\frac{1}{d^{2}}\right)\left(\left\|\frac{\Gamma_{AD}}{d}\right\|_{1}+\left\|\frac{I_{d^{2}}-\frac{\Gamma_{AD}}{d}}{d^{2}-1}\right\|_{1}\right)$ (179)
=$\stackrel{{\scriptstyle(c)}}{{=}}2p\left(1-\frac{1}{d^{2}}\right),$ (180)
where: (a) from $\Gamma_{AD}^{\mathcal{A}_{\mathrm{Dep}}^{p}}=(1-p)\Gamma_{AD}+\frac{p}{d}I_{d^{2}}$; (b) from $\frac{\Gamma_{AD}}{d}$, and $I_{d^{2}}-\frac{\Gamma_{AD}}{d}$ being orthogonal; and (c) from trace norm of quantum states being equal to one.
Combining the above chain of arguments together completes the proof. ∎
Remark 10 (Employing other metrics of utility: fidelity and trace distance).
Employing trace distance as the metric, we define $\beta$-utility as
$\sup_{\rho\in\mathcal{D}(\mathcal{H})}\frac{1}{2}\left\|\rho-\mathcal{A}(\rho)\right\|_{1}\leq\beta,$ (181)
where $\mathcal{D}(\mathcal{H})$ is the set of all density matrices. Then, for the depolarization mechanism we have
$\sup_{\rho\in\mathcal{D}(\mathcal{H})}\frac{1}{2}\left\|\rho-\mathcal{A}_{\mathrm{Dep}}^{p}(\rho)\right\|_{1}=\frac{p(d-1)}{d}\leq\beta,$ (182)
which follows by observing that the supremum is achieved on pure states (due to convexity of the trace norm) and by following arguments similar to those in the proof of Proposition 7. Observe that, similar to Proposition 7, the utility degrades linearly with $p$ (recall that utility degrades with increasing $\beta,\gamma$).
Alternatively, we can consider the following utility metric based on fidelity:
$\inf_{\rho\in\mathcal{D}(\mathcal{H})}\mathsf{F}\!\left(\rho,\mathcal{A}(\rho)\right)\leq\zeta.$ (183)
Compared to the other metrics, by fidelity being a similarity metric rather than a distinguishability metric, higher $\zeta$ implies higher utility gain. For the depolarization mechanism, we have
$\inf_{\rho\in\mathcal{D}(\mathcal{H})}\mathsf{F}\!\left(\rho,\mathcal{A}^{p}_{\mathrm{Dep}}(\rho)\right)=1-p+\frac{p}{d}\leq\zeta.$ (184)
This follows due to the joint concavity of the root fidelity, so that the infimum is achieved by pure states. Lastly, consider that
$\mathsf{F}\left(\mathcal{A}_{\mathrm{Dep}^{p}}(\psi),\psi\right)$ $=\langle\psi|\mathcal{A}^{p}_{\mathrm{Dep}}(\psi)|\psi\rangle$ (185)
$=\langle\psi|\left[(1-p)|\psi\rangle\!\langle\psi|+\frac{\mathrm{I}_{d}}{d}\right]|\psi\rangle$ (186)
$=(1-p)|\langle\psi|\psi\rangle|^{2}+\frac{p}{d}\langle\psi|\psi\rangle$ (187)
$=(1-p)+\frac{p}{d},$ (188)
to arrive at the desired expression.
Next, we focus on understanding the privacy-utility tradeoff in terms of the parameter $p$. This can also be understood as figuring out whether a given noisy quantum channel, represented as a depolarization mechanism, is suitable for a given application when privacy and utility gains are predetermined.
Figure 6: For fixed $d=2$, the figure depicts the optimum utility changing with parameter $p$ and the achievability of $\varepsilon$-QPP changing with parameter $p$ (any $\varepsilon,\gamma$ values above the drawn curves are achievable for a fixed $p$ value). The value of $K$ encodes the domain knowledge available, where $K=1$ corresponds to no such additional information being available.
From Proposition 7, to achieve $\gamma$-utility, we require that
$p\leq\frac{\gamma d^{2}}{(d^{2}-1)}.$ (189)
In terms of privacy, to achieve $\varepsilon$-QPP in the chosen privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$, from Theorem 3, we require that
$p\geq\frac{dK}{dK+e^{\varepsilon}-1}.$ (190)
These two inequalities showcase the privacy-utility tradeoff of the depolarization mechanism. Thus, it is important to tune $p$ based on the required utility characterized by $\gamma$, as well as the privacy parameter $\varepsilon$.
Furthermore, from Theorem 3, we can also deduce that for a fixed $p$ value, the achievable privacy regime includes
$\varepsilon\geq\ln\!\left(1+\frac{(1-p)dK}{p}\right).$ (191)
Fig. 6 depicts the privacy-utility tradeoff in terms of the parameter $p$. Lower $p$ values achieve lower privacy (higher $\varepsilon$) while the utility gain is higher (lower $\gamma$). Privacy is relatively more sensitive to $p$ compared to utility, which degrades linearly with $p$. We also observe that with the reduction of the value $K$, the achievable privacy regime grows. The parameter $K$ from Theorem 3 encodes the domain knowledge available, which is fed into the privacy model from the $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ QPP framework. This also highlights the importance of encoding such information when available to achieve higher utility gains while providing the required privacy guarantees.
Next, we emphasize the impact of the dimension on the achievability of privacy and utility. In Fig. 7, we observe that the privacy level is more sensitive to the dimension than the utility. Considering the utility of the depolarization mechanism (given by $\frac{p(d^{2}-1)}{d^{2}}$), we can always have the following upper bound on it for every $d$:
$\frac{p(d^{2}-1)}{d^{2}}\leq p,$ (192)
where the upper bound is achieved when $d\to\infty$. However, the achievable privacy level in (191) degrades no larger than order $\ln(d)$. Thus, it is important to identify the right/optimum privacy parameters, especially in high-dimensional setups.
Figure 7: For fixed $K=1$, the figure depicts the optimum utility changing with parameter $p$ and the achievability of $\varepsilon$-QPP changing with parameter $p$ (any $\varepsilon,\gamma$ values above the drawn curves are achievable for a fixed $p$ value) for $d\in\{2,4,8\}$.
In the above analysis for the depolarization mechanism in Fig. 4, we chose $\mathcal{E}=\mathcal{I}$. It is an interesting future work to analyze utility for user-specific $\mathcal{E}$ channels: i.e., choosing $\inf_{\mathcal{B}\in\mathrm{CPTP}}\|\mathcal{E}-\mathcal{B}\circ\mathcal{A}^{p}_{\mathrm{Dep}}\circ\mathcal{E}\|_{\diamond}$ as the utility metric. If $\mathcal{E}$ possesses certain symmetries, then one can make use of arguments similar to those presented previously.
VI-C Characterizing Optimal Privacy-Utility Tradeoff when $\mathcal{Q}$ is a Small Set
In this subsection, we focus on identifying the optimal utility that can be obtained by applying an $(\varepsilon,\delta)$-QPP mechanism. Here, we first focus on the setting in which $\mathcal{Q}=\{(\mathcal{R}_{1},\mathcal{R}_{2}),(\mathcal{R}_{2},\mathcal{R}_{1})\}$, ¯=$\bar{\mathcal{M}}=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$, and $\Theta=\{P_{X}\}$, but the following ideas can be extended to the case when $\mathcal{Q}$ is an arbitrary finite set and $\Theta$ includes finite number of probability distributions. However, the computational complexity involved in identifying the optimal utility increases with the cardinality of the set $\mathcal{Q}$ and $\Theta$, due to the addition of more constraints to the optimization problem.
To incorporate privacy requirements, we use the equivalent formulation of QPP via the DL divergence presented in Proposition 1. To this end, first, we employ the SDP formulated in Lemma 1 to compute the relevant DL divergence and then use that in the optimization of utility. We showcase the use of this SDP in characterizing optimal utility next. It is also important to note that this SDP formulation may also be used as a way to quantify privacy in a given $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ framework, as presented in Proposition 1.
Proposition 8 (Optimal utility for fixed privacy constraints).
The optimal utility, as quantified by the $\gamma$-utility metric, for every privacy mechanism that is $(\varepsilon,\delta)$-QPP in the $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ framework, where $\mathcal{Q}=\{(\mathcal{R}_{1},\mathcal{R}_{2}),(\mathcal{R}_{2},\mathcal{R}_{1})\}$, is given by the following:
$\mathsf{U}\!\left(\varepsilon,\delta,\mathcal{R}_{1},\mathcal{R}_{2}\right)\coloneqq\\ \inf_{\begin{subarray}{c}\mu\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\\ \Gamma_{AC}^{\mathcal{A}}\geq 0\\ \lambda_{1}\geq 0,Y_{1}\geq 0\\ \lambda_{2}\geq 0,Y_{2}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\mu:\\ Z_{AD}\geq\Gamma_{AD}-\mathrm{Tr}_{C}\!\left[\Gamma_{CD}^{\mathcal{B}}\mathrm{T}_{C}(\Gamma_{AC}^{\mathcal{A}})\right],\\ \mu I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{C},\\ \mathrm{Tr}_{D}\!\left[\Gamma_{AC}^{\mathcal{A}}\right]=I_{A},\\ \ln(\lambda_{1})\leq\varepsilon,\\ \mathrm{Tr}\!\left[Y_{1}\right]\leq\delta,\\ Y_{1}\geq\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda_{1}\mathrm{Tr}_{A}\!\left[(\mathrm{T}(\rho^{\mathcal{R}_{2}})\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right],\\ \ln(\lambda_{2})\leq\varepsilon,\\ \mathrm{Tr}\!\left[Y_{2}\right]\leq\delta,\\ Y_{2}\geq\mathrm{Tr}_{A}\!\left[(\mathrm{T}\big{(}\rho^{\mathcal{R}_{2}}\big{)}\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda_{2}\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\end{array}\right\}.$ (193)
Proof:
The proof follows from the SDP formulation of the $\gamma$-utility given in Proposition 6, and the privacy constraints (i.e., $\max\{\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}_{1}})\|\mathcal{A}(\rho^{\mathcal{R}_{2}})\right),\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}_{2}})\|\mathcal{A}(\rho^{\mathcal{R}_{1}})\right)\}\leq\varepsilon$) imposed through the SDP formulation of DL divergence presented in Lemma 1. We also used the fact that for a superoperator $\mathcal{A}$ from system $A$ to $C$, the following equality holds [27, Eq. (3.2.14)]
$\mathcal{A}(\rho^{\mathcal{R}_{1}})=\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right].$ (194)
This concludes the proof. ∎
Remark 11 (Privacy constraints via equivalent formulation through hockey-stick divergence).
Instead of using the DL divergence, we can also encode the privacy constraints through the equivalent formulation in Remark 5. To this end, the dual formulation of the hockey-stick divergence (can be obtained by (45)), as
$\mathsf{E}_{\lambda}(\rho\|\sigma)=\inf_{Z\geq 0}\left\{\mathrm{Tr}\!\left[Z\right]:Z\geq\rho-\lambda\sigma\right\},$ (195)
can also be incorporated to compute the optimum utility.
Remark 12 (Optimal privacy parameters for fixed utility).
To find out the optimal (minimal) privacy parameter $\varepsilon^{\star}$ for a given mechanism $\mathcal{A}$ with the utility constraint $\gamma$, and fixed tolerance $\delta$, first we compute the following quantity:
$\lambda^{\star}_{1}(\mathcal{A},\gamma,\delta)\coloneqq\\ \inf_{\begin{subarray}{c}\lambda\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\\ Y_{1}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\lambda:\\ Z_{AD}\geq\Gamma_{AD}-\mathrm{Tr}_{C}\!\left[\Gamma_{CD}^{\mathcal{B}}\mathrm{T}_{C}(\Gamma_{AC}^{\mathcal{A}})\right],\\ \gamma I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{C},\\ \mathrm{Tr}\!\left[Y_{1}\right]\leq\delta,\\ Y_{1}\geq\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda\mathrm{Tr}_{A}\!\left[(\mathrm{T}(\rho^{\mathcal{R}_{2}})\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right]\end{array}\right\}.$ (196)
Similarly $\lambda^{\star}_{2}$ can be obtained by exchanging $\rho^{\mathcal{R}_{1}}$ and $\rho^{\mathcal{R}_{2}}$. Then the optimal value is given by
$\varepsilon^{\star}(\mathcal{A},\gamma,\delta)\coloneqq\ln\!\left(\max\{\lambda^{\star}_{1}(\mathcal{A},\gamma,\delta),\lambda^{\star}_{2}(\mathcal{A},\gamma,\delta)\right)\}.$ (197)
The optimal (minimal) $\delta$ for a fixed $\varepsilon$ with a utility constraint can be obtained by encoding the privacy constraint through the dual form of hockey-stick divergence, as given in Remark 11.
VII Auditing Privacy Frameworks
Auditing for privacy aims to detect violations in privacy guarantees and reject incorrect algorithms (see [51, 52, 53, 54] for classical approaches). In this section, we focus on how we can employ the tools of quantum information theory, as well as quantum algorithms, to audit the privacy of quantum systems. Here we focus on auditing algorithms for QDP guarantees, and note that the following ideas can be extended to auditing algorithms for privacy guarantees demanded by QPP.
The main idea behind auditing classical privacy frameworks is to translate the privacy requirement to a weaker privacy notion that is efficiently computable. Then, for algorithms that do not satisfy the privacy demanded by the relaxed privacy notion, we can conclude that the original privacy requirement is also violated. The pitfall of this approach is the impossibility of quantifying the gap between the original and relaxed privacy notions. For instance, knowing that the relaxed privacy notion is satisfied does not determine whether the original privacy requirement is also satisfied. However, here we target auditing QDP without translating it to a relaxed privacy notion. Note that the following ideas can be extended to auditing for a relaxed privacy notion as well (for the variants of QPP in Section IX).
VII-A Techniques for Auditing QDP
Using the equivalent formulation in Proposition 1 and combining with the specific choices of $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ as given in Section III-C1, for $(\varepsilon,\delta)$-QDP, we have that
$\sup_{\rho\sim\sigma}\ \overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)\leq\varepsilon.$ (198)
Then, we can compute the left-hand side above by using the SDP formulation of $\overline{\mathsf{D}}^{\delta}\!\left(\cdot\|\cdot\right)$ presented in Lemma 1. This approach is helpful for low-dimensional setups because the time complexity of an SDP computation is polynomial in the dimension of the quantum states. However, the complexity of this approach grows exponentially with the number of qubits. Developing techniques to evaluate the DL divergence using a quantum computer would lead to methods for auditing procedures as well, which is beyond the scope of this work and thus left as a future work.
Next, using the equivalent formulation in Remark 5, consider that $(\varepsilon,\delta)$-QDP is equivalent to
$\sup_{\rho\sim\sigma}\ \mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)\leq\delta,$ (199)
where
$\mathsf{E}_{\gamma}(\rho\|\sigma)$ $\coloneqq\mathrm{Tr}\!\left[(\rho-\gamma\sigma)_{+}\right]$ (200)
$=\frac{1}{2}\left\|\rho-\gamma\sigma\right\|_{1}+\frac{(1-\gamma)}{2},$ (201)
with $\gamma\geq 1$ for quantum states $\rho$ and $\sigma$ [10, Eq. (II.2)]. As shown in (43) and (45), the quantity on the right-hand side of (200) can be evaluated by means of an SDP, which implies the same for the hockey-stick divergence $\mathsf{E}_{\gamma}(\rho\|\sigma)$. Then, auditing QDP reduces to computing $\mathsf{E}_{\gamma}(\rho\|\sigma)$ for $\rho\sim\sigma$. However, as before, the time complexity of this SDP grows exponentially in the number of qubits. Based on the results of [55, 56] and the connection of $\mathsf{E}_{\gamma}(\rho\|\sigma)$ to the trace distance given in (201), we expect that evaluating $\mathsf{E}_{\gamma}(\rho\|\sigma)$ should be computationally difficult even for quantum computers [55, 56]. Nevertheless, there are proposals for evaluating the trace distance using variational quantum algorithms [57, 58] (which however do not give particular runtimes), and for cases in which the quantum states have low rank [59]. In what follows, we consider one approach from [57, 58] (via variational algorithms with parameterized quantum circuits) to estimate the quantity $\left\|\mathcal{A}(\rho)-e^{\varepsilon}\mathcal{A}(\sigma)\right\|_{1}$. Using such an estimate, for a fixed value of the privacy parameter $\varepsilon$, we can validate on which values of $\delta$ the needed guarantees are satisfied.
VII-A1 Using Quantum Circuits to Audit QDP
Figure 8: Quantum circuit assisted in estimating QDP: $U^{\rho},U^{\sigma}$ are the unitaries used to prepare $\rho$ and $\sigma$ by tracing out $R_{1},R_{2}$ systems, respectively. Then $\mathcal{A}$ is applied on the systems $S_{i}$ for $i\in\{1,2\}$. The unitary $Q$ takes inputs $F_{i},B_{i}$ and outputs $F_{i}^{\prime},T_{i}$, where $F_{i}$ and $T_{i}$ are qubit systems. Finally, each of the $T_{i}$ systems is measured and the (classical) output random variable is denoted as $X_{\rho}$ for $i=1$ and $X_{\sigma}$ for $i=2$. Here $X_{\rho},X_{\sigma}\in\{0,1\}$. This procedure is repeated a sufficient number of times, and the outcomes of the trials are used to estimate $\mathbb{P}(X_{\rho}=0)$ and $\mathbb{P}(X_{\sigma}=1)$.
Firstly, let us focus on how to estimate $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$ for a fixed $\rho\sim\sigma$ and $\varepsilon$. With the ideas developed in [57, Algorithm 14], we discuss how a so-called quantum interactive proof can be used for estimating the privacy level. For that, refer to the quantum circuit in Fig. 8: the unitaries $U^{\rho}$ and $U^{\sigma}$ are used to prepare the states $\rho$ and $\sigma$, by tracing out systems $R_{1}$ and $R_{2}$, respectively. Then the algorithm $\mathcal{A}$ is applied on the systems $S_{i}$ for $i\in\{1,2\}$. The unitary $Q$ takes inputs $F_{i},B_{i}$ and outputs $F_{i}^{\prime},T_{i}$, where $F_{i}$ and $T_{i}$ are qubit systems. Finally, both of the $T_{i}$ systems are measured in the standard basis $\{|0\rangle,|1\rangle\}$, and the (classical) output random variable is denoted as $X_{\rho}$ for $i=1$ and $X_{\sigma}$ for $i=2$. Here $X_{\rho}$ and $X_{\sigma}$ take values in $\{0,1\}$. This procedure is repeated a sufficient number of times, and we use the results to estimate $\mathbb{P}(X_{\rho}=0)$ and $\mathbb{P}(X_{\sigma}=1)$.
Next, consider a scenario in which one could maximize the following utility function over all possible choices of $Q$:
$g(Q,\rho,\sigma,\mathcal{A},\varepsilon)\coloneqq\frac{1}{e^{\varepsilon}+1}\mathbb{P}(X_{\rho}=0)+\frac{e^{\varepsilon}}{e^{\varepsilon}+1}\mathbb{P}(X_{\sigma}=1).$ (202)
In quantum complexity terminology, this action could be conducted by a quantum prover who has unbounded computational power (we discuss how to relax this assumption in Remark 13). From [57, Eq. (128)] and the discussion therein, we conclude that
$f(\rho,\sigma,\mathcal{A},\varepsilon)$ $\coloneqq\sup_{Q}g(Q,\rho,\sigma,\mathcal{A},\varepsilon)$ (203)
$=\frac{1}{2}\left(1+\frac{1}{e^{\varepsilon}+1}\left\|\mathcal{A}(\rho)-e^{\varepsilon}\mathcal{A}(\sigma)\right\|_{1}\right),$ (204)
where the optimization is over every unitary $Q$.
If
$f(\rho,\sigma,\mathcal{A},\varepsilon)\leq\frac{1}{2}\left(1+\frac{2\delta+e^{\varepsilon}-1}{e^{\varepsilon}+1}\right),$ (205)
then $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)\leq\delta$, due to (200). Then, by changing the role of $\rho$ and $\sigma$ in (202), we obtain $f(\sigma,\rho,\mathcal{A},\varepsilon)$. Next, we select the maximum out of these quantities
${\widehat{f}}(\rho,\sigma,\mathcal{A},\varepsilon)\coloneqq\max\left\{f(\rho,\sigma,\mathcal{A},\varepsilon),f(\sigma,\rho,\mathcal{A},\varepsilon)\right\}.$ (206)
To this end, if
$\sup_{\rho\sim\sigma}\widehat{f}(\rho,\sigma,\mathcal{A},\varepsilon)\leq\frac{1}{2}\left(1+\frac{2\delta+e^{\varepsilon}-1}{e^{\varepsilon}+1}\right),$ (207)
it can be verified that $\mathcal{A}$ is $(\varepsilon,\delta)$-QDP.
Remark 13 (Relaxing the computationally unbounded assumption).
Finding the $Q$ which achieves the optimum utility in (202) is practically infeasible. To relax this assumption, we replace the action of the prover (who finds this $Q$) with a parameterized circuit $Q_{\theta}$. Then we can use (202) as a utility function for training a variational quantum algorithm [60, 61] to estimate a lower bound for (203). With that we obtain a lower bound on $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$, which we denote as $\mathsf{E}^{\mathrm{LB}}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$. A lower bound gives a sufficient condition for ruling out algorithms that do not satisfy $(\varepsilon,\delta)$-QDP. This claim follows because the estimated lower bound $\mathsf{E}^{\mathrm{LB}}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)>\delta$ implies that $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)>\delta$.
Remark 14 (Neighboring pairs).
To verify $(\varepsilon,\delta)$-QDP, it is required to compute whether $\widehat{f}(\rho,\sigma,\mathcal{A},\varepsilon)\leq\frac{1}{2}\left(1+\frac{2\delta+e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)$ for all neighboring pairs $(\rho,\sigma)\in\mathcal{Q}$. However, checking this requirement has increasing computational complexity as the cardinality of the set $\mathcal{Q}$ increases. If the privacy requirements can be relaxed so as to encode domain knowledge as in QPP framework, the effective set of $\mathcal{Q}$ may be a small set in some applications of interest.
VII-B Formal Guarantees for Auditing QDP
In the previous subsection, we presented possible avenues for auditing QDP, using some tools available in quantum information theory, as well as variational quantum algorithms. Now, we put forward the question “Can we quantify the success of a privacy auditing pipeline for correctly accepting and rejecting an algorithm with given privacy requirements?” The authors of [53, 54] have worked towards answering this question for auditing classical DP, but using a relaxed privacy definition.
In what follows, we propose a hypothesis testing based auditing pipeline for QDP with domain knowledge, which is represented in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$. To obtain a desired significance in the hypothesis test, we employ the trace-norm estimation quantum algorithm proposed in [59], which has at most $\alpha$ additive error from the exact value, with high probability.
Let us define
$\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})\coloneqq\frac{1}{e^{\varepsilon}+1}\left\|\mathcal{A}(\rho)-e^{\varepsilon}\mathcal{A}(\sigma)\right\|_{1}.$ (208)
Fix $(\rho,\sigma)\in\mathcal{Q}$. If algorithm $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-QDP, then $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)\leq\delta$. By applying (200), we arrive at
$\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})\leq\frac{2\delta+e^{\varepsilon}-1}{e^{\varepsilon}+1}\eqqcolon\mathsf{g}(\varepsilon,\delta).$ (209)
Next, by estimating the quantity on the left-hand side, and using $\mathsf{g}(\varepsilon,\delta)$ as a threshold, we design an auditing pipeline for QDP by means of the following null and alternative hypotheses:
$H_{0}$ $:\max\{\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A}),\mathsf{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})\}\leq\mathsf{g}(\varepsilon,\delta),$ (210)
$H_{1}$ $:\max\{\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A}),\mathsf{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})\}>\mathsf{g}(\varepsilon,\delta).$ (211)
Let the estimates of $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ and $\mathsf{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})$ from a randomized algorithm (in our analysis we use the algorithm corresponding to [59, Corollary 3.4]) be $\hat{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ and $\hat{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})$, respectively. We choose
$\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)\coloneqq\max\{\hat{T}^{\varepsilon}(\rho,\sigma,\mathcal{A}),\hat{T}^{\varepsilon}(\sigma,\rho,\mathcal{A})\}$ (212)
as our test statistic.
Estimation of the test statistic:
Lemma 3 (Estimating trace distance using samples of $\rho$, $\sigma$ — restatement of [59, Corollary 3.4]).
Given access to identical copies of $d$-dimensional quantum states $\rho$ and $\sigma$, there is a quantum algorithm that estimates $\mathsf{T}(\rho,\sigma)$ within additive error $\alpha$ and with probability not less than $(1-\beta)$, by using
$O\!\left(\log\!\left(\frac{1}{\beta}\right)\frac{r^{2}}{\alpha^{5}}\log^{2}\!\left(\frac{r}{\alpha}\right)\log^{2}\!\left(\frac{1}{\alpha}\right)\right)$ (213)
samples of $\rho$ and $\sigma$, where $r$ is an upper bound on the rank of $\rho$ and $\sigma$.
Lemma 3 is obtained by using the existing result in Corollary 3.4 of [59], and combining its argument in Theorem 2.6 therein on estimating $\mathrm{Tr}\!\left[A\rho\right]$ within additive error $\alpha$ with probability $1-\beta$, by using $O\!\left(\frac{1}{\alpha^{2}}\log\!\left(\frac{1}{\beta}\right)\right)$ identical samples of $\rho$. The algorithm proposed in [59] is designed based on the following idea. Let $V\coloneqq(\rho-\sigma)/2$. Consider its singular value decomposition as $V=W\Sigma U^{\dagger}$. Then the trace distance can be expressed by the following identity:
$\mathsf{T}(\rho,\sigma)=\frac{1}{2}\left(\mathrm{Tr}\!\left[\rho\,\mathrm{sgn}(V)\right]-\mathrm{Tr}\!\left[\sigma\,\mathrm{sgn}(V)\right]\right),$ (214)
where $\mathrm{sgn}(V)\coloneqq W\mathrm{sgn}(\Sigma)U^{\dagger}$, and $\mathrm{sgn}(\cdot)$ is the sign function. Then, $\mathrm{Tr}\!\left[\rho\,\mathrm{sgn}(V)\right]$ and $\mathrm{Tr}\!\left[\sigma\,\mathrm{sgn}(V)\right]$ can be estimated separately, by combining the techniques of quantum singular value transformation [62] and the Hadamard test [63]. To this end, to implement unitary block-encodings of $\rho$ and $\sigma$ approximately, the technique of density matrix exponentiation [64] is used. The same techniques can be employed to compute $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ since
$\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})=\\ \frac{1}{e^{\varepsilon}+1}\left(\mathrm{Tr}\!\left[\mathcal{A}(\rho)\ \mathrm{sgn}(\mathrm{V}^{\varepsilon})\right]-e^{\varepsilon}\mathrm{Tr}\!\left[\mathcal{A}(\sigma)\ \mathrm{sgn}(\mathrm{V}^{\varepsilon})\right]\right),$ (215)
where $\mathrm{V}^{\varepsilon}\coloneqq(\mathcal{A}(\rho)-e^{\varepsilon}\mathcal{A}(\sigma))/(e^{\varepsilon}+1)$.
Type-1 error analysis
We arrive at the following bound on the Type-1 error of the proposed hypothesis testing pipeline.
Proposition 9 (Type-I error).
Fix arbitrary $\alpha,\delta>0$ and consider the above hypothesis testing pipeline. Then
$\sup_{\rho\sim\sigma}\mathbb{P}\!\left(\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)>\mathsf{g}(\varepsilon,\delta)+\alpha\,\middle|\,H_{0}\right)\leq\beta,$ (216)
if the algorithm from Lemma 3 has access to
$O\!\left(\log\!\left(\frac{1}{\beta}\right)\frac{r^{2}}{\alpha^{5}}\log^{2}\!\left(\frac{r}{\alpha}\right)\log^{2}\!\left(\frac{1}{\alpha}\right)\right)$ (217)
identical copies of the states $\rho$ and $\sigma$, such that $\rho\sim\sigma$ and where
$r\coloneqq\sup_{\rho\sim\sigma}\max\{\mathrm{rank}\!\left(\mathcal{A}(\rho)\right),\mathrm{rank}\left(\mathcal{A}(\sigma)\right)\}.$ (218)
Proof:
Fix $\rho$ and $\sigma$ such that $\rho\sim\sigma$. Under the null hypothesis and the assumption that $\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)={\hat{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})$, we have that
$\mathbb{P}\!\left(\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)>\mathsf{g}(\varepsilon,\delta)+\alpha\right)$
$=\mathbb{P}\!\left({\hat{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})>\mathsf{g}(\varepsilon,\delta)+\alpha\right)$ (219)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\mathbb{P}\!\left({\hat{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})-{{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})>\alpha\right)$ (220)
$\leq\mathbb{P}\!\left(|{\hat{T}}^{\varepsilon}(\rho,\sigma,\mathcal{A})-{{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})|>\alpha\right)$ (221)
$\stackrel{{\scriptstyle(b)}}{{\leq}}\beta,$ (222)
where: (a) follows since ${{\mathsf{T}}}^{\varepsilon}(\rho,\sigma,\mathcal{A})\leq\mathsf{g}(\varepsilon,\delta)$ under the null hypothesis; (b) from the high probability statement in Lemma 3. Similarly, the above inequality holds when $\hat{T}(\rho,\sigma,\mathcal{A},\varepsilon)={\hat{T}}^{\varepsilon}(\sigma,\rho,\mathcal{A})$ concluding the proof. ∎
Proposition 9 provides a bound on the number of samples of the states required to achieve type-I error (significance) of $\beta$. In that case, we would use a threshold of $g(\varepsilon,\delta)+\alpha$ for accepting the null hypothesis, such that the null hypothesis is accepted when the test statistic is less than or equal to $g(\varepsilon,\delta)+\alpha$.
Remark 15 (Computational complexity with the rank $r$).
From Proposition 9, it is evident that the copy complexity of the algorithm grows as $O(r^{2}\log^{2}(r))$, where $r$ is the rank upper bound. Let $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ be a quantum channel. Then, $r\leq d_{B}$, where $d_{B}$ is the dimension of the Hilbert space $\mathcal{H}_{B}$. Considering computational complexity, one possibility is to compose $\mathcal{A}$ with another quantum channel $\mathcal{N}$ that translates the space to a low-dimensional setting. However, due to the data-processing inequality for the trace distance, it will only provide a lower bound. With that, it is possible to reject algorithms if $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{N}\circ\mathcal{A})>g(\varepsilon,\delta)$, which implies that $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})>g(\varepsilon,\delta)$. This approach will then have the limitations equipped with the classical auditing approaches using relaxed privacy notions, since the contraction gap between $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{N}\circ\mathcal{A})$, and $\mathsf{T}^{\varepsilon}(\rho,\sigma,\mathcal{A})$ is hard to quantify. In the quantification of the gap, finding the contraction coefficient $\eta_{\mathcal{N}}$ of the channel $\mathcal{N}$ would be useful if $\eta_{\mathcal{N}}<1$.
In summary, we presented in this section a hypothesis testing pipeline for auditing the privacy of quantum systems while providing formal guarantees on auditing QDP using quantum algorithms designed for trace distance. Analysing the Type-II error of this approach is also an important task to quantify the power of the test, which is left for future work.
VIII Information-Theoretic Bounds from QPP and Connections to Quantum Fairness
In this section, we first explore several information-theoretic bounds that arise from an algorithm satisfying QPP constraints. We then demonstrate how these bounds can be used to establish connections between QPP and quantum fairness [25, 24]. Later on we utilize the results derived in this section to determine the relative strength of the QPP variants introduced in Section IX.
VIII-A Information-Theoretic Bounds from QPP
It was pointed out in [10] that finding bounds on quantum relative entropy and mutual information due to QDP is an interesting open problem. In this section, we address this question in a general setting, which includes QDP as a special case, by giving bounds for relative entropy and Holevo information. We additionally provide bounds in terms of Rényi relative entropies, and trace distance. For the rest of the discussion, fix the privacy framework to be $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$.
Proposition 10 (Bounds on quantum Rényi relative entropy and quantum relative entropy due to QPP).
Fix $\alpha>1$. If $\mathcal{A}$ is $\varepsilon$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$, then
$\mathsf{D}_{\alpha}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\min\!\left\{\frac{\varepsilon^{2}\alpha}{2},\varepsilon\right\},$ (223)
where $\mathsf{D}_{\alpha}(\cdot\|\cdot)$ is an arbitrary quantum Rényi relative entropy satisfying data processing. [^4] Note that Petz–Rényi in (5) satisfies data processing for $\alpha\in(0,1)\cup(1,2]$ and sandwiched Rényi in (8) satisfies data processing for $\alpha\in[1/2,1)\cup(1,\infty)$. Furthermore,
$\mathsf{D}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\min\!\left\{\frac{\varepsilon^{2}}{2},\varepsilon\right\},$ (224)
where $\mathsf{D}$ is the quantum relative entropy in (6).
Proof:
$\varepsilon$-QPP of the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ implies that for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$
$\mathsf{D}_{T}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\varepsilon,$ (225)
where $\mathsf{D}_{T}$ is the Thompson metric from (17). By this implication and the fact that every quantum Rényi-divergence $\mathsf{D}_{\alpha}$ of order $\alpha$ satisfying data processing is bounded from above by $\mathsf{D}_{\max}$ [65, Eq. (4.36)], $\varepsilon$-QPP implies that for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$,
$\mathsf{D}_{\alpha}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\varepsilon.$ (226)
By a different argument via the maximal extension presented in Lemma 4, we can further arrive at the following, for all $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$,
$\mathsf{D}_{\alpha}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\frac{\varepsilon^{2}\alpha}{2}.$ (227)
This completes the proof of the first inequality.
Next, noting that the Petz–Rényi relative entropy in (5) satisfies data processing for $\alpha\in(0,1)\cup(1,2]$, and then taking the limit $\alpha\to 1^{+}$, we arrive at the bound on quantum relative entropy by using the equality in (6). ∎
Lemma 4.
Fix $\alpha>1$, and $\rho,\sigma$ PSD operators. The following inequality holds:
$\mathsf{D}_{\alpha}(\rho\|\sigma)\leq\frac{\alpha}{2}(\mathsf{D}_{T}(\rho\|\sigma))^{2}.$ (228)
Proof:
See Appendix E. ∎
Remark 16 (Operational interpretation of Thompson metric).
An operational interpretation of the Thompson metric has appeared in symmetric postselected hypothesis testing (a setting allowing for an inconclusive outcome along with two general conclusive outcomes and postselecting on the conclusive outcomes) as the asymptotic error exponent of discriminating two quantum states $\rho$ and $\sigma$ [41], as well as in the resource theory of symmetric distinguishability [40]. Here by referring to (225), the QPP framework also provides another operational interpretation of the Thompson metric. In the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, for fixed $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ (by definition also for $(\mathcal{T},\mathcal{R})\in\mathcal{Q}$), the Thompson metric given by $\mathsf{D}_{T}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)$ is equal to the minimal $\varepsilon$ needed to achieve $\varepsilon$-QPP for the secret pair $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$.
Proposition 11 (Bounds on trace norm).
If $\mathcal{A}$ is $\varepsilon$-QPP, then
$\sup_{\begin{subarray}{c}\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}\end{subarray}}\left\|\mathcal{A}(\rho^{\mathcal{R}})-\mathcal{A}(\rho^{\mathcal{T}})\right\|_{1}\leq\min\!\left\{\varepsilon,\sqrt{2\varepsilon}\right\},$ (229)
and if $\mathcal{A}$ is $(\varepsilon,\delta)$-QPP, we have
$\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\left\|\mathcal{A}(\rho^{\mathcal{R}})-\mathcal{A}(\rho^{\mathcal{T}})\right\|_{1}\leq 1-\frac{2(1-\delta)}{e^{\varepsilon}+1}.$ (230)
Proof:
The first inequality holds by applying the quantum Pinsker inequality ($\frac{1}{2}\left\|\rho-\sigma\right\|^{2}_{1}\leq\mathsf{D}(\rho\|\sigma)$) [66, Theorem 1.15] and Proposition 10.
For the second inequality, next we prove a more general relation: $(\varepsilon,\delta)$-QPP implies $(\varepsilon^{\prime},\delta^{\prime})$-QPP for $\varepsilon^{\prime}<\varepsilon$ with
$\delta^{\prime}=1-(e^{\varepsilon^{\prime}}+1)(1-\delta)/(e^{\varepsilon}+1),$ (231)
similar to [6, Property 3] for classical DP. Using that, fixing $\varepsilon^{\prime}$ therein to zero leads to the desired result.
Proof that $(\varepsilon,\delta)$-QPP implies $(\varepsilon^{\prime},\delta^{\prime})$-QPP:
Fix $P_{X}\in\Theta$ and $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$. By assumption, we have that
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]\leq e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\delta.$ (232)
With the above choice for $\delta^{\prime}$, we have $\delta-\delta^{\prime}=(1-\delta)(e^{\varepsilon^{\prime}}-e^{\varepsilon})$. Plugging this in, we find that
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]$
$\leq e^{\varepsilon^{\prime}}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\delta^{\prime}+(\delta-\delta^{\prime})$
$\qquad+(e^{\varepsilon}-e^{\varepsilon^{\prime}})\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]$ (233)
$=e^{\varepsilon^{\prime}}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\delta^{\prime}$
$\qquad+(e^{\varepsilon}-e^{\varepsilon^{\prime}})\left(\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]-\frac{1-\delta}{e^{\varepsilon}+1}\right).$ (234)
Since $\varepsilon^{\prime}\leq\varepsilon$, we get the desired inequality if $\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]\leq\frac{1-\delta}{e^{\varepsilon}+1}$.
By choosing the measurement operator $\mathrm{I}-\mathrm{M}$, we also have
$\mathrm{Tr}\!\left[(\mathrm{I}-\mathrm{M})\mathcal{A}(\rho^{\mathcal{T}})\right]\leq e^{\varepsilon}\mathrm{Tr}\!\left[(\mathrm{I}-\mathrm{M})\mathcal{A}(\rho^{\mathcal{R}})\right]+\delta.$ (235)
Rewriting (235), we arrive at
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]\leq 1-e^{-\varepsilon}(1-\delta)+e^{-\varepsilon}\ \mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right].$ (236)
Similar to the previous manipulations, we get
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]\leq e^{\varepsilon^{\prime}}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\delta^{\prime}\\ +(e^{\varepsilon^{\prime}}-e^{-\varepsilon})\left(-\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\frac{1-\delta}{e^{\varepsilon}+1}\right).$ (237)
Since $\varepsilon^{\prime}\leq\varepsilon$, we arrive at the desired inequality when $\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]\geq\frac{1-\delta}{e^{\varepsilon}+1}$. By these two arguments, the desired inequality holds for either of the cases, proving its validity.
A similar inequality holds for every $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ and $P_{X}\in\Theta$. Thus, the desired implication has been proved. ∎
Bounds on mutual information
Let $X\sim P_{X}$ be a random variable, which can take values in an alphabet $\mathcal{X}$. Depending on $X$, the state $\rho^{X}$ is chosen from the set $\{\rho^{1},\ldots,\rho^{|\mathcal{X}|}\}$. Then the state $\rho^{X}$ is sent through a quantum channel $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ satisfying QLDP (recall this notion from Section III-C2). Afterwards the goal is to identify $X$ by performing a measurement described by the POVM $\{\mathrm{M}_{y}\}_{y\in\mathcal{Y}}$, which realizes the output $Y$. The flow diagram relevant to this setup is shown in Fig. 9.
Here, we focus on how much information about $X$ can be learned from the output of the quantum privacy mechanism $\mathcal{A}(\rho^{X})$ and the classical output $Y$, with an emphasis on the quantities $\mathsf{I}\!\left(X;B\right)_{\sigma}$ and $\mathsf{I}(X;Y)$, respectively, where
$\sigma_{XB}=\sum_{x\in\mathcal{X}}P_{X}(x)\,|x\rangle\!\langle x|\otimes\mathcal{A}(\rho^{x}).$ (238)
We define the Holevo information of the classical–quantum state $\sigma_{XB}$ as
$\mathsf{I}\!\left(X;B\right)_{\sigma}\coloneqq\mathsf{S}\!\left(\sum_{x\in\mathcal{X}}P_{X}(x)\,\mathcal{A}(\rho^{x})\right)\\ -\sum_{x\in\mathcal{X}}P_{X}(x)\,\mathsf{S}\!\left(\mathcal{A}(\rho^{x})\right).$ (239)
By data processing of the Holevo information, we have that $\mathsf{I}(X;Y)\leq\mathsf{I}\!\left(X;B\right)_{\sigma}$. Next, we provide bounds for $\mathsf{I}\!\left(X;B\right)_{\sigma}$ when $\mathcal{A}$ satisfies $\varepsilon$-quantum local DP (QLDP). Recall that, to achieve QLDP, the algorithm $\mathcal{A}$ is designed so that the output of $\mathcal{A}$ may not release much information about the input states (recall Section III-C2).
Figure 9: Setup relevant to Proposition 12: $X$ is a classical random variable that determines $\rho^{X}$, which is the input into the channel $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$. Then $Y$ is a random variable describing the outcome after applying the POVM $\{\mathrm{M}_{y}\}_{y\in\mathcal{Y}}$. Note that here the classical systems related to $X,Y$ are also given by the same system labels $X,Y$.
Proposition 12 (Bounds on Holevo information due to QLDP).
Let $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ be a quantum channel. If $\mathcal{A}$ satisfies $\varepsilon$-QLDP, then
$\mathsf{I}\!\left(X;B\right)_{\sigma}\leq\min\!\left\{\varepsilon,\frac{\varepsilon^{2}}{2}\right\}.$ (240)
Furthermore, if $\mathcal{A}$ is $(\varepsilon,\delta)$-QLDP, then
$\mathsf{I}\!\left(X;B\right)_{\sigma}\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (241)
if $\delta^{\prime}\in[0,1-1/d]$, where $\delta^{\prime}=1-{2(1-\delta)}/{(e^{\varepsilon}+1)}$, $d$ is the dimension of the Hilbert space $\mathcal{H}_{B}$, and
$\mathsf{h}(\delta^{\prime})\coloneqq-\delta^{\prime}\ln(\delta^{\prime})-(1-\delta^{\prime})\ln(1-\delta^{\prime})$ (242)
is the binary Shannon entropy in nats.
Proof:
The following proof ideas are inspired by the works on mutual information based classical DP and PP presented in [6, Theorem 1] and [54, Theorem 1], respectively. For the proof of the first part, consider that
$\mathsf{I}\!\left(X;B\right)_{\sigma}$
$=\sum_{x\in\mathcal{X}}P_{X}(x)\,\mathsf{D}\!\left(\mathcal{A}(\rho^{x})\middle\|\sum_{x^{\prime}\in\mathcal{X}}P_{X}(x^{\prime})\mathcal{A}(\rho^{x^{\prime}})\right)$ (243)
$\leq\sum_{x\in\mathcal{X}}\sum_{x^{\prime}\in\mathcal{X}}P_{X}(x)P_{X}(x^{\prime})\ \mathsf{D}\!\left(\mathcal{A}(\rho^{x})\middle\|\mathcal{A}(\rho^{x^{\prime}})\right)$ (244)
$\leq\min\!\left\{\varepsilon,\frac{\varepsilon^{2}}{2}\right\},$ (245)
where the first inequality follows from the joint convexity of quantum relative entropy [67]. The final inequality follows because $\mathcal{A}$ satisfies $\varepsilon$-QLDP, as well as from Proposition 10 and the setup for QLDP in Section III-C2.
For the second part, consider that
$\mathsf{I}\!\left(X;B\right)_{\sigma}$ $=\mathsf{S}\!\left(\mathcal{A}(\bar{\rho})\right)-\sum_{x\in\mathcal{X}}P_{X}(x)\mathsf{S}\!\left(\mathcal{A}(\rho^{x})\right)$ (246)
$\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (247)
where
$\bar{\rho}=\sum_{x\in\mathcal{X}}P_{X}(x)\rho^{x},$ (248)
and the final inequality follows from the fact that $\left\|\mathcal{A}(\bar{\rho})-\mathcal{A}(\rho^{x})\right\|_{1}\leq\delta^{\prime}$ due to $\mathcal{A}$ satisfying $(\varepsilon,\delta)$-QPP along with Proposition 11. Then, to arrive at the final expression, we use the entropy continuity result known as the Fannes–Audenaert Inequality [68]:
$\left|\mathsf{S}\!\left(\mathcal{A}(\bar{\rho})\right)-\mathsf{S}\!\left(\mathcal{A}(\rho^{x})\right)\right|\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (249)
concluding the proof. ∎
Next, we extend the earlier setup to the case in which we generate $n$ i.i.d. random variables from the distribution of $P_{X}$. Then, $\{X_{1},\ldots,X_{n}\}$ forms a database. Depending on the value of the database, we choose $\rho^{X^{n}}\coloneqq\rho^{X_{1}}\otimes\cdots\otimes\rho^{X_{n}}$. Then $\rho^{X^{n}}$ is passed through a quantum algorithm $\mathcal{A}$ that satisfies QDP, followed by a POVM $\{\mathrm{M}_{y}\}_{y\in\mathcal{Y}}$ that generates the random outcome $Y$ taking values in $\mathcal{Y}$. In this formulation, we take the convention that $\rho\sim\sigma$ (i.e., $\rho$ and $\sigma$ are neighbors) in the QDP framework if $\mathrm{Tr}_{i}\!\left[\rho\right]=\mathrm{Tr}_{i}\!\left[\sigma\right]$ for all $i\in\{1,\ldots,n\}$. Let
$\sigma_{X^{n}B}=\sum_{x\in\mathcal{X}}P_{X^{n}}(x^{n})\,|x^{n}\rangle\!\langle x^{n}|\otimes\mathcal{A}(\rho^{x^{n}}).$ (250)
Proposition 13 (Bounds on mutual information due to QDP).
Let $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ be a quantum channel. For all $i\in\{1,\ldots,n\}$, suppose that $X_{i}$ is drawn i.i.d. from the distribution $P_{X}$. If $\mathcal{A}$ satisfies $\varepsilon$-QDP, then
$\sup_{i\in\{1,\ldots,n\},P_{X}}\mathsf{I}(X_{i};B|X^{n\backslash i})_{\sigma}\leq\min\!\left\{\varepsilon,\frac{\varepsilon^{2}}{2}\right\}.$ (251)
Furthermore, if $\mathcal{A}$ satisfies $(\varepsilon,\delta)$-QDP, then
$\sup_{i\in\{1,\ldots,n\},P_{X}}\mathsf{I}(X_{i};B|X^{n\backslash i})_{\sigma}\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (252)
if $\delta^{\prime}\in[0,1-1/d]$, where $\delta^{\prime}=1-{2(1-\delta)}/{(e^{\varepsilon}+1)}$.
Proof:
First, let us consider $\mathsf{I}(X_{i};B|X^{n\backslash i}=z^{n\backslash i})_{\sigma}$. Define the shorthands
$\rho^{n\backslash i}$ $\coloneqq\rho_{1}^{z_{1}}\otimes\ldots\rho_{i-1}^{z_{i-1}}\otimes\rho_{i+1}^{z_{i+1}}\otimes\rho_{n}^{z_{n}},$ (253)
$\omega$ $\coloneqq\sum_{x\in\mathcal{X}}P_{X_{i}}(x)\rho_{i}^{x}\otimes\rho^{n\backslash i},$ (254)
$\omega^{x}$ $\coloneqq\rho_{i}^{x}\otimes\rho^{n\backslash i}.$ (255)
Then $\omega\sim\omega^{x}$ because $\mathrm{Tr}_{i}\!\left[\sigma\right]=\mathrm{Tr}_{i}\!\left[\sigma^{x}\right]$.
For the first part, it thus follows that
$\mathsf{I}(X_{i};B|X^{n\backslash i}=z^{n\backslash i})_{\sigma}\\ =\sum_{x\in\mathcal{X}}P_{X_{i}}(x)\ \mathsf{D}\!\left(\mathcal{A}(\omega^{x})\middle\|\mathcal{A}(\omega)\right)\leq\min\!\left\{\varepsilon,\frac{\varepsilon^{2}}{2}\right\}.$ (256)
The last inequality holds because $\mathcal{A}$ satisfies $\varepsilon$-QDP, along with Proposition 10. Since this inequality holds for all possible $z^{n\backslash i}$ such that $\{X^{n\backslash i}=z^{n\backslash i}\}$, the desired relation holds.
For the second part, consider that
$\mathsf{I}(X_{i};B|X^{n\backslash i}=z^{n\backslash i})_{\sigma}$
$=\mathsf{S}\!\left(\mathcal{A}({\omega})\right)-\sum_{x\in\mathcal{X}}P_{X_{i}}(x)\mathsf{S}\!\left(\mathcal{A}(\omega^{x})\right)$ (257)
$\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (258)
where the last inequality holds because $\left\|\mathcal{A}(\omega)-\mathcal{A}(\omega^{x})\right\|_{1}\leq\delta^{\prime}$ with $\mathcal{A}$ being $(\varepsilon,\delta)$-QDP for the pair $\sigma\sim\sigma^{x}$, and again applying the continuity result for quantum entropy, as in the proof of Proposition 12. ∎
By the data processing inequality for mutual information, we have $\mathsf{I}(X_{i};Y|X^{n\backslash i})\leq\varepsilon^{\prime}$, where $\varepsilon^{\prime}$ is determined by the values on the right-hand side of the inequalities in Proposition 13. This showcases that the setup of Proposition 13 satisfies $\varepsilon^{\prime}$-mutual information differential privacy, as proposed in [6], where a randomized mechanism $A:\mathcal{X}^{n\times k}\to\mathcal{Y}$ is defined to be $\varepsilon$-mutual information differentially private if
$\sup_{\begin{subarray}{c}P_{X}\in\mathcal{P}(\mathcal{X}^{n\times k}),\\ i\in\{1,\ldots,n\}\end{subarray}}\mathsf{I}(X_{i};A(X)|X^{n\backslash i})\leq\varepsilon.$ (259)
VIII-B Quantum Fairness and QPP
In this subsection, we see that quantum fairness can be formulated as a special case of QPP. This also encourages the design of customized fairness models depending on the application of interest, with the assistance of the QPP framework.
Quantum fairness aims to treat all input states equally, meaning that no pair of input states that are close in some sense (e.g., close in trace distance) should have a large difference in their outcomes when processed by an algorithm [24]. For a quantum decision model $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$, where a quantum channel $\mathcal{E}$ is followed by a POVM $\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}$ (i.e., a quantum algorithm described by a quantum to classical channel), quantum fairness is defined in [24] as follows.
Definition 6 ($(\alpha,\beta)$-fairness [24]).
Suppose we are given a quantum decision model $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$, two distance metrics $D(\cdot\|\cdot)$ and $d(\cdot\|\cdot)$ on $\mathcal{D}(\mathcal{H})$ and $\mathcal{D}(\mathcal{O})$ respectively. Fix $0<\alpha,\beta\leq 1$. Then the decision model $\mathcal{A}$ is $(\alpha,\beta)$ fair if for all $\rho,\sigma\in\mathcal{D}(\mathcal{H})$ such that $D(\rho\|\sigma)\leq\alpha$, then
$d\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)\leq\beta.$ (260)
Remark 17 (Privacy implying fairness).
Choose $D(\rho\|\sigma)=\left\|\rho-\sigma\right\|_{1}/2$ and $d\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)=\frac{1}{2}\sum_{i}\!\left|\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho-\sigma)\right]\right|$, which are commonly used distance metrics [24]. With the QPP framework
$\mathcal{S}$ =$=\{\rho:\rho\in\mathcal{D}(\mathcal{H})\},$ (261)
$\mathcal{Q}$ =$=\{(\rho,\sigma):\rho,\sigma\in\mathcal{D}(\mathcal{H}),\ D(\rho\|\sigma)\leq\alpha\},$
$\Theta$ =$=\{P_{X}:P_{X}\in\mathcal{P}(\mathcal{X})\},$
$\mathcal{M}$ =$=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\},$
if $\mathcal{E}$ satisfies $\varepsilon$-QPP, then the quantum decision model $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$ is also fair.
Proposition 14 (Fairness guarantee from QPP).
If $\mathcal{E}$ satisfies $\varepsilon$-QPP with $(\mathcal{S},\mathcal{Q},\Theta,{\mathcal{M}})$ defined above, then $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$ is $(\alpha,\sqrt{\varepsilon^{\prime}/2})$-fair, where $\varepsilon^{\prime}=\min\{\varepsilon,\varepsilon^{2}/2\}$.
Proof:
From Proposition 11, the fact that $\mathcal{E}$ is $\varepsilon$-QPP implies
$\|\mathcal{E}(\rho)-\mathcal{E}(\sigma)\|_{1}\leq\min\{\varepsilon,\sqrt{2\varepsilon}\}=\sqrt{2\varepsilon^{\prime}},$ (262)
for $\rho$ and $\sigma$ such that $D(\rho\|\sigma)\leq\alpha$.
Then consider the measurement channel that performs the following transformation:
$\mathcal{E}(\rho)\to\sum_{i\in\mathcal{O}}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho)\right]|i\rangle\!\langle i|.$ (263)
It follows from the data-processing inequality for the trace distance that
$\left\|\sum_{i\in\mathcal{O}}\left(\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho)\right]-\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\sigma)\right]\right)|i\rangle\!\langle i|\right\|_{1}\leq\sqrt{2\varepsilon^{\prime}}.$ (264)
This leads to
$d\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)=\frac{1}{2}\sum_{i}\left|\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho-\sigma)\right]\right|\leq\sqrt{\frac{\varepsilon^{\prime}}{2}},$ (265)
concluding the proof. ∎
This shows how $\varepsilon$-QPP also provides fairness to the decision models of interest. However, fairness guarantees may not be sufficient to guarantee privacy in general. Next, we show that when $\mathcal{M}$ satisfies a relationship related to the POVM of the quantum decision model where fairness needs to be ensured, fair algorithms also act as privacy mechanisms.
Proposition 15 (Privacy guarantees obtained from fairness).
Consider the same privacy framework as in (261) but with the modification =$\mathcal{M}=\{\cup_{i\in\mathcal{B}}\mathrm{M}_{i}|\mathcal{B}\subseteq\mathcal{O}\}$. If $\mathcal{A}=\{\mathcal{E},\{\mathrm{M}_{i}\}_{i\in\mathcal{O}}\}$ is $(\alpha,\beta)$-fair, then $\mathcal{E}$ is $(\varepsilon,2\beta)$-QPP for every $\varepsilon\geq 0$.
Proof:
Since $\mathcal{A}$ is $(\alpha,\beta)$-fair, it follows that
$\frac{1}{2}\sum_{i\in\mathcal{O}}\left|\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho-\sigma)\right]\right|\leq\beta.$ (266)
Then, for every $\mathcal{B}\subseteq\mathcal{O}$
$\left|\mathrm{Tr}\!\left[\sum_{i\in\mathcal{B}}\mathrm{M}_{i}\mathcal{E}(\rho-\sigma)\right]\right|$ $\leq\sum_{i\in\mathcal{B}}\left|\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho-\sigma)\right]\right|$ (267)
$\leq\sum_{i\in\mathcal{O}}\left|\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{E}(\rho-\sigma)\right]\right|$ (268)
$\leq 2\beta,$ (269)
where the first inequality follows from the triangular inequality, the second from $\mathcal{B}\subseteq\mathcal{O}$, and the last from (266). ∎
IX Variants of Quantum Pufferfish Privacy Framework
Inspired by variants of the classical DP framework proposed using divergences as privacy metrics [69, 70], and variants of quantum DP [10], we propose variants of quantum PP with the use of generalized quantum divergences [28] (see also [27, Definition 7.13] and recall (1)). According to [43], post-processing is an important property that should be satisfied in any privacy framework. A generalized divergence by its very definition satisfies this requirement for any processing by means of a quantum channel. Thus, quantum generalized divergences by design can be employed for defining privacy metrics.
In this section, we focus on QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, where ¯$\bar{\mathcal{M}}\coloneqq\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$. Here we formulate QPP based on generalized divergences, which we denote by $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP, where $\boldsymbol{\mathsf{D}}$ is a placeholder for the generalized divergence being used.
Definition 7 ($(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP).
Fix a privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ and $\varepsilon>0$. An algorithm $\mathcal{A}$ is $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP if
$\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\boldsymbol{\mathsf{D}}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)\leq\varepsilon,$ (270)
where $\boldsymbol{\mathsf{D}}$ is an arbitrary generalized divergence and $\rho^{\mathcal{R}}$ and $\rho^{\mathcal{T}}$ are defined as in Definition 4. Note that these two density matrices depend on elements of $\Theta$ and $\mathcal{Q}$.
Remark 18 (Properties of $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP).
$(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP satisfies the following properties, which follow from the properties of the underlying generalized divergence $\boldsymbol{\mathsf{D}}$:
  • Post-processing: Given a mechanism $\mathcal{A}$ satisfying $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M})}$, for a quantum channel $\mathcal{N}$, the processed mechanism $\mathcal{N}\circ\mathcal{A}$ also satisfies $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$.
  • Convexity: Let $\mathcal{A}_{1},\ldots,\mathcal{A}_{k}$ be $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$. Take $I$ as a $k$-ary categorical random variable with parameters $(p_{1},\ldots,p_{k})$. Then the mechanism $\mathcal{A}\coloneqq\mathcal{A}_{I}$ also satisfies $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP in the same framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, if $\boldsymbol{\mathsf{D}}$ satisfies the direct-sum property, as defined in [27, Eq. (4.3.7)], from which it follows that $\boldsymbol{\mathsf{D}}$ is jointly convex [27, Proposition 4.15].
  • Parallel composability: If $\mathcal{A}_{i}$ satisfies $(\varepsilon_{i},\delta_{i})$-QPP in $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ for $i\in\{1,\ldots,k\}$, then the composed mechanism, as defined in Theorem 1, satisfies $\left(\boldsymbol{\mathsf{D}},\sum_{i=1}^{k}\varepsilon_{i}\right)$-QPP in the framework $\left(\mathcal{S},\mathcal{Q}^{(k)},\Theta,\bar{\mathcal{M}}^{k}\right)$, if $\boldsymbol{\mathsf{D}}$ satisfies subadditivity (i.e., if $\boldsymbol{\mathsf{D}}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})\leq\boldsymbol{\mathsf{D}}(\rho_{1}\|\sigma_{1})+\boldsymbol{\mathsf{D}}(\rho_{2}\|\sigma_{2})$ for all states $\rho_{1}$, $\rho_{2}$, $\sigma_{1}$, and $\sigma_{2}$).
Remark 19 (Auditing variants of QPP).
The methodologies proposed in Section VII can be used to audit the variants of QPP (based on generalized divergences) as well. In this regard, quantum algorithms and procedures for estimating respective generalized divergence (e.g., Rényi relative entropies) would be useful. This provides more motivation for designing efficient and accurate techniques for estimating them, beyond those already established in [71].
Choosing various generalized quantum divergences, we arrive at a wide spectrum of privacy frameworks. Then we can utilize the tools of quantum information theory, particularly relationships between divergences, to characterize the relative strength of those privacy frameworks. This then provides an operational interpretation of those divergences as privacy metrics.
IX-A Variants Based on Rényi Divergences
First, we define the following quantities. The measured Rényi divergence of order $\alpha\in(0,1)\cup(1,\infty)$ is defined as [72, Eq. (17)]
${\check{\mathsf{D}}}_{\alpha}(\rho\|\sigma)\coloneqq\sup_{\mathcal{M}}\mathsf{D}^{c}_{\alpha}\!\left(\mathcal{M}(\rho)\|\mathcal{M}(\sigma)\right),$ (271)
where
$\mathsf{D}^{c}_{\alpha}(\mathcal{M}(\rho)\|\mathcal{M}(\sigma))\coloneqq~{}\frac{1}{\alpha-1}\ln\!\left(\sum_{x\in\mathcal{X}}\!\left(p(x)\right)^{\alpha}\!\left(q(x)\right)^{1-\alpha}\right),$ (272)
with $p(x)\coloneqq\mathrm{Tr}[\mathrm{M}_{x}\rho]$ and $q(x)\coloneqq\mathrm{Tr}[\mathrm{M}_{x}\sigma]$ for $\mathcal{M}$ corresponding to a POVM $\{\mathrm{M}_{x}\}_{x\in\mathcal{X}}$. The Rényi preparation divergence of order $\alpha\in(0,1)\cup(1,\infty)$ is defined as [73, 74]
${\hat{\mathsf{D}}}_{\alpha}(\rho\|\sigma)\coloneqq\inf_{P,Q,\mathcal{P}}\mathsf{D}^{c}_{\alpha}(P\|Q),$ (273)
where $\mathcal{P}$ is a classical–quantum channel, $\mathcal{P}(P)=\rho$, $\mathcal{P}(Q)=\sigma$, and the classical Rényi divergence is defined as
$\mathsf{D}^{c}_{\alpha}(P\|Q)\coloneqq~{}\frac{1}{\alpha-1}\ln\!\left(\sum_{x\in\mathcal{X}}\!\left(P(x)\right)^{\alpha}\!\left(Q(x)\right)^{1-\alpha}\right).$ (274)
The quantities in (271) and (273) satisfy the data-processing inequality for all $\alpha\in(0,1)\cup(1,\infty)$ by construction, following from this property holding for the underlying classical divergence. Moreover, the following bounds hold
$\check{\mathsf{D}}_{\alpha}(\rho\|\sigma)\leq{\mathsf{D}}_{\alpha}(\rho\|\sigma)\leq{\hat{\mathsf{D}}}_{\alpha}(\rho\|\sigma).$ (275)
where ${\mathsf{D}}_{\alpha}$ is an arbitrary quantum Rényi divergence that satisfies data processing [75, Eq. (3.7)].
Remark 20 (Relative strength of privacy metrics).
Choosing the preparation divergence $\hat{\mathsf{D}}_{\alpha}=\boldsymbol{\mathsf{D}}$ in $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP gives the strongest privacy metric that satisfies the post-processing property for the family of quantum Rényi divergences of order $\alpha$, while $\check{\mathsf{D}}_{\alpha}=\boldsymbol{\mathsf{D}}$ gives the weakest privacy metric from that same family of divergences. That is, we have that
$\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\check{\mathsf{D}}_{\alpha}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)$
$\leq\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\mathsf{D}_{\alpha}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right)$ (276)
$\leq\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\hat{\mathsf{D}}_{\alpha}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\|\mathcal{A}(\rho^{\mathcal{T}})\right),$ (277)
so that
[…]
Remark 21 (Operational interpretation as privacy metrics).
Choose
$\mathcal{S}$ =$=\{\rho,\sigma\},$ (279)
$\mathcal{Q}$ =$=\{(\rho,\sigma)\},$ (280)
$\Theta$ =$=\{\{P_{X}(x),\rho^{x}\}_{x\in\mathcal{X}}:P_{X}\in\mathcal{P}(\mathcal{X}),\ \rho^{x}\in\{\rho,\sigma\}\},$ (281)
$\mathcal{M}$ $=\bar{\mathcal{M}}.$ (282)
The strongest privacy metric that can be generated from the family of quantum Rényi divergences of order $\alpha$, which is also a generalized divergence, is devised by setting $\boldsymbol{\mathsf{D}}=\hat{\mathsf{D}}_{\alpha}$. In the chosen QPP framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$, the value of $\varepsilon$ that divides the region where $(\mathsf{D},\varepsilon)$-QPP is achieved and the region where it is violated, for an identity channel, is $\hat{\mathsf{D}}_{\alpha}(\rho\|\sigma)$.
The sandwiched Rényi relative entropy in (8) satisfies data processing for $\alpha\in[1/2,1)\cup(1,\infty)$ and the quantum relative entropy in (6) also satisfies data processing. Thus, both of these are candidates for a generalized divergence.
Proposition 16.
Fix $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, $\alpha\in[1/2,1)\cup(1,\infty)$, and $\delta\in(0,1)$. Then, for an algorithm $\mathcal{A}$, we have that
[…]
where
$\varepsilon^{\prime}$ $\coloneqq\min\!\left\{\varepsilon,\frac{\varepsilon^{2}\alpha}{2}\right\},$ (284)
$\varepsilon^{\star}$ $\coloneqq\varepsilon^{\prime}+\frac{1}{\alpha-1}\ln\!\left(\frac{1}{\delta^{2}}\right)+\ln\!\left(\frac{1}{1-\delta^{2}}\right).$ (285)
We also have
[…]
where
$\varepsilon^{\prime\prime}$ $\coloneqq\min\!\left\{\varepsilon,\frac{\varepsilon^{2}}{2}\right\},$ (287)
$K$ $\coloneqq\sup_{\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}}\mathsf{T}\!\left(\mathcal{A}(\rho^{\mathcal{R}}),\mathcal{A}(\rho^{\mathcal{T}})\right)$ (288)
$\hat{\varepsilon}$ $\coloneqq\frac{1}{\delta^{2}}\left(\varepsilon^{\prime}+K\right)+\ln\!\left(\frac{1}{1-\delta^{2}}\right).$ (289)
Proof:
The first implication in both (283) and (286) follows from Proposition 10. Then for the second implication, recall from item 3 of Proposition 2 that $\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)\leq\mathsf{D}_{\max}^{\delta}(\rho\|\sigma)$. Then applying [38, Propositions 5 and 6], we arrive at:
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $\leq\widetilde{D}_{\alpha}(\rho\|\sigma)+\frac{1}{\alpha-1}\ln\!\left(\frac{1}{\delta^{2}}\right)+\ln\!\left(\frac{1}{1-\delta^{2}}\right),$ (290)
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)$ $\leq\frac{1}{\delta^{2}}\left(\mathsf{D}(\rho\|\sigma)+\mathsf{T}(\rho,\sigma)\right)+\ln\!\left(\frac{1}{1-\delta^{2}}\right).$ (291)
Finally, to conclude the proof, invoke Proposition 1 to establish the required relationship to the QPP framework from there. ∎
Note that the chain of implications in (283) holds for every Rényi divergence $\mathsf{D}_{\alpha}$ satisfying data processing, beyond just $\widetilde{\mathsf{D}}_{\alpha}$, because data processing is the key property to adapt Proposition 10 and [38, Proposition 6] (see Eqs. (K51) and (K52) therein).
Remark 22 (Comparison to existing results for QDP and classical DP).
For the specific instantiation of QDP, Lemmas V.4 and V.5 of [10] show that
[…]
with
$\bar{\varepsilon}\coloneqq\varepsilon+\frac{\ln\left(1/(1-\sqrt{1-\delta^{2}})\right)}{\alpha-1}\approx\varepsilon+\frac{\ln(2/\delta^{2})}{\alpha-1},$ (293)
where the approximation holds for small $\delta$. Our first implication in (283) is tight compared to this since $\varepsilon^{\prime}\leq\varepsilon$ and the second implication is also tight (i.e, $\varepsilon^{\star}\leq\bar{\varepsilon}$) if we choose $\delta^{2}\leq 1-2^{\left(-\frac{1}{\alpha-1}\right)}$ in the small $\delta$ regime.
IX-B Variant Incorporating Entanglement
In this subsection, we propose a variant of QPP that incorporates reference systems, which enables us to observe the effect of entanglement on the privacy of the system of interest. This variant is applicable in identifying whether there is an advantage in designing private systems with the association of other quantum systems.
Definition 8 ($(\boldsymbol{\mathsf{D}}^{R},\varepsilon)$-QPP with reference systems).
With the same setup $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ in Definition 4, a quantum algorithm $\mathcal{A}:\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ is $(\varepsilon,\boldsymbol{\mathsf{D}})$-QPP with reference systems if
$\sup_{\Theta,(\omega_{RA}^{\mathcal{R}},\omega_{RA}^{\mathcal{T}})\in\mathcal{G}}\boldsymbol{\mathsf{D}}\!\left((\mathcal{I}_{R}\otimes\mathcal{A})(\omega_{RA}^{\mathcal{R}})\|(\mathcal{I}_{R}\otimes\mathcal{A})(\omega_{RA}^{\mathcal{T}})\right)\leq\varepsilon,$ (294)
where the set $\mathcal{G}$ is defined in (27).
Since there is no restriction on the reference systems, the supremum in Definition 8 is also taken over the dimension of the reference system $R$. Thus, the supremum here is over an unbounded reference system $R$. However, following from isometric invariance of generalized divergences, together with purification and the Schmidt decomposition, it suffices to take the supremum over pure bipartite states with the reference system $R$ isomorphic to the channel input system $A$. Similar to the QPP framework, and other variants of QPP, the above variant incorporating reference systems also satisfies post-processing, convexity, and parallel composability (similar to Remark 18).
Remark 23 (Adaptive composition with reference systems).
Let $\mathcal{A}_{i}$ be a $(\boldsymbol{\mathsf{D}}^{R},\varepsilon_{i})$-QPP algorithm for $i\in\{1,2\}$. Consider the adaptive composition of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, as shown in Fig. 10: First $\mathcal{A}_{1}$ is applied on the upper system, then the quantum channel $\mathcal{N}$ acts on both systems, and lastly $\mathcal{A}_{2}$ acts on the upper system. Adaptive composition of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ also satisfies $(\boldsymbol{\mathsf{D}}^{R},\varepsilon_{1})$-QPP. This shows that the adaptive composition illustrated in Fig. 10 does not degrade privacy, showcasing the strength of the privacy framework in Definition 8.
Figure 10: Adaptive composition with reference systems: First $\mathcal{A}_{1}$ is applied on the upper system, then the quantum channel $\mathcal{N}$ on both the systems, and lastly $\mathcal{A}_{2}$ on the upper system. Then the adaptive composition is QPP if $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are.
Next, we observe that Definition 8 is a stronger privacy guarantee than Definition 7, which does not take into account reference systems.
Corollary 3 (Strength compared to $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP).
$(\boldsymbol{\mathsf{D}}^{R},\varepsilon)$-QPP implies $(\boldsymbol{\mathsf{D}},\varepsilon)$-QPP.
This follows from the data-processing inequality for the underlying generalized divergence $\boldsymbol{\mathsf{D}}$, choosing the channel as the partial trace over the reference system $R$. Also it indicates that in certain scenarios where the system of interest has entanglement with some other reference systems, general QPP guarantees defined in earlier sections may not be sufficient. It is important to note that the choice of $\boldsymbol{\mathsf{D}}$ could heavily affect the design of useful privacy frameworks with entanglement (recall the example in Remark 3 with the choice $\boldsymbol{\mathsf{D}}=\mathsf{D}_{\max}$, along with the equivalence of $\mathsf{D}_{\max}$ to $\varepsilon$-QPP, as given in (225)).
X Concluding Remarks and Future Directions
In conclusion, our proposed notion of QPP provides a flexible privacy framework for quantum systems. It gives rise to an operational interpretation of the DL divergence, which was used to study properties of QPP mechanisms and characterize privacy-utility tradeoffs. We explored quantum privacy mechanisms (e.g., via the depolarization channel) and presented a methodology for auditing quantum privacy.
Future research directions are abundant and include privacy-utility analysis of specific quantum estimation tasks, designing efficient quantum algorithms that achieve QPP, analysis of the type-II error of our quantum privacy auditing pipeline, tight characterizations of parallel composability of QPP mechanisms, and many more. Another interesting research direction is to devise efficient methods for computing the DL divergence using quantum algorithms, which would enable the auditing of quantum privacy. In the longer term, the proposed framework could lay the foundations for privacy-preserving learning in quantum systems, and we aim in future work to explore various directions in that space.
Acknowledgements—We acknowledge helpful discussions with Rochisha Agarwal, Hansadi Jayamaha, Kaiyuan Ji, Margarite LaBorde, Hemant Mishra, Dhrumil Patel, Aby Philip, Soorya Rethinasamy, and Vishal Singh. We are also indebted to Prof. Çişik Balyk for sharing his valuable insights. TN and MMW acknowledge support from the National Science Foundation under Grant No. 1907615. ZG is partially supported by NSF grants CCF-2046018, DMS-2210368, and CCF-2308446, and the IBM Academic Award.
Appendix A Proof of Lemma 2
The proof given below is closely related to the proof of [65, Lemma 6.21], but there are some subtle differences and so we provide it here for completeness.
Let
$\Sigma$ $\coloneqq\left(\rho-\lambda\sigma\right)_{+},$ (295)
$G$ $\coloneqq\left(\lambda\sigma\right)^{1/2}\left(\lambda\sigma+\Sigma\right)^{-1/2}.$ (296)
Note that
00 0 $\leq G^{{\dagger}}G$ (297)
$=\left(\lambda\sigma+\Sigma\right)^{-1/2}\left(\lambda\sigma\right)\left(\lambda\sigma+\Sigma\right)^{-1/2}$ (298)
$\leq\left(\lambda\sigma+\Sigma\right)^{-1/2}\left(\lambda\sigma+\Sigma\right)\left(\lambda\sigma+\Sigma\right)^{-1/2}$ (299)
$\leq I.$ (300)
From the fact that $\rho-\lambda\sigma\leq\Sigma$, it follows that
$\rho\leq\lambda\sigma+\Sigma.$ (301)
Define the following state:
$\widetilde{\rho}\coloneqq\frac{G\rho G^{{\dagger}}}{\operatorname{Tr}[G^{{\dagger}}G\rho]}.$ (302)
Consider that
$1-\operatorname{Tr}[G\rho G^{{\dagger}}]$
$=\operatorname{Tr}[(I-G^{{\dagger}}G)\rho]$ (303)
$\leq\operatorname{Tr}[(I-G^{{\dagger}}G)\left(\lambda\sigma+\Sigma\right)]$ (304)
$=\operatorname{Tr}[(I-G^{{\dagger}}G)\left(\lambda\sigma+\Sigma\right)]$ (305)
$=\operatorname{Tr}[\lambda\sigma+\Sigma]$
$\quad-\operatorname{Tr}\!\left[\left(\lambda\sigma+\Sigma\right)^{-1/2}\left(\lambda\sigma\right)\left(\lambda\sigma+\Sigma\right)^{-1/2}\left(\lambda\sigma+\Sigma\right)\right]$ (306)
$=\operatorname{Tr}[\lambda\sigma+\Sigma]-\operatorname{Tr}\!\left[\lambda\sigma\right]$ (307)
$=\operatorname{Tr}[\Sigma]$ (308)
$=\delta$ (309)
This implies that
$\operatorname{Tr}[G\rho G^{{\dagger}}]\geq 1-\delta.$ (310)
Then it follows that
$\widetilde{\rho}$ $=\frac{G\rho G^{{\dagger}}}{\operatorname{Tr}[G^{{\dagger}}G\rho]}$ (311)
$\leq\frac{G\left(\lambda\sigma+\Sigma\right)G^{{\dagger}}}{\operatorname{Tr}[G^{{\dagger}}G\rho]}$ (312)
$=\frac{\lambda\sigma}{\operatorname{Tr}[G^{{\dagger}}G\rho]}$ (313)
$\leq\frac{\lambda\sigma}{1-\delta}.$ (314)
Let $\psi_{RA}=\sqrt{\rho_{A}}\Gamma_{RA}\sqrt{\rho_{A}}$ be the canonical purification of $\rho_{A}$, with $\Gamma_{RA}\coloneqq\sum_{i,j}|i\rangle\!\langle j|_{R}\otimes|i\rangle\!\langle j|_{A}$, and let $\widetilde{\psi}_{RA}=\frac{G_{A}\psi_{RA}G_{A}^{{\dagger}}}{\operatorname{Tr}[G^{{\dagger}}G\rho]}$ purify $\widetilde{\rho}$. Then
$\sqrt{F(\rho,\widetilde{\rho})}$ $\geq\frac{1}{\operatorname{Tr}[G^{{\dagger}}G\rho]}\left|\langle\psi|_{RA}I_{R}\otimes G_{A}|\psi\rangle_{RA}\right|$ (315)
$\geq\left|\langle\psi|_{RA}I_{R}\otimes G_{A}|\psi\rangle_{RA}\right|$ (316)
$=\left|\langle\Gamma|_{RA}I_{R}\otimes\sqrt{\rho_{A}}G_{A}\sqrt{\rho_{A}}|\Gamma\rangle_{RA}\right|$ (317)
$=\left|\operatorname{Tr}[G\rho]\right|$ (318)
$\geq\operatorname{Re}[\operatorname{Tr}[G\rho]]$ (319)
$=\operatorname{Tr}[\overline{G}\rho]$ (320)
$=1-\operatorname{Tr}[\left(I-\overline{G}\right)\rho],$ (321)
where
$\overline{G}\coloneqq\frac{G+G^{{\dagger}}}{2}.$ (322)
The first inequality follows from Uhlmann’s theorem for fidelity, and the second follows because $\operatorname{Tr}[G^{{\dagger}}G\rho]\leq 1$. Observe that $\overline{G}\leq I$ because $\left\|G\right\|_{\infty}\leq 1$ and by applying the triangle inequality. So this means that $I-\overline{G}\geq 0$. Now consider that
$\operatorname{Tr}[\left(I-\overline{G}\right)\rho]$
$\leq\operatorname{Tr}[\left(I-\overline{G}\right)\left(\lambda\sigma+\Sigma\right)]$ (323)
$=\operatorname{Tr}[\lambda\sigma+\Sigma]-\operatorname{Tr}[\overline{G}\left(\lambda\sigma+\Sigma\right)]$ (324)
$=\operatorname{Tr}[\lambda\sigma+\Sigma]$
$-\frac{1}{2}\operatorname{Tr}\Big{[}\left(\left(\lambda\sigma\right)^{1/2}\left(\lambda\sigma+\Sigma\right)^{-1/2}+\left(\lambda\sigma+\Sigma\right)^{-1/2}\left(\lambda\sigma\right)^{1/2}\right)\times$
$\qquad\left(\lambda\sigma+\Sigma\right)\Big{]}$ (325)
$=\operatorname{Tr}[\lambda\sigma+\Sigma]-\frac{1}{2}\operatorname{Tr}\left[\left(\lambda\sigma\right)^{1/2}\left(\lambda\sigma+\Sigma\right)^{-1/2}\left(\lambda\sigma+\Sigma\right)\right]$
$\qquad-\frac{1}{2}\operatorname{Tr}\!\left[\left(\lambda\sigma+\Sigma\right)^{-1/2}\left(\lambda\sigma\right)^{1/2}\left(\lambda\sigma+\Sigma\right)\right]$ (326)
$=\operatorname{Tr}[\lambda\sigma+\Sigma]-\frac{1}{2}\operatorname{Tr}\left[\left(\lambda\sigma\right)^{1/2}\left(\lambda\sigma+\Sigma\right)^{1/2}\right]$
$\qquad-\frac{1}{2}\operatorname{Tr}\!\left[\left(\lambda\sigma\right)^{1/2}\left(\lambda\sigma+\Sigma\right)^{1/2}\right]$ (327)
$=\operatorname{Tr}[\lambda\sigma+\Sigma]-\operatorname{Tr}\!\left[\left(\lambda\sigma\right)^{1/2}\left(\lambda\sigma+\Sigma\right)^{1/2}\right]$ (328)
$\leq\operatorname{Tr}[\lambda\sigma+\Sigma]-\operatorname{Tr}\!\left[\left(\lambda\sigma\right)^{1/2}\left(\lambda\sigma\right){}^{1/2}\right]$ (329)
$=\operatorname{Tr}[\lambda\sigma+\Sigma]-\operatorname{Tr}\!\left[\lambda\sigma\right]$ (330)
$=\operatorname{Tr}[\Sigma]$ (331)
$=\delta.$ (332)
So all of this implies that
$\sqrt{\mathsf{F}(\rho,\widetilde{\rho})}\geq 1-\delta,$ (333)
and in turn that
$\mathsf{F}(\rho,\widetilde{\rho})\geq\left(1-\delta\right)^{2}.$ (334)
By applying the inequality
$\frac{1}{2}\left\|\rho-\widetilde{\rho}\right\|_{1}\leq\sqrt{1-\mathsf{F}(\rho,\widetilde{\rho})},$ (335)
we conclude that
$\frac{1}{2}\left\|\rho-\widetilde{\rho}\right\|_{1}$ $\leq\sqrt{1-\left(1-\delta\right)^{2}}$ (336)
$=\sqrt{1-\left(1-2\delta+\delta^{2}\right)}$ (337)
$=\sqrt{2\delta-\delta^{2}}$ (338)
$=\sqrt{\delta\left(2-\delta\right)}.$ (339)
Putting everything together, we see that $\widetilde{\rho}$ is a quantum state satisfying
$\frac{1}{2}\left\|\rho-\widetilde{\rho}\right\|_{1}$ $\leq\sqrt{\delta\left(2-\delta\right)},$ (340)
$\widetilde{\rho}$ $\leq\frac{\lambda\sigma}{1-\delta}.$ (341)
This means that $\widetilde{\rho}$ and $\frac{\lambda}{1-\delta}$ are feasible for ${\mathsf{D}_{\max}^{\sqrt{\delta\left(2-\delta\right)}}(\rho\|\sigma)}$, and so it follows that
$\mathsf{D}_{\max}^{\sqrt{\delta\left(2-\delta\right)}}(\rho\|\sigma)$ $\leq\ln\!\left(\frac{\lambda}{1-\delta}\right)$ (342)
$=\ln\lambda+\ln\!\left(\frac{1}{1-\delta}\right).$ (343)
This concludes the proof.
Appendix B Subadditivity of Smooth Max-Relative Entropy
Lemma 5.
Given $\delta_{1},\delta_{2}\in\left[0,1\right]$, states $\rho_{1}$ and $\rho_{2}$, and PSD operators $\sigma_{1}$ and $\sigma_{2}$, the following subadditivity relation holds
$\mathsf{D}_{\max}^{\delta_{1}+\delta_{2}}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})\leq\mathsf{D}_{\max}^{\delta_{1}}(\rho_{1}\|\sigma_{1})+\mathsf{D}_{\max}^{\delta_{2}}(\rho_{2}\|\sigma_{2}).$ (344)
Proof:
Let $\overline{\rho}_{i}$ and $\lambda_{i}$ be optimal choices for $D_{\max}^{\delta_{i}}(\rho_{i}\|\sigma_{i})$, for $i\in\left\{1,2\right\}$. Then, consider that
$\overline{\rho}_{1}\otimes\overline{\rho}_{2}\leq\lambda_{1}\sigma_{1}\otimes\overline{\rho}_{2}\leq\lambda_{1}\sigma_{1}\otimes\lambda_{2}\sigma_{2}=\lambda_{1}\lambda_{2}\sigma_{1}\otimes\sigma_{2}.$ (345)
Furthermore, consider that
$\frac{1}{2}\left\|\overline{\rho}_{1}\otimes\overline{\rho}_{2}-\rho_{1}\otimes\rho_{2}\right\|_{1}$
$=\frac{1}{2}\left\|\overline{\rho}_{1}\otimes\overline{\rho}_{2}-\overline{\rho}_{1}\otimes\rho_{2}+\overline{\rho}_{1}\otimes\rho_{2}-\rho_{1}\otimes\rho_{2}\right\|_{1}$ (346)
$=\frac{1}{2}\left\|\overline{\rho}_{1}\otimes\left(\overline{\rho}_{2}-\rho_{2}\right)+\left(\overline{\rho}_{1}-\rho_{1}\right)\otimes\rho_{2}\right\|_{1}$ (347)
$\leq\frac{1}{2}\left\|\overline{\rho}_{1}\otimes\left(\overline{\rho}_{2}-\rho_{2}\right)\right\|_{1}+\frac{1}{2}\left\|\left(\overline{\rho}_{1}-\rho_{1}\right)\otimes\rho_{2}\right\|_{1}$ (348)
$=\frac{1}{2}\left\|\overline{\rho}_{1}\right\|_{1}\left\|\overline{\rho}_{2}-\rho_{2}\right\|_{1}+\frac{1}{2}\left\|\overline{\rho}_{1}-\rho_{1}\right\|_{1}\left\|\rho_{2}\right\|_{1}$ (349)
$=\frac{1}{2}\left\|\overline{\rho}_{2}-\rho_{2}\right\|_{1}+\frac{1}{2}\left\|\overline{\rho}_{1}-\rho_{1}\right\|_{1}$ (350)
$\leq\delta_{1}+\delta_{2},$ (351)
where (346) follows from the triangular inequality for the trace norm and the final inequality from the assumption that $\overline{\rho}_{i}$ are the optimizers for $D_{\max}^{\delta_{i}}(\rho_{i}\|\sigma_{i})$, for $i\in\left\{1,2\right\}$.
Finally we have shown that $\overline{\rho}_{1}\otimes\overline{\rho}_{2}$ and $\lambda_{1}\lambda_{2}$ are candidates for the optimization for $D_{\max}^{\delta_{1}+\delta_{2}}(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2})$, thus concluding the proof. ∎
Appendix C Proof of Theorem 1
For (1): Fix $(\rho^{\mathcal{R}},\rho^{\mathcal{T}})$ in (23), $\mathrm{M}\in\mathcal{M}$. Consider
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]$ $=\mathrm{Tr}\!\left[\mathrm{M}\sum_{i=1}^{k}p_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}})\right]$ (352)
$=\sum_{i=1}^{k}p_{i}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}_{i}(\rho^{\mathcal{R}})\right]$ (353)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{i=1}^{k}p_{i}\left(e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}_{i}(\rho^{\mathcal{T}})\right]+\delta\right)$ (354)
$=\sum_{i=1}^{k}p_{i}e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}_{i}(\rho^{\mathcal{T}})\right]+\delta$ (355)
$=e^{\varepsilon}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\delta,$ (356)
where (a) follows due to each $\mathcal{A}_{i}$ being $(\varepsilon,\delta)$-QPP. This relation is true for every $(\rho^{\mathcal{R}},\rho^{\mathcal{T}})$, and so it is true for all such pairs generated from $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$.
For (2): Fix $0\leq\mathrm{M}^{\prime}\leq\mathrm{I}$ such that $\mathrm{M}^{\prime}\in\mathcal{M}^{\prime}$ as stated in the property. With that assumption, there exists $\mathrm{M}\in\mathcal{M}$ such that $\mathrm{M}=\mathcal{N}^{\dagger}(\mathrm{M}^{\prime})$. Consider that
$\mathrm{Tr}\!\left[\mathrm{M}^{\prime}\mathcal{N}\!\left(\mathcal{A}(\rho^{\mathcal{R}})\right)\right]$ $=\mathrm{Tr}\!\left[\mathcal{N}^{\dagger}(\mathrm{M}^{\prime})\mathcal{A}(\rho^{\mathcal{R}})\right]$ (357)
$=\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right],$ (358)
where $\mathcal{N}^{\dagger}$ is the adjoint of $\mathcal{N}$, implying that
$0\leq\mathcal{N}^{\dagger}(\mathrm{M}^{\prime})=\mathrm{M}\leq\mathrm{I}$ (359)
because $\mathcal{N}^{\dagger}$ is positive and unital by the assumption $\mathcal{N}$ is a quantum channel. Similarly, we have that $\mathrm{Tr}\!\left[\mathrm{M}^{\prime}\mathcal{N}\left(\mathcal{A}(\rho^{\mathcal{T}})\right)\right]=\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]$, and we conclude that the processed mechanism satisfies $(\varepsilon,\delta)$-QPP with the choice of $\mathcal{M}^{\prime}\subseteq\{\mathrm{M}^{\prime}:\mathcal{N}^{\dagger}(\mathrm{M}^{\prime})\in\mathcal{M}\}$.
For (3): Fix $(\mathcal{R}_{i},\mathcal{T}_{i})\in\mathcal{Q}$ for $i\in\{1,\ldots,k\}$, and $\ \bigotimes_{i=1}^{k}\mathrm{M}_{i}\in\bigotimes_{i=1}^{k}\mathcal{M}_{i}$. Denote
$\mathcal{A}^{(k)}(\rho^{\mathcal{R}^{(k)}})\coloneqq\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}})\otimes\mathcal{A}_{2}(\rho^{\mathcal{R}_{2}})\otimes\ldots\mathcal{A}_{k}(\rho^{\mathcal{R}_{k}})$ (360)
and $\mathcal{A}^{\otimes k}(\rho^{\mathcal{T}^{(k)}})$ similarly by replacing $\mathcal{R}$ with $\mathcal{T}$.
Fix $i\in\{1,\ldots,k\}$. Consider that $\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]\leq 1\leq 1+\delta_{i}$ because $\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]$ is a probability. Combining with the inequality $\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]\leq e^{\varepsilon_{i}}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{T}_{i}})\right]+\delta_{i}$, which holds from the assumption that QPP holds, we conclude that
$\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]\leq\min\!\left\{e^{\varepsilon_{i}}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{T}_{i}})\right],1\right\}+\delta_{i}.$ (361)
Consider that
$\prod_{i=1}^{k}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]$
$\leq\left(\min\!\left\{1,e^{\varepsilon_{1}}\mathrm{Tr}\!\left[\mathrm{M}_{1}\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}})\right]\right\}+\delta_{1}\right)\prod_{i=2}^{k}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]$ (362)
$\leq\min\!\left\{1,e^{\varepsilon_{1}}\mathrm{Tr}\!\left[\mathrm{M}_{1}\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}})\right]\right\}\prod_{i=2}^{k}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]+\delta_{1}$ (363)
$\leq\min\!\left\{1,e^{\varepsilon_{1}}\mathrm{Tr}\!\left[\mathrm{M}_{1}\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}})\right]\right\}\times$
$\quad\left(\min\!\left\{1,e^{\varepsilon_{2}}\mathrm{Tr}\!\left[\mathrm{M}_{2}\mathcal{A}_{2}(\rho^{\mathcal{T}_{2}})\right]\right\}+\delta_{2}\right)\prod_{i=3}^{k}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]$
$\qquad+\delta_{1}$ (364)
$\leq\prod_{j=1}^{2}\min\!\left\{e^{\varepsilon_{j}}\mathrm{Tr}\!\left[\mathrm{M}_{j}\mathcal{A}_{j}(\rho^{\mathcal{T}_{j}})\right],1\right\}\prod_{i=3}^{k}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{R}_{i}})\right]$
$\qquad+\delta_{1}+\delta_{2}$ (365)
$\leq e^{\sum_{i=1}^{k}\varepsilon_{i}}\prod_{i=1}^{k}\mathrm{Tr}\!\left[\mathrm{M}_{i}\mathcal{A}_{i}(\rho^{\mathcal{T}_{i}})\right]+\sum_{i=1}^{k}\delta_{i},$ (366)
where the last inequality follows by proceeding with similar expansions for each remaining term of the product as carried out in the first three steps.
Appendix D Proof of Proposition 3
Fix $\mathrm{M}^{\prime},\mathrm{M}\in\bar{\mathcal{M}}$, $P_{X}\in\Theta$, and $(\mathcal{R}_{1},\mathcal{T}_{1}),(\mathcal{R}_{2},\mathcal{T}_{2})~{}\in\mathcal{Q}$. Let $\mathrm{M}_{y}\coloneqq(\langle y|\otimes\mathrm{I})\mathrm{M}^{\prime}(|y\rangle\otimes\mathrm{I})$ and note that $\mathrm{M}_{y}$ is a measurement operator in $\bar{\mathcal{M}}$. Recall the definition of the channel
¯$\overline{\mathcal{E}}\coloneqq\sum_{y\in\mathcal{Y}}\mathcal{E}^{y}.$ (367)
Consider that
$\operatorname{Tr}\!\left[(\mathrm{M}\otimes\mathrm{M}^{\prime})\left(\sum_{y\in\mathcal{Y}}\mathcal{E}^{y}(\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}}))\otimes|y\rangle\!\langle y|\otimes\mathcal{A}_{2}^{y}(\rho^{\mathcal{R}_{2}})\right)\right]$
$=\sum_{y\in\mathcal{Y}}\operatorname{Tr}\!\left[\mathrm{M}\mathcal{E}^{y}(\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}}))\otimes\mathrm{M}^{\prime}\!\left(|y\rangle\!\langle y|\otimes\mathcal{A}_{2}^{y}(\rho^{\mathcal{R}_{2}})\right)\right]$ (368)
$=\sum_{y\in\mathcal{Y}}\operatorname{Tr}[\mathrm{M}\mathcal{E}^{y}(\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}}))]\operatorname{Tr}[\mathrm{M}^{\prime}\!\left(|y\rangle\!\langle y|\otimes\mathcal{A}_{2}^{y}(\rho^{\mathcal{R}_{2}})\right)]$ (369)
$=\sum_{y\in\mathcal{Y}}\operatorname{Tr}[\mathrm{M}\mathcal{E}^{y}(\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}}))]\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{R}_{2}})]$ (370)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{y\in\mathcal{Y}}\operatorname{Tr}[\mathrm{M}\mathcal{E}^{y}(\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}}))]$
×$\qquad\qquad\times\left(\min\!\left\{1,e^{\varepsilon_{2}}\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}_{2}})]\right\}+\delta_{2}\right)$ (371)
$=\sum_{y\in\mathcal{Y}}\operatorname{Tr}[\mathcal{E}^{y{\dagger}}(\mathrm{M})\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}})]$
×$\qquad\qquad\times\left(\min\!\left\{1,e^{\varepsilon_{2}}\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}_{2}})]\right\}+\delta_{2}\right)$ (372)
=$\stackrel{{\scriptstyle(b)}}{{=}}\operatorname{Tr}[\mathrm{M}\overline{\mathcal{E}}(\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}}))]\delta_{2}$
$+\sum_{y\in\mathcal{Y}}\left(\operatorname{Tr}[\mathcal{E}^{y{\dagger}}(\mathrm{M})\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}})]\min\!\left\{1,e^{\varepsilon_{2}}\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}_{2}})]\right\}\right)$ (373)
$\stackrel{{\scriptstyle(c)}}{{\leq}}\delta_{2}$
$\ +\sum_{y\in\mathcal{Y}}\operatorname{Tr}[\mathcal{E}^{y{\dagger}}(\mathrm{M})\mathcal{A}_{1}(\rho^{\mathcal{R}_{1}})]\min\!\left\{1,e^{\varepsilon_{2}}\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}_{2}})]\right\}$ (374)
$\stackrel{{\scriptstyle(d)}}{{\leq}}\delta_{2}+\sum_{y\in\mathcal{Y}}\left(e^{\varepsilon_{1}}\operatorname{Tr}[\mathcal{E}^{y{\dagger}}(\mathrm{M})\mathcal{A}_{1}(\rho^{\mathcal{T}_{1}})]+\delta_{1}\right)$
×$\qquad\qquad\qquad\qquad\times\min\!\left\{1,e^{\varepsilon_{2}}\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}_{2}})]\right\}$ (375)
$\leq\sum_{y\in\mathcal{Y}}\left(e^{\varepsilon_{1}}\operatorname{Tr}[\mathcal{E}^{y{\dagger}}(\mathrm{M})\mathcal{A}_{1}(\rho^{\mathcal{T}_{1}})]e^{\varepsilon_{2}}\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}_{2}})]+\delta_{1}\right)$
$\qquad\qquad+\delta_{2}$ (376)
$=\sum_{y\in\mathcal{Y}}\left(e^{\varepsilon_{2}}e^{\varepsilon_{1}}\operatorname{Tr}[\mathcal{E}^{y{\dagger}}(\mathrm{M})\mathcal{A}_{1}(\rho^{\mathcal{T}_{1}})]\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}_{2}})]+\delta_{1}\right)$
$\qquad\qquad+\delta_{2}$ (377)
$=e^{\varepsilon^{\prime}}\times$
$\operatorname{Tr}\!\left[(\mathrm{M}\otimes\mathrm{M}^{\prime})\!\left(\sum_{y\in\mathcal{Y}}\mathcal{E}^{y}(\mathcal{A}_{1}(\rho^{\mathcal{T}_{1}}))\otimes|y\rangle\!\langle y|\otimes\mathcal{A}_{2}^{y}(\rho^{\mathcal{T}_{2}})\right)\right]$
$\qquad\qquad+\delta_{2}+\delta_{1}\left|\mathcal{Y}\right|,$ (378)
where: (a) from $\operatorname{Tr}[\mathrm{M}_{y}^{\prime}\mathcal{A}_{2}^{y}(\rho^{\mathcal{R}_{2}})]\leq 1\leq 1+\delta_{2}$ and $\mathcal{A}_{2}^{y}$ being $(\varepsilon_{2},\delta_{2})$-QPP; (b) and (c) from
$\operatorname{Tr}\!\left[\sum_{y\in\mathcal{Y}}\mathcal{E}^{y{\dagger}}(\mathrm{M})\mathcal{A}_{1}(\rho^{\mathcal{T}_{1}})\right]$ $=\operatorname{Tr}\!\left[\mathrm{M}\sum_{y\in\mathcal{Y}}\mathcal{E}^{y}\mathcal{A}_{1}(\rho^{\mathcal{T}_{1}})\right]$ (379)
$=\operatorname{Tr}[\mathrm{M}\overline{\mathcal{E}}(\mathcal{A}_{1}(\rho^{\mathcal{T}_{1}}))]$ (380)
$\leq 1;$ (381)
and (d) from $\mathcal{A}_{1}$ being $(\varepsilon_{1},\delta_{1})$-QPP and the fact that $\mathcal{E}^{y{\dagger}}(\mathrm{M})$ is a measurement operator in $\bar{\mathcal{M}}$.
Appendix E Proof of Lemma 4
The proof follows analogously to the classical version of this bound in [70, Proposition 3.3], along with the upper bound for an arbitrary $\mathsf{D}_{\alpha}(\cdot\|\cdot)$ satisfying data processing. Set $\alpha>1$. Then, for such $\mathsf{D}_{\alpha}(\cdot\|\cdot)$, from [65, Equation 4.34], which is obtained by choosing a specific preparation channel, we have that
$\mathsf{D}_{\alpha}(\rho\|\sigma)\leq\frac{1}{\alpha-1}\log\mathrm{Tr}\!\left[\sigma^{1/2}\!\left(\sigma^{-1/2}\rho\sigma^{-1/2}\right)^{\alpha}\sigma^{1/2}\right].$ (382)
Let us use the following substitution:
$\mathsf{D}_{T}(\rho\|\sigma)=\varepsilon.$ (383)
Then we have $\mathsf{D}_{\max}(\rho\|\sigma)\leq\varepsilon$ and $\mathsf{D}_{\max}(\sigma\|\rho)\leq\varepsilon$. Moreover, with the definition of $\mathsf{D}_{\max}(\cdot\|\cdot)$ in (14), we have $\rho\leq e^{\varepsilon}\sigma$ and $\sigma\leq e^{\varepsilon}\rho$. Then we find that
$e^{-\varepsilon}\mathrm{I}\leq\sigma^{-1/2}\rho\sigma^{-1/2}\leq e^{\varepsilon}\mathrm{I}.$ (384)
Suppose that $\sigma^{-1/2}\rho\sigma^{-1/2}$ has the following spectral decomposition $\sum_{i}t_{i}|\phi_{i}\rangle\!\langle\phi_{i}|$. Then $e^{-\varepsilon}\leq t_{i}\leq e^{\varepsilon}$, and so for all $i$, $\exists\lambda_{i}\in[0,1]$ such that
$t_{i}=\lambda_{i}e^{\varepsilon}+(1-\lambda_{i})e^{-\varepsilon}.$ (385)
Consider that
$e^{(\alpha-1)\mathsf{D}_{\alpha}(\rho\|\sigma)}$
$\leq\mathrm{Tr}\!\left[\sigma^{1/2}\!\left(\sigma^{-1/2}\rho\sigma^{-1/2}\right)^{\alpha}\sigma^{1/2}\right]$ (386)
$=\mathrm{Tr}\!\left[\sigma^{1/2}\sum_{i}\!\left(\lambda_{i}e^{\varepsilon}+(1-\lambda_{i})e^{-\varepsilon}\right)^{\alpha}|\phi_{i}\rangle\!\langle\phi_{i}|\ \sigma^{1/2}\right]$ (387)
$=\sum_{i}\!\left(\lambda_{i}e^{\varepsilon}+(1-\lambda_{i})e^{-\varepsilon}\right)^{\alpha}\mathrm{Tr}\!\left[\sigma|\phi_{i}\rangle\!\langle\phi_{i}|\right]$ (388)
$\leq\sum_{i}\!\left(\lambda_{i}e^{\varepsilon\alpha}+(1-\lambda_{i})e^{-\varepsilon\alpha}\right)\mathrm{Tr}\!\left[\sigma|\phi_{i}\rangle\!\langle\phi_{i}|\right]$ (389)
$=e^{\varepsilon\alpha}c_{1}+e^{-\varepsilon\alpha}c_{2}$ (390)
where the first inequality follows from the inequality in (382), the second from the convexity of the function $x^{\alpha}$ for $\alpha>1$, and the definitions
$c_{1}$ $\coloneqq\sum_{i}\lambda_{i}\mathrm{Tr}\!\left[\sigma|\phi_{i}\rangle\!\langle\phi_{i}|\right],$ (391)
$c_{2}$ $\coloneqq\sum_{i}(1-\lambda_{i})\mathrm{Tr}\!\left[\sigma|\phi_{i}\rangle\!\langle\phi_{i}|\right].$ (392)
In (390), we arrive at a function of $\alpha$ (i.e., $e^{\varepsilon\alpha}c_{1}+e^{-\varepsilon\alpha}c_{2}$). Observing that $c_{1}+c_{2}=1$, we can find $c_{1}$ and $c_{2}$ by evaluating this function of $\alpha$ at $\alpha=1$, which turns out to be equal to one because
$\sum_{i}t_{i}\mathrm{Tr}[\sigma|\phi_{i}\rangle\!\langle\phi_{i}|]=\mathrm{Tr}[\rho]=1,$ (393)
where $t_{i}$ is given in (385). Proceeding with this we get $c_{1}=\frac{1-e^{-\varepsilon}}{e^{\varepsilon}-e^{-\varepsilon}}$. Then collecting all these relations and simplifying we obtain,
$e^{(\alpha-1)\mathsf{D}_{\alpha}(\rho\|\sigma)}\leq\frac{\sinh(\alpha\varepsilon)-\sinh\!\left((\alpha-1)\varepsilon\right)}{\sinh(\varepsilon)}.$ (394)
Together with [70, Lemma B.1], we can further bound (394) from above by $e^{\alpha(\alpha-1)\varepsilon^{2}/2}$. With the substitution in (383) we conclude the proof.